A Complete Bibliography of Publications in Annals of Statistics (1970–1979)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

19 May 2021 Version 1.05

Title word cross-reference

(M, S) [Jac78]. (v, k, λ) [RF73]. 2 [Lor76]. 2^{m} [Shi76c, Shi76d]. 2^{n} [RF73]. 2l [Shi76c]. 6 ≤ m ≤ 8 [Shi76d]. [0, 1] [Wit75]. α [Sha78]. C(α) [Ray74]. χ_{r}^{2} [Jen74]. D [Atw73, Paz74]. F [AHS78, HRS75]. F (4t; 2t, 2t) [Fed77]. h [TG74]. K [KR77, OS79, Con73, Goo77, Mal74, dP79]. L [Boo79, Tsa77]. L_{1} [DW79, EK76]. L_{2}(7) [Joh75a]. L_{2}(s) [Rag73b]. M [Car78, Col77, Jur77, Mar76, YM79, SKD73, Sus74]. m, n [Moo74]. mgf [Per78a]. MLR [GH77a]. μ [Kag73a]. μ_{l} = 0 [Shi76c]. N [Jos75b, PF74]. ν [SW78]. p_{1}^{n_{1}}p_{2}^{n_{2}} ⋯ p_{k}^{n_{k}} [ST75a]. ϕ [FM74]. R [Jur77, Joh74]. s [Law75]. S_{1}S_{2}^{−1} [Pil75]. Σ_{1} = Σ_{2} [Sub75]. t [BV73, JS78, LG77, Rob76, Rob82, Wij75a]. t, χ^{2} [AHS78]. θ > 0 [Bic73b]. θ ≤ 0 [Bic73b]. U [BK78b, CJ78, GS73, Gre77, Joh75b, MWS77, Sen74a]. V_{NN} [Maa73]. ε [Lem74]. W^{1} [And73a].

-Class [SKD73]. -Estimates [Car78, Jur77]. -Estimators [Col77, Mar76, YM79]. -Mixing [FM74]. -Optimal [Lem74, Atw73, Tsa77]. -Optimum [Paz74]. -Ordered [Law75]. -Patterns [Moo74]. -Resolvability [Sha78]. -Resolvable [Kag73a]. -Sample [KR77]. -Spacings [dP79]. -SPRT’S [Lor76]. -Squares [Fed77, HRS75]. -Statistic [Joh75b, MWS77]. -Statistics [Goo77, TG74, Boo79, BK78b, CJ78, GS73, Gre77, Sen74a]. -Test [LG77, Wij75a]. -Tests [JS78]. -Variate [Sus74]. -Way [PF74].

1977 [Efr79].

42 [Mar74a].

80th [LL74].

A. [DOG77]. Absolute [AHS78]. Absolutely [GC78]. Absorbing [Koh74]. abstract [Tuk58]. Acknowledgement [Ber73a, Mar74a, Moh77]. Acknowledgment [Fab74a, GH79, Mar74b]. Action [Ing79]. Adaptation [Alb79]. Adaptive [Ber74a, Ber78a, FR73, KG73, LR79, Mac74b, Sto75, Ton78, Wei78]. Addendum [DW78]. Additional [Le 74]. Additions [TW79]. Additive [HS78]. Adequacy [TA75]. Admissibility [Ber76a, Blu74, BF74b, BF74a, BCS79b, Bro79, CS73a, ES78, GM77, HS78, Jos75c, Jos79b, Kie76, Mak77, Por75, Ram73, SS74, Sco75]. Admissible [Ala73a, Ber76b, Bro73b, CS73b, Ing79, OS79, ST75b, Str78, Var78]. Affine [AB76, HT75a, Kag73a, Sha78]. After [SS78]. Against [BH79, Goo75, Ray74, TH73]. Agreeing [Aum76]. Algorithm [Lju78]. Algorithmic [Wu78]. Algorithms [Kus74, WK74, WW78, Yak74]. Aliasing [HRF74]. Aligned [Sen77b]. Allocation [PT77]. Almost [Car78, DHT73, Pol78, Sen74a]. Alternative [Bre74, Rag73a, Ram73, Sch79a]. Alternatives [BN75, Hus77, HK75, Joh75b, Par79, Pir74, Pir75, Ray74, Sug73, TH73]. Among [Goo64, Goo76b]. Analogues [Bic73a]. Analysis [Alb76, And79, Ber73b, BKR79, Bha76, Dyk79, Fer73, Hab76, Lai75a, LS77, LS79a, MBH75, Mil77b, PF74, Rob78b, Sen77a, Sed77]. Anderson [Fab74a, Fab74b]. Ann [Mar74a]. Antipodally [Bin74]. any [WB74]. Application [Anb76, Arn73, Dun79, FH79, Goo76a, Mac74b, RS77a, SY79, SG73, Sha76, Sho73]. Applications [Ant74b, Arn76, Ber76e, BCS76, Bro79, Efr75, FM77, FH77b, Goo73, HW78, HPS77, Hol73c, Hol79, Hud78, JM77, Lai75a, LS77, LS79a, Mac78, MN79, Mak73, OQ73, SST79, Sen77b, Wij77, Wij79, Woo78, dWV73, dP79]. Approach [Arn79, FK75, GR77, Hab75b, Hil75, San75, Sen79b, Ste79, VS77, Win73, WB74]. Approaches [Wie76]. Approximate [Gol75a, Kie74, Kre79, MI77, Que49, Sha75, Spr76b]. Approximately [SY78]. Approximating [Anb77, GP73, GH78]. Approximation [Abd73, Anb73, And73c, Dut73, Fab73a, HM76, KG73, Kus74, LR79, Lju78, Rob79c, Rup79, Uos74]. Approximations [CR78, For79, JM75, PS75b, Sen77b, Sha76, Woo77]. Arbitrary [Ber76b, BBB^{+}77, BF79, Dos76, Enn70, Enn73]. Areas [And73c, GH78]. Arguments [OW75]. Arising [Sen79c]. ARMA [Han81, Dav77b, Han75, PD78, RC79, Sha76]. Armed [BF79, Jos75a, Jos85, Kel74]. Array [Ple76]. Arrays [RS74a, Shi76c]. Aspects [Ber73b, Wu78]. Associate [Kag74]. Association [Jog77, SKD73]. Assuming [Goo75]. Assumption [Con73]. Asymmetrical [Che78b]. Asymmetry [Col76]. Asymptotic [ABvZ76, ABvZ78, Alb79, And73a, Ant76, Beh76, Ber75a, Ber72, Ber73a, Ber73b, Ber76f, Ber78b, BM73b, Bic73b, BvZ78, BH79, Bro76, CPL76, Col77, CM75, DZ76, DuM73, Eic79, Fed75b, Fle78a, GR75a, GHH76, Haj74, Han76a, Har75, Har78, Hec76, Hol73b, Hsi79b, Jae79, Jir76, JT74, Joh74, JT78, Jur77, Kal74, KS74, KR77, Kra73, Maa73, MJ76, Mic78, Mic79, Mil77b, MC75, Mui78, Pfa73, Pfa74, PR73, Por77a, PV76, Rao78, Rao80, Rie78, Rob78a, Rus76a, Rus82, Ruy74, RvZ78, Sch77, Sen76, Sen79a, Sha74, SH79, Sha73, Shi76a, Ste76, Sti73, Sto74a, Sub75, Sug73, Sza79, Var79b, Var79a, Wei74, YM77, YM79, dWV73, dP79]. Asymptotically [Ade75, And73b, AL75, Ber74a, Fab73a, Fab78, Fen76, GK76, GO78, MJ76, Mic78, PS73, Sac75, Str78, Ton73, Tus77, Yoh73]. Asymptotics [Hub73]. Attainment [Jos76b, Wij73]. Autocorrelation [BH79, MVK78, Ram74]. Autocorrelations [Men77]. Autocovariances [Han74]. Autoregression [SST79]. Autoregressions [Pag78]. Autoregressive [Alb78b, And75a, And77, Ber74b, Bha78, HK74a, HF79, Pag74, Sha75, YM77]. Average [And75a, And77, Men77, Sha75]. Averaging [Fly74].

B [DOG77]. Bahadur [BB78, Dut73, Fu73, Haj74, Hsi79a, HK75, JS78, Kre79, Spr76b, Wie76]. Balanced [AB76, CS77, HT75a, Joh76b, Kag73a, Kie75, Mos75, RS74a, RF73, Shi76c, Shi76d]. Bandit [Jos75a, Jos85, Kel74]. Bandits [BF79]. Bands [Bre78, GF79, Sen73]. Based [AG79, AG80, FP79, FM75, Joh74, Lai74, Lai77, M.75, MWS77, Men77, Mie79, PK79, Per78b, PS75a, RSL75, SKD73, Sch77, Sen79a, Spr76b, SS73, SV78b]. Bayes [Ber76a, Ber76c, BS78a, BC79, BCS79a, DS73, DZ76, EM76b, Fox78, GHH76, GH77a, GK76, Gol75a, Gol77, Goo75, JT74, JT78, Kel77, KH76, LT73, Lin75, MB79, O’B76, Ruk78, Sch73c, Sin76b, Sin79, Sni77, Str73a, SV78a, VS77, Wag73, Win73, Yoh73, vH76]. Bayesian [Alv77, Ant74b, Bic73b, Bre78, Fer73, FP79, For79, GR77, Gol75b, Hsu79, Moo74, Rub78, Sha74, SV78b]. Behavior [Bro76, GR75b, JT74, JT78, Kou77, MI77, SBS73, YM77, YM79]. Behaviour [Sen74a]. Behrens [Rob82, Rob76]. Below [Anb77]. Bernoulli [Jos85, BF79, Cab77, Dev76, Jos75a, Kel74, Klo73]. Bernstein [HM76]. Berry [Jos85, CJ78, Jos75a]. Best [BK78a, Ber76d, BY77, Bur73, GR77, Jos79a, Por75, Ste69, Ste73, Sin79].

Beta [Aro41, Aro59, TW75, TW79]. Between [BK78a, Gre74b, HPL74, IT78, Kel77, Meh75, SST79]. Bias [Que56, Tuk58]. Biased [Wei78]. BIB [FH79, FH77b, HJ74, HL79]. BIBDs [Kag73b]. Binary [Goo73, Mar74b, Mar74a]. Bingham [JM79]. Binomial [BY77, BC79, Hol73b, OS79, SS74]. Biometric [Yan78]. Birth [Kei75, Kei78]. Birth-and-Death [Kei75, Kei78]. Birthday [LL74]. Bishop [Hab76]. Bivariate [AHS78, AL77, Blo77, Div79, KP77, SH79, TDT76, Whi76]. Black [BHS74, Klu77]. Block [AB76, Ham74, HT75a, Jac78, Joh76b, Kag73a, Kag76, Kie75, Mos75, Nig70, Nig76, Sha78]. Blocks [CS77, FH77b, FH79, Pir74, Pir75]. Bootstrap [Efr79]. Borel [Bla76, Bon76a]. Both [Spi79]. Bound [BK78a, Fuj76, HI75, IT78, Jog78, Jos76b, RW78b, Wij73]. Boundaries [Ruy75]. Boundary [Bro73b, Fab74a, Fab74b, Kao78, Rao78, Rao80]. Bounded [Eat74, Per78a, Wij75a, Wij77]. Boundedly [Hoe77]. Boundedness [Lai75c, RS77b]. Bounds [AG79, AG80, Bje77, Boh73, BS78b, Col77, Joh75b, KS76b, MB79, Mey77a, SG76a, SG77, Ski76, Wel77a]. Bradley [Daw77]. Branching [Dio75, DE79]. Broad [EO78]. Broken [Goe75]. Brownian [Mac74a]. Buehler [Fra77]. Bundle [Sen73].

C [Cra78, DOG77]. Calculation [Div79]. Calibration [Sch73b]. Canonical [LaM77]. Cantelli [Wel77b, Wel78a]. Capacities [HS73, HS74, Rie77]. Carlo [Hub73, Sie76a]. Carver [Cra78]. Case [ARS76, Blu74, CS75, Dok74, Fab73a, Fed75b, Ing79, KZ74, KKMP79, Lin75, Mey77b, O’B76, Ork74, PR75, Per78a, PR73, RvZ78, Sha79b, Sho73, Str73a, TH78, Var79a, vZ76]. Cauchy [RW74a]. Causal [Rub78]. Caused [Str74a]. CDF [GR75a]. Cell [BKR79, CS75, Hab77a, Klo78, Ruy75, Spr76a]. Cells [SU74]. Censored [BJM77, FP79, MJ76, Sen76, Sen79c, Shi75, SV78a, SV78b]. Censorship [BC74a]. Center [RSL75, SN73]. Centering [Hoe73]. Central [BN75, Dun79, Jur73, Mor75, TW75, Wel78b]. Certain [Arn73, Che78b, HPL74, Kag74, Koz78, Lai75c, Le 79, PS75b, Sch79a, SBS73, Var79b]. Chain [Bai75]. Chains [Goo73, Lem74]. Change [Mac74a, Mac78, SS75]. Character [Hol73b]. Characteristic [Ahs77, FM77, Hok75, PD78, Pil75, Ste69, Ste73]. Characterization [AM75, Blo77, BL77, EH74, GH77b, GH79, Hec76, Jam79, KMR79, MVK78, Ram74, Rao76a, RHP74, DOG77]. Characterizations [AS79, Fab73c, Fab81, Gor73, TA75]. Characterizes [Fin78]. Characterizing [Sam75]. Charts [Lai74]. Chernoff [FM74, Lai75b, Spr76a]. Chi [CS75, MS75, OQ73, Ruy75, Spr76b, Spr76a]. Chi-Square [CS75, OQ73, Ruy75, Spr76b, Spr76a]. Chi-Squared [MS75]. Choice [Bro73a, Muc73, Ram75]. Circle [KMR79]. Circulant [GP73]. Circular [Epl79]. Class [Ber75c, BCS76, Eat78, EO78, KS77, KS83, Lin76, SKD73, ST75b, Sni77, TH73, Wij79]. Classes [Hwa73, Kag74, Wah74]. Classical [Fin78]. Classification [And73a, BDS76, FC74, GO78, Gup74, IT78, Sha74, Sha77b]. Classifications [PF74]. Clustering [Har78]. Co [Por75]. Co-Ordinate [Por75]. Coefficient [Alb78a, Hal74, PV76, Str74a, Zid78]. Coefficients [ARS76]. Coherence [HS78]. Coherent [Bue76]. Coin [HC75, Wei78]. Coincide [Wie76]. Coincidence [Adi75]. Collector [Sen79b]. Combinations [Bic73a, Bje77, CH79, Eat74, Hec76, Sch77]. Combining [BB76, BS79b]. Comments [Daw77]. Common [BC74b, CS74, CS76, CS77, NH77, Str73a]. Commutation [MN79]. Compact [M.75, MP73]. Comparing [Ber75d, Pir74, Pir75]. Comparison [Adi74, Beh76, FK75, GD79, HT74, LS79b, Sch74, SG76a, SG77, Sie78, Sin76a, Spr76b]. Comparisons [Lat79, Sch73a]. Competing [Fle78a, Fle78b, Mil77a]. Competitors [Mau75]. Complete [BCS76, Eat78, Fed77, Hab75a, Hoe77, Sni77]. Completely [Gor74a]. Completeness [BJM77]. Component [AL75, BS79a, GHH76, O’B76]. Components [Bro76, OSB76, Sch77, Tho75]. Composite [For79, OQ73]. Composition [Lat79]. Compound [Arn76, Cop74, GHH76, Hor75, Ing79, LR77, Rao76a, Sim76a, Sus74, Var78]. Computation [Paz78]. Concave [HP76]. Concept [Sha77a]. Concerning [Bro75, GG73]. Concomitant [DOY77]. Concomitants [Yan77]. Condition [Wie76]. Conditional [Bon77, Bro78, Eat78, Hol79, Kie76, KN78, Mic79, OQ73, PW75, RGQO79, Rob79a, Rob79b, Sen79a]. Conditionally [Ton77]. Conditioning [Bla79]. Conditions [Bly74, BK77, Fly76, Har75, Kha62, Kha73c, Sha79b, TDT76, Wij77]. Confidence [Bon77, Bre78, BC74b, Bro78, Cau75a, Cau75b, CS73b, CS73a, Fra77, GF79, Goo64, Goo76b, Jos75c, Jos76a, Kie76, Lai76, LJJ73, Mey77b, Sen73, SS73, Wij79, Tuk58]. Configurations [RF73]. Confounding [WB74]. Conic [Bla79]. Conjecture [Hol77, Jos75a, Jos85, LR77]. Conjectures [Hub73, SK74]. Conjugate [DY79]. Connected [EH74, Mak74]. Connectedness [RF75]. Consistency [AT76, AT79, Ber72, Ber73a, Cai75, DS73, DW77, DW78, Hal74, HPW73, HP76, MY77, Ple76, Rei73, RC79, RW75, Rob78b, Sag75, Sil78, Win75]. Consistent [Ano77, Ber74b, BHO^{+}77, Fu75, Sin77, Sto77]. Constant [Ade75]. Constrained [KS77, KS83, Kus74, ST75a]. Constraint [Klo78]. Constraints [GC78]. Constructing [Wij79]. Construction [AB76, Fed77, Ham74, HL79, RS74a, Rag73b, RS74d, SW78, Kie75]. Constructions [Hwa73]. Contained [Le 74]. Containing [GR77]. Contents [Ano76]. Contiguous [BN75]. Contingency [Goo76a, Shi76b]. Continued [Aro41, Aro59, TW79]. Continuous [Con73, Dos76, Jaq75, Lem74, Lin75, Ngu79, Rob75, vH76]. Continuous-Time [Ngu79]. Contractions [Gol77]. Contrast [Pfa73, Pfa74]. Contrasts [BDS76, Goo64, Goo76b]. Contribution [Bro78, Sti76]. Contributions [AH75, HRS75, Kha73d, RS74a, SY79]. Control [sC79, CW80, Dos76, Lai74]. Convergence [BP77, Bha74, BK77, Bra76, CH79, DW79, Dur73, EH77, EK76, FM74, Fu73, Fu75, GS73, HL77, Han74, Han76b, HK74b, Hus77, LeC73, Lin75, Lju78, Mak74, Neu75, Paz74, Per78b, RS75, Rec75, Sen76, Sen79c, Sho73, Sin76b, Str73b, Sus74, Tsa77, Wah75b, Wit75, WW78]. Convergent [Atw76]. Converging [Atw73]. Convex [EH77, Wel78b, Wyn77]. Coordinate [Arn79, Hab75b]. Coordinate-Free [Arn79, Hab75b]. Coordinates [Ber76a, Ber76d, Ber76c]. Correction [ABvZ78, Ano76, AG80, Aro59, Ber73a, BG80, Bro73b, Cau75a, CW80, CS76, Enn73, Fab81, FH79, Goo76b, Han81, HP75, Hin80, HS74, Jos85, Kad73, Kei78, KHS74, Nig76, OD80, Pir75, Pre79, Rao79, Rao80, Sen74c, SG76b, SG77, Ste73, Sti79, Wel78a, Wij75b]. Corrections [Aro59, BR75, Haf81, KS83, Kha73c, Kha73d, Pfa74, Rob82, Ros82, Rus82, TW79]. Correlation [ARS76, Alb78a, Jir76, Kar77, Lig77, PV76, Que49]. Cost [Pat76]. Counterexample [Fly74]. Counting [Aal78b]. Counts [Hab77a]. Coupon [Sen79b]. Covariance [And73b, And75a, Arn73, BBB^{+}77, EP73, EM76b, FR78, Gle79, GG73, Haf79, Haf81, Hol73a, IT78, KS74, LT73, MVK78, MC75, Nag73, Roy73, Sha75, Sha76, She79, Str74a, Sug73, Sza79]. Covariances [Han76a]. Covering [MP73]. Cramér [Bly74, FH77a, Jos76b, Sch77, Wij73]. Criteria [Atw76, CS73a, Fly76, Har78, KS74, Nag73, PS75a, Sug73, WW78]. Criterion [BK76, Hor75, Jaq73, Jaq75, Nor74, O’R75, O’R76, OD80]. Cross [Bon76a]. Cross-Sections [Bon76a]. Crossings [Kao78]. Crowded [SU74]. Cumulants [GCL75]. Cumulative [Per77, Pha73]. Curvature [Efr75]. Curve [GF79, SV78b]. Curvilinear [Bro73a]. Cyclic [Joh73, SKD73].

Darmois [Lor73, MB79]. Data [Ber79a, Epl79, FP79, GR75b, Hab73, Lig77, Rob75, Shi75, Sti77, Wel79]. Death [Kei75, Kei78]. Decision [Ala73b, BR77, Cop74, Enn70, GHH76, GH77a, Hor75, Ing79, Jaq73, Jaq75, KC78, Lem74, Mac74b, Nor74, Rao76a, Rhe74, Sch79b, Sni77, VS77, Var78]. Decreasing [HPS77]. Decrement [Aal78a]. Deficiencies [Alb79]. Defining [Efr75]. Degree [Kha62, Kha73c, Kha73d, KS76a]. Deleted [Wag73]. Delta [WB79]. Densities [Ber75c, Rei73, dMTT75]. Density [Ber79b, BR73, BR75, BS78b, But77, Dav75, Dav77a, DW77, DW78, DW79, Fab73b, Lii78, Mey77a, MY77, Ngu79, Ros75a, Ros82, Sil78, Sin77, Wah75a, Wah75b, Wal77, WB79, WD79, Win75]. Depend [Tor77b]. Dependence [AHS78, Alb78b, Alb78c, Dev76, DH78, KS78, Klo73, KP77, KKMP79, Sha77a]. Dependency [HPL74]. Dependent [AG79, AG80, Dyk79, GR75b, Kou77, MM76, Por77b, Por79]. Derivative [Fab73b, Mal74]. Derivatives [Mac74b, Mey77a, Sil78, Sin77]. Derived [Eat78, Hab74, HT75a, SH79]. Deriving [Zid78]. Descriptive [BL75a, BL75b, BL76]. Design [AH75, Atw76, BH79, sC79, CW80, Fed77, Gil77, HRS75, HS79, Kag74, LR79, McD75, RS77a, Sed77, Tsa77, Wah74, Wei78, Wei79, WW78, Yoh73]. Designs [AB76, Arn79, Atw73, Che78b, Che78a, CS77, EH74, FH77b, FH79, Fuj76, HK74a, Ham74, HT75a, HJ74, HF75, HA78, HL79, Hwa73, Jac78, Joh73, Joh75a, Joh76a, Joh76b, Kag73a, Kag76, Kie74, Kie75, KS76a, Mos75, Nig70, Nig76, Paz74, Paz78, Rag73b, RF75, RS74d, SKD73, SW78, SRK76, Sha78, Shi76c, Shi76d, ST76, Wu78]. Detecting [SS75]. Determine [JLL79]. Determining [Bro79, Cov73, GT74b]. Deviation [Ros75a, Ros82, Sie78]. Deviations [BR73, BR75, KHS72, KHS74, Sie75, Sie76b, Sto74b, Woo78]. Diagonal [FR78]. Dichotomies [Cop74]. Differ [Hab75b]. Difference [BK78a, Kel77]. Differences [Bon76b, Rag73b]. Different [CS73a, Gre74b]. Differentiable [BK77, Sen74a]. Differential [Boo79, JLL79, Muc73]. Difficulties [Pie73]. Diffuse [GHH76, LS78]. Diffusions [Bro73b]. Digits [Mar74b, Mar74a]. Dimension [Ros82, Sch78b]. Dimensional [Bla79, BF74b, Ros75a]. Dimensionality [LeC73]. Direct [Hab75a]. Directional [Ber79a, Wel79]. Directions [Pre78, Pre79]. Dirichlet [Fab81, Ant74b, BC79, CH78, Fab73c, Goo75, Goo76a, SU74]. Disagree [Aum76]. Discount [BF79]. Discounted [Dos76, Klu77]. Discounting [Fly74]. Discovering [Sta79]. Discrete [Bro75, Jaq73, Kle75, O’B76, Rob75, SK75, Hab76]. Discreteness [Bla73]. Discriminants [EG74]. Discriminating [IT78]. Discrimination [Gre74b, RW78b]. Discussion [Ano77, BHO^{+}77, Sti76]. Disease [Klo78]. Dispersion [BL76]. Distance [Ber77b, Sus74]. Distinctness [Oka73]. Distinguishability [Kha73a]. Distributed [PR73, Rob79c, Ton77]. Distribution [Ahs77, Ala73a, ABvZ76, ABvZ78, And73a, Ber77a, BvZ78, Bin74, Bla79, Blo77, Boc75, Bre78, Bre74, Che73, Con73, CM75, DOY77, De 74, DuM73, Dur73, EH77, EM76a, Eic79, Epl79, Fab73c, Fab81, Fed75b, Fin78, Fle78a, GR75a, GH77b, GH79, GT74b, Han76a, Hec76, Hil75, Jae79, Jen74, KMR79, KH76, Lau74, Maa73, Mak74, Mal74, MI77, Meh75, OS79, PR75, Par73, Par79, PW75, Pha73, PR73, Pil75, PS75a, PS73, RS75, Rao78, Rao80, RHP74, RW78b, Sax76, SN73, Sti73, SV78a, WW73, Wel77a, Wel78b, Woo74, Yan77, Zid73, dP79, vZ76]. Distribution-Free [ABvZ76, RW78b, Sax76, ABvZ78, BvZ78]. Distributions [AHS78, AM75, Bas77, BV73, BM73c, BS78c, Bro75, CPL76, CS74, CS76, Cra74, EM74, Fer74, GT73, Gol77, Goo76a, Gre74a, GH78, GG73, Har78, HK74b, HM76, Hoe77, HS79, Hsi79b, Jam79, Joh75b, JM79, KS74, Kle75, KP77, Kul73, Lai79, Law75, LG77, Mac74a, MA73, MB79, MO74, MC75, OS79, Pol73, Rag73a, Rao76a, RGQO79, Rus76a, Rus82, Sam75, Sax76, SU74, Sub75, Sug73, Sus74, Sza79, Tho78, TW75, TDT76, Wel77a, Whi76, dWV73]. Divisible [Ham74, Joh76a]. Do [Hab75b, Sti77]. Does [Tor77b]. Domain [Bha78]. Domains [And77]. Dominated [Tor77a]. Dose [GC78]. Dose-Response [GC78]. Double [Uth73]. Down [Sen77b]. Duals [HT75a]. Due [HRF74]. During [Fer76]. Dynamic [AT79, Bla76, Fly74, Fly76, HT75b, Rup79, Sch79b, Uos74].

Early [Sti78]. Economical [DH79]. Edgeworth [BG80, Alb78a, BG78, Bic74, GCL75, Pra76]. Effect [RF73]. Effective [Mau75]. Effects [Har76, Rub78, Sch73a, TH78]. Efficiencies [Ber76f, Ber78b, Hsi79a, JS78, Lai78]. Efficiency [BB78, Efr75, Gor74b, Hol77, Hos79, HK75, Kre79, Mic78, Mic79, Por77a, Pra76, Wie76]. Efficient [Ade75, And73b, Bai75, Ber74a, Ber78a, Fab73a, Fab78, Fen76, GO78, Sac75, Sch74]. Efron [Daw77]. Eigenvalues [Oka73]. Elementary [KHS72, KHS74, Saw73, SK74]. Elimination [Che78a, RF75]. Ellipsoidal [Ber79b]. Empiric [GR75a, Rec75]. Empirical [BC79, DZ76, Dok74, EH77, EM76b, Eic79, FM74, FM77, Fox78, GH77a, Jae79, KH76, Lin75, Neu75, O’B76, RS75, Sho73, Sie76b, Sin76b, Sin79, Sni77, Sto74b, SV78a, VS77, Wel77a, Win73, Wit75, vH76, vZ76]. Ended [Lor73]. Environments [DE79]. Equal [CS75, Pol73]. Equations [GT74a, Hab73, Hab74, Muc73]. Equi [Wah75a]. Equi-Spaced [Wah75a]. Equiprobability [Goo75]. Equivalence [Fly76, Kal74, Kie74, KR77, Sha73, Wah74]. Equivariant [BZ74]. Erdos [Cso79]. Error [Bje77, Dav75, Dav77a, Dut73, Fox78, IT78, Lau74, Pie79, Spi79, Str74b]. Errors [Gup74, Kou77, Str74a]. Esseen [CJ78]. Estimability [AS79, BK76]. Estimable [BDS76]. Estimate [BM73a, DuM73, GF79, Lii78, MB78, Sil78]. Estimated [Dur73]. Estimates [AT76, AT79, Ber74a, Ber77b, Ber77c, Ber74b, BR73, BR75, BC74a, Cai75, Car78, Dav75, Dav77a, DW77, DW78, DW79, Gol75b, Hab75b, Hab77b, Jur77, LeC73, Mee76, Mil77b, Ree78, Rei73, Ros75a, Ros82, Sch78a, Sie78, SS73, Str78, Wag73, YM77]. Estimating [Ber77a, CS74, CS76, GT74a, Hoy75, Jos79a, Kim74, NH77, Sch78b, Spi79, Zid73]. Estimation [Aal78a, Ade74, AL77, Ala79, Alv77, Anb73, And73b, And75a, And77, Bai75, Ber75c, Ber76b, BB76, BBB^{+}77, BC79, Blu76, Bon75, BS78b, BS78c, BC74b, Cab77, CH78, CD74, CS73b, CS73a, Col76, Cra74, Dav77b, Dev76, Dio75, DE79, Dry76, Dun79, EM76b, Fab73b, Fab78, Fen76, FP79, Fle78a, Fle78b, Fox78, FC74, FR78, GK76, Gle79, GT73, GC78, Haf79, Haf81, Hal74, Han75, Han81, Har76, HF79, Hos79, Hud78, Jos75b, Kei75, Kei78, Kel77, KN78, KH76, Lin75, Mak74, Mal74, Men77, Mey77a, Ngu79, OS79, OSB76, Pag74, Pat76, PK77, Pha73, PD78, Ple76, Por77b, Por79, Pra76, Que56, RSL75, Rao76b, Rao79, RS74b, RW74b, Roy73, SY78, Sag75, Sag78, SS78, Sär76, SG76a, SG77, Sha79a, Sim76a, Sin76b, Sin79, Sta79, Ste79, Str74b, Sus74]. Estimation [SV78a, TH78, Ton73, Var79a, Wah75a, Wah75b, Wal77, WB79, WW73, Woo74, Woo77, Yak79, Yan78, Yoh73, Yoh74, dMTT75]. Estimator [Ant74a, Ber78a, Ber76d, BK78b, But77, Kha73b, MWS77, PS74, Por75, Ram73, Rao78, Rao80, SN73, SZ76, SV78b]. Estimators [Ade75, Ala73a, BK78a, Bar73, Ber76a, Ber76c, BS78a, Boc75, Bre74, BZ74, Bro73b, Bro76, Bro79, Col77, EM76a, Fu73, Fu75, GR75b, GM77, Gol77, HI75, Jos77, Jos79b, JM79, Kou77, LT73, Mar76, Meh75, Mic78, MY77, MB78, Obe78, Per78b, Pfa73, Pfa74, Pfa79, Por77a, RC79, Ruk78, Sac75, Sch74, ST75b, Sen79a, Sha73, Sin77, Sti77, Sto74a, Sto75, Str73a, Tak79, Var79b, WD79, Win75, YM79, Zeh75, Zid78]. Euclidean [BK78a]. Even [HF75, Shi76c]. Events [DHT73]. Exact [ARS76, Hsi79a, Koz78, Par79, PS75a]. Example [Fra77]. Examples [Sch73c]. Excess [Kao78]. Exchangeable [GT73, Sha77a, Tho78]. Exclusion [MM76]. Existence [Fed77, HP74, HP75]. Exit [Lai75a]. Expansion [Alb78a, And73a, BG78, BG80, Maa73, PV76, Rob78a]. Expansions [ABvZ76, ABvZ78, Bic74, BvZ78, Car78, CPL76, Cra74, GCL75, KS74, MC75, Pfa73, Pfa74]. Expectation [DS73, Saw73]. Expectations [AG79, AG80]. Expected [Ber79d, DOY77, Fis75, Hab77a, Lai73, Lor76, PS75b, RS74c, SK74, Sha79b]. Experimental [Che78b, Paz74]. Experiments [Atw73, Gil77, GD79, HT74, HS79, Joh73, Le 79, Nig70, Nig76, Tor77a, Wei78, WB74, Yoh73]. Exponential [Ahs77, Ber79a, Ber72, Ber73a, Bic73b, Blo77, BS79a, Bre74, CM75, DY79, Efr78, EM74, GM77, Gol77, Hab77b, Hip74, Hud78, JT78, JLL79, Lai75c, Mic79, O’B76, RW78a, RS77b, Sam75, Sha73, Sin76b, Sin79, Str73b, Uth73, Zid73, vH76]. Exponentially [Per78a, Wij75a, Wij77]. Extended [HS78]. Extension [Har76, Szy73]. Extensions [BK76, KG73, Per77]. Extraction [Pie79]. Extrapolation [O’R75, O’R76, OD80]. Extrema [BP73]. Extreme [Kul73, III75]. Extremum [SBS73].

F. [Pra76]. Factor [Goo75, SRK76, WB74]. Factorial [Fuj76, Hab75a, Joh73, RF73, Shi76c, Shi76d, WB74]. Factorials [ST75a]. Factorization [FM75]. Factors [RS75]. Failure [AL77]. Families [Bic73b, Bon75, Cai75, DY79, EM76a, Efr78, Gre74a, Hip74, Hoe77, HS79, Hud78, JT78, Lan74, Lor73, Mic79, O’B76, RW78a, RS74d, Sin76b, Sin79, Str73b]. Family [Aal78b, Ala73a, GM77, GH77a, Gol77, Jam79, JLL79, Sam75, Sha73, Sus74, Tak79, vH76]. Favorable [Rie77]. Feature [GPW75]. Ferguson [Bla73, BM73c]. Fewer [Mau75]. Fiducial [Cau75a, Cau75b]. Fienberg [Hab76]. Filaments [Sen73]. Finding [Arn79, ST76]. Finite [Cai75, HC75, Joh78, Jos75c, Jos77, Jos79a, Jos79b, Lai79, Per78a, Rob78a, RW78b, Sär76, Sco75, ST75b, Sha79b, Var78, Wyn77, Yak74]. Finitely [HS78, Lor77]. First [JLL79, Jos75b, Lai75a, Men77, Ski76]. Fisher [Rob82, Hub74, JM79, OW75, PS74, Pra76, Rob76, Sav76, Sti76]. Fisherian [Bas77]. Fit [CS75, Gre77, Maa73, MS75, OQ73, RGQO79, Spr76b, Ste76, Wei74, dWV73, dP79]. Fitting [Bha78, HRF74, SST79]. Fixed [BCS79b, MBH75, Mey77b, Pat76, Sen73]. Forecasts [Pie75]. Form [HT75b, LaM77, Oka73]. Formal [BG78, BG80]. Forms [Ber75b, GP73, dWV73]. Formula [Goo77, MM76]. Four [PS75a]. Fractional [Fuj76, Shi76c, Shi76d, SRP76]. Fractions [Aro41, Aro59, TW79]. Free [ABvZ76, Arn79, Hab75b, PS73, RW78b, Sax76, ABvZ78, BvZ78]. Frequencies [Klo78]. Frequency [And77, Hab73, Hab74, Hab77a, Hab77c, Pie73]. Friedman [Jen74, Sch79a]. Function [Anb77, Aro41, Aro59, BM73a, Ber77a, BR73, BR75, Bre78, CPL76, Con73, De 74, Dur73, EH77, FM77, Goo73, Jae79, KH76, Mak74, Mal74, Meh75, MY77, Pha73, PR73, PV76, Ram74, Roy73, SV78a, Tor77b, TW79, Wel77a, Wes79, Yan78, dMTT75, vZ76]. Functionals [Kim74, Mac74a]. Functions [BK77, Div79, EP74, Fuj73, GC78, HPS77, Hoy75, Jos76a, KP77, KKMP79, LJJ73, LR77, Lin76, MR75, Mey77a, NPS77, PK77, PK79, Pie75, PS77, RS75, RW74a, Sam75, Saw73, SBS73, Sen74a, SK74, SH79, Sho73, Sti74, Sti79, TA75, TG74, TW75, WK74, Wel77b, Wel77c, Wel78a, Yak79]. Further [Daw77, HRS75, Kha73d, Por79].

Gambling [Bla76]. Games [Koh74, KZ74, Lin74, Mau75, Sen75, Zam73]. Gamma [Fin78]. Gauss [Dry76, Hab75b, Har76]. Gaussian [Deo73, GP73, Gre74b]. General [Adi78, AS79, And73c, Fuj73, Hal74, Hil75, Kar78, Kie74, LaM77, McD75, MS75, Ork74, Per78a, Str74a, Wij77, WB74, WW78, Yan77]. Generalizations [FR79, Uos74]. Generalized [Arn76, Arn77, BM73a, Ber76a, Ber76c, BS78a, DS73, EP73, EGS76, FK75, Kag74, Kie75, LT73, RW75, RS74d, SW78, SZ76, TG74, Wei79]. Generating [Goo73, Hoy75, Rus76b, Sam75]. Generators [BS79b]. Geometric [SW78]. Geometrical [Ham74]. Geometry [Efr78, HT75a]. Given [Vit79, Whi76, Wij79]. Glivenko [Wel77b, Wel78a]. Global [BR73, BR75, Lii78]. Good [Obe78, LR77]. Goodness [CS75, Maa73, OQ73, RGQO79, Spr76b, Ste76, dP79]. Goodness-of-Fit [OQ73, RGQO79, Spr76b, Ste76, dP79]. Group [Arn77, Ham74, Joh76a]. Groups [Gor74a, Hip75, Rus76b].

H. [Sed77]. Hadamard [HW78]. Hajek [Dup75, Szy73]. Hammersley [Kha73b]. Harold [Smi78]. Harry [Cra78]. Hellinger [Ber77b]. Henry [DL79]. Hermite [Wal77]. Heterogeneity [Che78a, RF75]. Heteroscedasticity [Bic78]. Heuristic [Bro79]. Hiding [MP73]. High [Hos79]. High-Order [Hos79]. Hodges [Ant74a, BK78b]. Holland [Hab76]. Homogeneity [TH73]. Homogeneous [Roy73]. Horvitz [Ram73]. Hotelling [Smi78]. Hsu [And79, ACL79, Chu79, Leh79]. Hyperadmissibility [LR74]. Hypergeometric [Lai79]. Hypotheses [Adi78, CS77, For79, JM75, JT78, Koz78, PS75a, Rus76b, Tho75]. Hypothesis [Hal74, Hsi79a, Hus77, KS76b, McD75, Yak74].

I. [LR77]. I.I.D. [Neu75, RvZ78, Sho73, vZ76]. Identically [PR73, Ton77]. Identification [Mie79, SY79]. Identified [Fed75b]. Identity [Haf79, Haf81, Hud78]. IFRA [BS79a]. II [Ber78b, HA78, LS79a, Le 79, Woo74]. II. [BL75b, NPS77]. III. [BL76]. Imperfect [Sen75]. Implications [CS73a]. Importance [Sie76a]. Imprecise [PK79]. Improved [GH77a, SZ76, Sim76b]. Improvement [Sin77]. Improving [BZ74]. Inadmissibility [Bar73, Ber76d, Ber76c, BCS79b]. Include [Har76]. Inclusion [MM76]. Inclusion-Exclusion [MM76]. Incomplete [AB76, Aro41, Aro59, CS77, Ham74, HT75a, HRF74, Hoe77, Joh76b, Kag73a, KZ74, Mos75, Rhe74, Sha78, TW75, TW79]. Incompletely [MBH75]. Inconsistencies [GT74b]. Incorrect [FS77]. Independence [Jam79, Kha62, Kha73c, Kha73d, Mil77a, Ros75a, Ros82, Ruy74, Shi75, Shi76b]. Independent [Bar73, BB76, Bon76b, CH79, DHT73, Mar74b, OS79, PR73, Rec75, Ton77, Mar74a]. Index [Shi76c]. Indirect [Hab74, Hab77c]. Induced [BR77, Bha74]. Inequalities [Bly74, Gup74, Jog77, Joh76a, Kar74, Law75, Ser74]. Inequality [Eat74, FH77a, Kag73b, RS77a, Ste69, Ste73, Szy73]. Inference [Aal78b, DS73, Dok74, FM75, Hil75, JM77, Joh74, Klo73, LS78, Leh79, III75, Pie73, Rub78]. Inferences [Wes79]. Infinite [Shi76b, YM77]. Infinitely [Zam73]. Information [Bas77, Ber79d, BC74b, GD79, Hok75, Hub74, KZ74, Le 74, Paz78, PS74, Rhe74, Sen75]. Innovation [Dav77b]. Insoluble [Bro73b]. Integer [Kha73b, WB74]. Integral [CPL76, Mak74, OQ73, Ple76, RGQO79]. Integrated [Dav77a]. Interblock [BC74b]. Interpolating [Wah75a]. Interpolation [Hub74]. Interpretations [WK74]. Intersection [Lau74]. Interval [CS73b, Dav77b, GK76, RS74b, Ski76, TH78, Woo77, Zeh75]. Intervals [Blu74, Goo64, Goo76b, Jos75c, Jos76a, LJJ73, Sax76, SS73]. Introduction [BL75a]. Invariance [Bha76, LR73, Sen74b, Sen74c, Sen77a, Sen79b, SRP76]. Invariant [And75b, Ber76d, Blu74, EP74, Lai75c, M.75, OSB76, Per78a, Por75, Pre78, Pre79, Rob79c, RS77b, Sch73c, Uth73]. Invariants [Bon76a]. Inverse [BY77, Haf79, Haf81, PT74, PT77, Sha75]. Inverses [Sha76]. Involving [BF74a, Hab77c]. Irrationality [Cov73, HC75]. Isomorphic [Joh75a]. Isotonic [RW73, RW75]. Isotonized [RW74a]. Iterated [Mak73, Nor74, SG73, Wel77c]. Iterative [HT75b].

J [Hip75, LR77]. J. [LL74]. Jackknife [Efr79, Mil74a, Mil74b]. Jackknifing [Ree78, Sen77a]. James [MB78]. Jaroslav [Dup75]. Joint [Goo73, Har75, Jen74, Jos79b, MC75]. Jointly [GP73].

Kagan [DOG77]. Kaplan [GF79]. Kendall [Alb78a, Jir76, PV76]. Kernel [DW79, Sil78, Wah75b]. Kernels [Kim74, Wah74]. Kesten [KG73]. Kiefer [Bro78, Lor76, SS79]. Knot [Wah75b]. Knots [Wah75a]. Known [Blu76, Gil77, Jos75b, Sin77, Ski76]. Knut [HF75]. Kolmogoroff [Ste73, Ste69]. Kolmogoroff-Type [Ste73, Ste69]. Kolmogorov [Rag73a, RSL75]. Konijn [Sed77]. Koopman [Lor73, MB79].

Lag [Rob79c]. Large [Ade74, BC74a, Gre77, Gre74b, KS76a, KHS72, KHS74, MI77, MS75, Rob73, Sie75, Sie76b, Sie78, Sto74b, Str73b, SV78b, Wel77b, Wel78a, Woo78, Tuk58]. Large-Sample [Ade74, MS75, Rob73, Str73b]. Largest [ARS76, Blu76, Ton73]. Latent [CPL76, MC75, Mui78]. Law [Fen76, SG73, Wel77c]. Laws [Cso79, Mak73, Wel77b, Wel78a]. Layouts [Tho75]. Least [AT76, AT79, BK78a, Hab75b, KN78, Rie77, YM77].

Least-Squares [YM77]. Lebesgue [Sin76b, Sin79]. Lebesgue-Exponential [Sin79]. Lecture [Efr79]. Lehmann [Ant74a, BK78b, Spr76a]. Lemma [HS73, HS74]. Length [WW78]. Lengths [Goo73]. Levels [LS79b, WB74]. Life [BS79a, BC74a, Sen76]. Lifetimes [Mil77a]. Likelihood [And75a, Bar73, BCS76, Cai75, Che73, CS75, DuM73, Fed75a, FK75, FS77, FM75, GC78, Hab73, Hab74, Hab77b, HL77, Hin79, Hin80, Hsi79b, JM79, Kei75, Kei78, KS74, LR77, Mil77b, PR73, Pra76, Ree78, Rei73, RC79, RW74b, RW78a, Sen76, Sen79a, Sim76a, Sto75, Sug73, Sza79, WW73, Woo74, Woo78, dMTT75]. Limit [BN75, BC74a, Deo73, Dun79, Hol73c, Hol79, Jur73, Lai78, Mor75, PV76]. Limiting [BV73, Hol77, Rag73a, SBS73]. Limits [Cau75a, Cau75b]. Line [Spi79]. Linear [Adi78, AS79, And73b, And75a, AT76, AG79, AG80, BK78a, BP77, Bha78, Bic73a, BH79, Bje77, Bra76, CH79, Dry76, Eat74, EG74, EK76, FR78, GM77, GPW75, Hab73, Hab74, Hab75a, Hab77a, Hal74, HT74, Hec76, Hoe73, Hos79, Hsi79a, Hus77, HK75, JM75, Jos76a, Jur77, Kao78, Klo78, KR77, Kre79, LaM77, LJJ73, Mac74b, McD75, Mee76, PT74, PT77, Per78a, PS73, Rao76b, Rao79, Rob79c, RvZ78, SY78, Sch77, Sta79, Sti74, Sti79, Str74a, Wel77a, Win73, Yoh74, YM79]. Linearity [Ant76]. Linearized [Ant74a]. Lines [Adi75, Pot74]. Linnik [DOG77]. Lives [BS79a]. Local [Ber75a, Kre79, PV76, RW78b, Sug73]. Locally [Ber75e, Shi75]. Location [Ber74a, Ber77c, Ber78a, Ber75c, Ber76a, Ber76d, Ber76c, Ber76e, BL75b, BH79, Blu74, Bon75, BS78c, Bre74, BF74b, BF74a, Col76, Col77, DH79, Fen76, Gho73, MWS77, Mar76, PS74, Por75, RS74b, Rob79b, Sac75, SS78, Sch74, SG74, SG76b, Sie78, Sto74a, Sto75, Uth73, Zeh75, Zid73]. Log [Fed75a, Hab73, Hab74, Hab77a]. Log-Linear [Hab73, Hab74, Hab77a]. Logarithm [Fab73b, Mak73, SG73, Wel77c]. Logistic [MA73, Wes79]. Look [Efr79]. Loss [BM73a, Ber76b, BBB^{+}77, Ber79c, BS78c, Fox78, Str74b, Sus74, Tor77b]. Lower [BS78b, HI75, Jos76b, Kle75, Sen75, Wes79, Wij73]. Lu [ACL79].

M [DOG77, Hab76]. M.L.E. [SS74]. MA [PD78]. Main [RF73]. Majorization [MO74, NPS77]. Manifolds [M.75]. Mann [Epl79]. Manova [Fuj73, TW75, EP74, Kar78, Wij79]. Many [Lor77, SK75, Zam73]. Marginal [MC75]. Marginals [Vit79, Whi76]. Markov [Har76, Bai75, Dos76, Dry76, Fle78b, FR73, Goo73, Hab75b, Jaq73, Jaq75, Lem74, Ngu79, Yak79]. Markovian [Rhe74]. Marshall [Jog78]. Martingale [Sen79b]. Martingales [Ste69, Ste73]. Math [Mar74a]. Mathematical [DOG77, Sti78]. Matrices [And73b, And75a, Arn73, CPL76, EP73, EM76b, GG73, HW78, IT78, KS74, MVK78, Paz78, Sha76]. Matrix [BBB^{+}77, CPL76, DH78, FR78, Gle79, Haf79, Haf81, Hok75, JM79, LT73, Mac74b, MN79, MC75, Mui78, Nag73, Saw73, SK74, Sha75, Str74a, Sug73, Sza79]. Maxima [Deo73]. Maximal [Bon76a, Lem74]. Maximising [Enn73]. Maximization [CPL76, Sha79b]. Maximizing [Enn70]. Maximum [And75a, Bar73, Cai75, DuM73, FK75, GC78, Hab74, Hab77b, JM79, Kei75, Kei78, Lai73, MB79, Mil77b, Pra76, Ree78, Rei73, RC79, RW74b, Sen79a, Sha77b, Sim76a, Sto75, Szy73, WW73, Woo74, dMTT75]. Maximum-Likelihood [DuM73]. Maxmin [Ray74]. Mean [Ade74, Ala73a, Alb78c, BM73a, Ber76b, BB76, BBB^{+}77, Blu76, Boc75, Bon76b, BC74b, CS74, CS76, CS77, Cov73, Dav75, Dav77a, Dut73, EM76a, Gle79, Jos75c, Jos79a, Kha73b, LJJ73, LT73, Mee76, Meh75, NH77, PK77, Per78b, RW74a, SS75, Sti73, Str73a, Sza79, Ton73, Zid78]. Mean-Estimation [PK77]. Mean-Square [Dut73]. Means [Cau75a, Cau75b, IT78, Jos79b, Kel77, Rob76, Rob82, Sha79a, Sho74]. Measurability [Rei73]. Measurable [BP73, Dar74]. Measure [HRF74, Lii78, Ros75a, Ros82]. Measured [Ber79c]. Measurements [HA78]. Measures [Arn79, BR73, BR75, Fer74, GD79, Lan74, Sie76b, Sto74b]. Median [Meh75, Mic78, PR75, Pfa79, RW73, Sch78a]. Medical [Wei79]. Meier [GF79]. Membrane [BKR79]. Memory [HC75, Yak74]. Method [AB76, And73c, Bro79, GT74b, Ham74, HL79, HT75b, KG73, Rag73b, ST75a]. Methods [Anb73, Efr79, KR77, Wah75a, Wah75b, dMTT75]. Metric [MP73]. Milliken [BK76]. Minimal [BJM77, HP74, HP75, Hol77, Lan74]. Minimax [Ala73a, BM73a, Ber75c, Ber76b, BBB^{+}77, Ber79c, Boc75, BS78c, EM76a, Gle79, Hol73b, HS73, HS74, Jos75b, LT73, Nor74, OS79, Pha73, Ste79, Str73a, Str74b, Tak79, Zeh75]. Minimaxity [Ber76e]. Minimize [Lai73]. Minimizing [GPW75, Lor76]. Minimum [Ber77b, Bon75, Pfa73, Pfa74, Por77a, Sär76]. Minimums [EM74]. Minque [Bro76]. Minque-Type [Bro76]. Misclassification [GPW75]. Mises [GH79, JM79, BK77, GH77b, HM76, Sch77, Sen74a]. Missing [Dry76]. Mixed [Mil77b, WB74]. Mixing [FM74, GR75a, MR75, Wit75]. Mixture [LG77, Nig70, Nig76, Pol78]. Mixtures [Ant74b, BC79, EO78, Goo76a, Haf79, Haf81, Tor77a, TW75]. MLE [Ber73a, Ber72]. Mode [Sag75, Sag78]. Model [AS79, Bha78, Bic73a, Dav77b, Dev76, FS77, FR78, Hal74, HRF74, Jur77, Kou77, LaM77, McD75, Men77, Mil77b, PD78, Rao76b, Rao79, Rie78, Sch78b, Str74a, TH78, Yoh74, YM79]. Models [Aal78a, And77, AT79, And75b, AL75, Ber74a, Ber77b, Ber79a, Ber72, Ber73a, BL75a, BL75b, BL76, BDS76, CM75, Dok74, Dry76, Dun79, Fuj73, Hab73, Hab74, Hab75a, Hab77a, Hab77b, Hab77c, Han75, Han81, Hok75, LS78, Lau74, MBH75, Mil77a, Pag74, Per78a, Rhe74, Rob73, SY78, Sha76, Tho75]. Modified [Anb77, HT75b, Lor76]. Moment [Kad71, Kad73, Sam75, Wij77]. Moments [AM75, HI75, Hir73, JLL79, Jos75b, Lai75c, Ski76]. Monotone [BCS76, CC76, Che73, GH77a, KS78, She79, Wel78b, vH76]. Monotonic [GC78, HPW73, KP77, KKMP79]. Monotonicity [BCS79a, EP74, Fuj73, Wij60, Wij75b]. Monro [Anb77]. Monte [Hub73, Sie76a]. Morimoto [Bro75]. Most [Ber75e, DH79, Mau75, MJ76, Shi75, Uth73]. Motion [Mac74a]. Move [Lin74]. Moving [And75a, And77, Men77, Sha75]. Much [Hab75b]. Multi [Che78a]. Multi-Way [Che78a]. Multinomial [Ala79, Goo75, Goo64, Goo76b, Klo78, LR77, Mor75, OS79]. Multinormal [Zid78]. Multiparameter [Hud78, JT78, JLL79]. Multiple [Aal78a, Ala73b, ARS76, Bar73, Bro73a, FC74, GH77a, JM77, McD75, Pag78, PS73, RW73, Tak79, TG73, VS77, Win73, Yak74]. Multistage [AH75, LS79b]. Multivariate [Ala73a, And79, Bas77, Ber79b, Ber76b, BS78a, Boc75, DJ76, Dry76, Dyk79, EM76a, EM76b, EM74, FR79, Hab76, Hsi79b, Hsi79a, IT78, KKMP79, Koz78, Lig77, LT73, MA73, Mar76, MO74, MM76, McD75, Mil77a, Oka73, PS75a, RGQO79, RS77a, Rus76a, Rus76b, Rus82, RvZ78, Sag78, Sie75, Str73a, Sza79].

Natural [Hud78]. Near [Sin79]. Nearest [DW77, DW78, Enn70, Enn73, MY77]. Nearly [Lor77]. Nearly-Optimal [Lor77]. Necessary [Bly74, Har75, TDT76]. Neighbor [DW77, DW78, Hwa73, MY77]. Neyman [HS73, HS74, LL74]. Noise [BKR79, Dun79, Pag74]. Non [Ber75d, EP73, KS74, Neu75, Pot74, RvZ78, Sho73, Sug73, TH73, vZ76]. Non-I.I.D. [Neu75, RvZ78, Sho73, vZ76]. Non-Null [KS74, Sug73]. Non-Parametric [Pot74, TH73]. Non-Sequential [Ber75d]. Non-Singularity [EP73]. Noncentral [Ber75b, Lau74, SK74, Str73b, TW75]. Noncontinuous [Sin76b]. Nonexistence [HF75]. Nonfull [O’R76, OD80]. Nonhomogeneous [Fle78b]. Noninvertible [Pie75]. Nonlinear [Dok74, Kim74, LS77, LS79a, PW73]. Nonlinearity [Bic78]. Nonnull [Sza79]. Nonorthogonal [PF74]. Nonparametric [Aal78a, Aal78b, AL77, Ano77, Ant74b, Bic74, BL75a, BL75b, BL76, BHO^{+}77, BS78b, Fer73, FP79, Fle78b, Gho73, Gol75a, Gol75b, Lai78, Par79, Ruy74, Sag75, SS78, SN73, SY79, Sin77, Sto77, SV78b, Win75, Yak79, dMTT75]. Nonstationary [Deo73, Pie79]. Norm [BK78a, RW74a]. Normal [AHS78, Ala73a, Alb78c, AT76, And75b, BM73a, Ber76b, BB76, BBB^{+}77, Bla79, Blu76, Boc75, Can79, CH79, CS74, CS76, Div79, EM76a, EO78, Gle79, GG73, HT74, Hsi79b, JM75, Jos76a, Kha62, Kha73c, Kha73d, LJJ73, LG77, LT73, Mee76, Meh75, NH77, O’R75, O’R76, OD80, RGQO79, RS77a, Sha79a, Str73a, Str74b, Sus74, Sza79, Uth73]. Normal-Theory [JM75]. Normality [Ber72, Ber73a, BM73b, CM75, DuM73, Har75, Hec76, Jir76, Ruy74, RvZ78, SH79, Shi76a]. Normalized [Bha74]. Norms [M.75]. not-quite [Tuk58]. Note [AB76, ABvZ78, Alb78a, Ano76, BY77, BK77, Cai75, Can79, Cau75a, Cob75, Dev76, DW78, Fab74a, Fab74b, FH79, Gil77, Gol75b, Gre74a, HK74b, Hin80, Hip75, HS74, Joh78, Jos77, Kad73, Kal74, Kei78, Kel74, Kha73c, Kha73d, Kul73, Lai75a, LJJ73, Mar74b, Mil77a, Moh77, Par79, PW75, Pre79, PS73, Ros75b, Ruy75, SG77, Sin76a, Sti79]. Notes [Aro59, Ber73a, BR75, Bro73b, CS76, Enn73, Goo76b, HP75, KHS74, Nig76, Pfa74, Pir75, Que56, Sen74c, SG76b, Ste73, Wij75b]. Notion [Zam73]. Nuisance [GT74a, Pfa79]. Null [KS74, Sug73]. Number [BS79b, Enn70, Enn73, Wei74, WB74]. Numbers [Wel77b, Wel78a].

Objective [FM75]. Objects [SK75]. Observation [Hab74, Hab77c, MI77, Ple76]. Observations [Abd73, BJM77, Che73, Dry76, Goe75, Joh74, Le 74, PK79, PT77, SV78a]. Observed [Dun79]. Obtain [OQ73]. Obtained [Shi76c]. Obtaining [Enn70, Enn73]. Occasion [LL74]. Off [HL79]. Olkin [Jog78]. One [ABvZ76, ABvZ78, Alb78b, Alb79, BF79, Bur73, Con73, GM77, Lai77, LG77, Por75, RS74c, SG73, Sen74b, Sen74c, SK75, Szy73, TH78, WBG74, vH76]. One-Armed [BF79]. One-Parameter [vH76]. One-Sample [ABvZ76, ABvZ78, Alb78b, Alb79, LG77, SG73, Sen74b, Sen74c, WBG74]. One-Way [TH78]. Open [Lor73]. Open-Ended [Lor73]. Operator [PD78]. Operators [PW73]. Optimal [Anb73, Arn79, Atw73, sC79, CW80, Che78a, EG74, GK76, Hal77, Jac78, KS77, KS83, KS76a, Lai73, Lem74, Lor77, Muc73, Obe78, PT77, Pfa79, RF73, Ros75b, Shi76d, Str78, ST76, Ton73, Tsa77, Tus77, Var79a, Wah75b, Wu78, Yoh73]. Optimality [Atw76, BH79, Boh73, Che78b, DZ76, EH74, Fly76, Hor75, Hsi79b, Jaq73, Jaq75, MB78, Pol78, Sch79a, SRK76, Sim76b, Var79b]. Optimization [Kus74]. Optimum [Bas77, Enn73, Kie74, Kul73, LS79b, Paz74, Paz78, Ram75, WW78]. Order [AM75, Bha74, Bic73a, Bje77, CH78, DOY77, Dut73, Efr75, Hec76, Hir73, Hok75, Hos79, Kad71, Kad73, MM76, MR75, Men77, III75, RW78a, Rus76a, Rus82, Sch73a, SG73, Sen74b, Sen74c, Sen77b, Sho73, Sti74, Sti79, Wei74, Wel77b, Wel77c, Wel78a, Woo77, Yan77]. Order-Effects [Sch73a]. Order-Statistic [Dut73]. Ordered [HS79, Law75, Par79, Pir74, Pir75, RS74b, RW74b, SS74, Sax76, TG73, TH73]. Ordering [Ton77]. Orders [HF75]. Ordinate [Por75]. Orientations [Pre78, Pre79]. Original [JLL79]. Orthogonal [Cra74, Fed77, Gil77, Kim74, Sch77, Wah75b]. Other [And75a, CS75, TG74]. Outlier [Gre74a]. Outlier-Prone [Gre74a]. Overlook [Hal76, Hal77].

P [Hab76]. Paired [FK75, Lat79, Sch73a, Sin76a]. Pairing [Che73, Goe75]. Pairs [Rie77]. Pao [ACL79]. Pao-Lu [ACL79]. Paper [Daw77, Hip75]. Paradox [Mak77]. Parallel [Pot74]. Parameter [BC79, Blu74, Bon75, Bre74, BF74b, BF74a, Col76, Dun79, FM75, Gho73, GM77, GT74a, HC75, Lor77, Mac74a, Par73, PS74, Rao78, Rao80, Sac75, SS78, Sto74a, Sto75, Tor77b, WW73, Woo74, Zid73, vH76]. Parameters [Adi74, And75a, BS78c, CS73b, DH79, Dev76, Dur73, JLL79, Jos76a, Kag73b, Pfa79, PD78, Rao76b, Rao79, RS74b, Rob79b, SS74, Ste76, Zeh75]. Parametric [Ber77b, Pot74, TH73]. Pareto [RHP74]. Partial [Aal78a, Mac78, Man77, MVK78, Ram74, Szy73]. Partitioning [Can79]. Partitions [Gil77, TG73]. Patchwork [Kie75]. Paths [Bha74]. Patterned [Arn73, Sza79]. Patterns [Moo74]. PBIB [Kag74]. Pearson [HS73, HS74]. Penalty [dMTT75]. Percentile [CC76]. Performance [FS77, RW78b]. Periodic [CD74, Pag78]. Periodogram [SST79]. Permutation [OW75, Rob73, SH79, Wel79]. Permuted [Mak73]. Perturbations [RS75].

Pfanzagl [Hip75]. Pitman [Lai78, PS74, Wie76]. Plan [Joh78]. Plans [Kar74, Mau75, RF73, WB74, Wyn77]. Plots [Dok74]. Plus [Pag74]. Point [BC74b, CS73b, Fu73, Fu75, HI75, Woo77]. Points [CV76, Ple76]. Poisson [Sha79a, Sim76a, Var79a]. Policies [FR73, Hal77, Lem74, Ork74]. Pólya [BM73c, Wei79]. Polykays [TG73]. Polynomial [Hok75, KS76a, Kim74, Mac78, PD78]. Polynomials [Kha62, Kha73c, Kha73d, TDT76, Zid78]. Population [Alb78c, BY77, GR77, Goo75, Joh78, Jos75c, Jos79a, Rob78a, Sco75, Str74b, Wyn77]. Populations [Can79, Gho73, Goo64, Goo76b, Gor73, IT78, Jos77, Jos79b, Lai79, NH77, Sär76, ST75b]. Positive [AHS78, DH78, Sha77a, WB74]. Possible [Sin79, Ste69, Ste73]. Posterior [CM75, FM75, Fra77, Hal74, HM76, Kle75, MI77, Sha77b]. Power [ABvZ76, ABvZ78, Ber75a, BvZ78, EP74, Fuj73, KS76b, Lai77, MBH75, Par73, RS74c, Rob73, SW78]. Power-One [Lai77]. Powerful [Ber75e, MJ76, Shi75, Uth73]. Powers [Str74b]. Practice [Hab76]. Prediction [Bha78, Bro73a, Lin74, PK77, PK79, TA75, Tor77b]. Predictive [EG74, Hin79, Hin80, LS78]. Preferences [Bue76]. Preliminary [SS78]. Presence [Col76, GT74a, Kra73, Pfa79]. Present [Beh76]. Preservation [PS77]. Prime [SW78]. Principle [Bha76, Bon77, Sen74b, Sen74c]. Principles [Sen77a, Sen79b]. Prior [CH78, Fer74, GT73, Goo75, GT74b, Rao76a, Tho78]. Priority [Ber73a, Fab74a, GH79, Mar74b, Moh77, Mar74a]. Priors [BR77, DY79, GHH76, HS78]. Probabilities [Aal78a, Ala79, Bai75, Ber75b, Hal76, Hal77, Kle75, Lat79, RS77a, Sie75, Sie76b]. Probability [BHS74, Bur73, Chu79, Div79, Dok74, Eat74, EGS76, Enn70, Enn73, Fer74, Fra77, Goo73, GH78, Gup74, GPW75, Jog77, Jog78, Jos75b, KHS72, KHS74, Lai75c, OQ73, RGQO79, Ser74, Sha77b, Sim76b, Ski76, Sta79, Sto74b, Wal77, WB79, WD79, Wes79, Wij60, Wij75b, dMTT75]. Problem [ABvZ76, ABvZ78, Arn76, Beh76, BvZ78, sC79, CW80, Efr75, Fle78b, GHH76, GH77a, GO78, Ing79, Kad71, Kad73, Kar78, Kel74, Kre79, Lor76, Mac74b, PT74, Rao76a, RS77a, Rob76, Rob82, Sen79b, Shi76a, Sus74, Sza79, Yah79]. Problems [Ant74b, Arn73, Bar73, BR77, BB76, Ber76e, BS78a, Bro73b, BF74b, BF74a, DOG77, DE79, Fed75b, Fer73, Fox78, Gol75a, HPS77, Hor75, KS77, KC78, KS83, Kus74, Lin75, Muc73, O’B76, PS75a, San75, Sni77, Tak79, Ton78, VS77, Var78]. Procedure [Anb77, DZ76, Hsu79, Rup79, ST76, Ton73]. Procedures [Anb73, Arn77, Arn79, BY77, BM73b, BF74b, Bro78, Car77, DH79, Fab74a, Fab74b, Gho73, GK76, Joh74, JT78, Kie76, LS79b, MBH75, Mie79, Rob79a, Rob79b, Sch73c]. Process [CR78, Dio75, Enn70, FM74, Jur73, Kei75, Kei78, Neu75, Ngu79, Rao78, Rao80, Rec75, Roy73, Sen77b, Sim76a, SS79]. Processes [Aal78b, And75a, Ant74b, BC79, Bla76, Cai75, Deo73, DE79, Dos76, Eic79, Fle78b, GR75a, Gre74b, HK74a, HL77, HF79, Hos79, Jaq73, Jaq75, KN78, MR75, RC79, Sen77b, Sen79c, Sha75, Sho73, Uos74, Wit75, YM77]. Product [BC74a, Hab77c, Lan74]. Products [Arn73, Arn76, Hab75a, Tor77a, TG73].

Programming [Bla76, Fly74, Fly76, HT75b, Sch79b]. Programs [FR73]. Progressively [Sen76, Sen79c]. Prone [Gre74a]. Proof [Fin78, LR77, Moh77, Ram73, SK74]. Proper [Jac78, Str73a]. Properties [Dav75, Dav77a, EGS76, Gor73, GG73, Jac78, Kha73b, Mac78, MN79, Mil77b, MY77, Rob76, Rob79a, Rob79b, Rob82, Sch79a, Sch77, Sen79a, Sto74a, Wah75b, Wal77, vZ76]. Property [Ahs77, Boh73, EP74, Kar77, Mee76, RW74a, Ste69, Ste73, Wij60, Wij75b]. Proportions [GHH76]. Pseudorandom [BS79b].

Quadratic [BM73a, Ber75a, Ber75b, Ber76b, BBB^{+}77, BS78c, GP73, Oka73, OSB76, Ros75a, Ros82, dWV73]. Quantile [CR78, Sen79c, Sho73]. Quantiles [De 74, Kul73]. quite [Tuk58].

R [Sti76]. R. [Pra76, Sav76]. Radhakrishna [DOG77]. Radial [Lau74]. Radii [Mey77b]. Ramachandran [DOG77]. Random [Ade74, BM73a, BC74a, CH79, Cov73, DE79, Eat74, Enn70, Enn73, Goe75, Haf79, Haf81, Hal76, Hal77, Har76, Jos75b, Lau74, Mak73, Mar74a, Mar74b, Neu75, RS75, Rec75, RW74b, Ruy75, Sha77a, Sho74, TH78, Ton77]. Randomization [Rob73, Rub78, ST75a, SRP76, Ste79]. Randomized [Pir74, Pir75]. Range [Cau75a, Cau75b]. Rank [Adi74, Adi78, Alb78a, Alb78b, Alb79, Ant76, Beh76, Ber74a, Ber75a, BP77, Bra76, CH78, Con73, DOY77, EK76, Hoe73, Hus77, HK75, Jir76, Jur73, KR77, Kre79, Lai75b, O’R76, OD80, Pir74, Pir75, PV76, Rey75, Rus76a, Rus82, RvZ78, SG73, Sen74b, Sen74c, SG74, SG76b, Sen77b, Shi75, Shi76a, WBG74]. Rank-Order [Sen74b, Sen74c]. Ranking [FK75, HPS77, KR77, San75, Ton78]. Rankings [Sin76a]. Ranks [Adi75, Haj74]. Rao [DOG77, Bly74, FH77a, Jos76b, Wij73]. Rapid [Klu77]. Rate [AL77, Fu73, Fu75, Hus77, Mak74, Per78b, Sin79, Sus74, Win75]. Rates [BS78b, EK76, GH77a, GS73, Lin75, Sin76b, Sin79]. Ratio [BCS76, Che73, CS75, EGS76, Fed75a, FS77, Hsi79b, KS74, Lai75c, RW78a, Sen76, Sim76b, Sug73, Sza79, Wij60, Wij75b, Woo78]. Rational [Nor74]. Ratios [HL77]. Re [Goe75]. Re-Pairing [Goe75]. Real [Enn70, Enn73, Sti77]. Reconstructions [Moo74]. Recovery [BC74b]. Recurrent [Bro73b]. Recursions [Han76b]. Recursive [Fab78, WD79, WK74]. Red [BHS74, Klu77]. Red-and-Black [BHS74, Klu77]. Reduced [BK78b, Neu75, Sho73]. Reducing [RW74a]. Reduction [Kag74, Le 79]. Regarding [Jos75a, Jos85, MB79]. Regenerative [Goo73]. Regression [Ade75, Adi74, Adi75, Adi78, Anb77, AT76, Ano77, Bar73, Beh76, Bha76, BHO^{+}77, BH79, Bro73a, CC76, sC79, CW80, Fed75b, Fed75a, FR78, Gor73, HK74a, HPW73, HP76, Hub73, Jur77, KS76a, Kou77, Mac78, Mak73, Mak74, O’R75, O’R76, OD80, PT74, PT77, Ple76, Pot74, PS73, RW73, RW75, SS78, Sch78a, SY79, She79, Sto77, Tak79, Vit79, Wah74, Win73, Zid78]. Regressions [FC74, Rob75]. Regret [Nor74]. Regular [Atw76, WW78].

Regularly [De 74]. Related [GS73, HL77, Mac74a, Pfa73, Pfa74, Wel77a]. Relating [Sen77a]. Relation [RF73, SST79]. Relations [JLL79, Jur77]. Relationships [Meh75]. Relatively [Wei74]. Remainder [BP77]. Remarks [Dar74, Por79, WD79]. Remez [ST76]. Renewal [FR73, LS77, LS79a]. Rényi [Cso79, Szy73]. Repeated [Arn79, FH77b, FH79, HA78, Koh74, KZ74, Man77]. Replacement [Cob75, Kal74, Man77, Ser74]. Replication [SRP76]. Reply [Ano77]. Representation [Str78, TDT76]. Rereading [Sav76, Sti76]. Residuals [Ade75, And75a, Bic78, Mac78]. Resistance [Mos75]. Resistant [HJ74]. Resolution [Fuj76, Shi76c, Shi76d]. Resolvability [Kag76, Sha78]. Resolvable [AB76, Kag73a]. Response [GC78, Hab77b]. Restricted [Alb79, San75]. Restrictions [LeC73, RW78a]. Result [Can79, Moh77]. Results [Ber76a, Ber76d, Ber76c, Fle78a, JM77, Joh74, Mui78, O’B76, Sha74, Ste76]. Review [DOG77, Hab76, Mil74a, Mui78, Sed77]. Rewards [Dos76, Lem74]. Ridge [Obe78]. Ridits [BL77]. Riemannian [M.75]. Rietz [Efr79]. Right [SV78a]. Risk [JT74, MB78, TA75, Wag73, Zid78]. Risks [Fle78a, Fle78b, Mil77a]. Robbins [Anb76, Anb77]. Robust [AL75, Ber77c, Ber78a, Col76, DH79, Dav77b, GR75b, Hub73, KE77, Kou77, Mar76, Por77b, Por79, Rie78, Sti77, Yoh74]. Robustly [Bic78]. Robustness [BH79, HPL74, Joh76b, Kar77, PS75a]. Role [Rub78, Sen76, Sen77a]. Roots [CPL76, DH78, HF79, MC75, Mui78, Pil75]. Rule [Ala73b, Lai75c]. Rules [Cop74, Ing79, Lat79, Pol78, RW78b]. Run [Fer76, Goo73]. Run-Lengths [Goo73]. Runs [Ros75b]. rv [Wit75].

S [Hab76, Sed77]. Sample [Ade74, ARS76, ABvZ76, ABvZ78, Alb78b, Alb79, Bha74, BvZ78, BC74a, BCS79b, Con73, De 74, Dok74, Dur73, EP73, Epl79, FM74, FR79, Goe75, Gre77, Gre74b, Jos79b, Kao78, KR77, Kre79, Kul73, Lai73, Lai77, LG77, Lor76, MB79, MJ76, Men77, Mey77b, Moh77, MS75, O’B76, Oka73, Per78b, PS75b, RS74c, Rob73, RW78b, Sed77, SBS73, SG73, Sen74b, Sen74c, Shi76a, Str73b, SV78b, WBG74]. Samples [AG79, AG80, Con73, HPL74, Rob78a, SV78b, Tuk58]. Sampling [BY77, Cob75, DuM73, GT73, Hol73c, Kal74, Kar74, LS78, LR74, LG77, Man77, Pat76, Ram75, Sco75, Ser74, Sie76a, Ste79, Tho78]. Saturated [RF73]. Savage [FM74, Lai75b]. Scale [Bre74, BF74a, CS73b, EO78, HK75, RS75, RS74b, Rob79b, Uth73, Zid73]. Scaled [Mey77b]. Scatter [Mar76]. Scheffe [Boh73, DL79]. Scheme [SKD73]. Schemes [BM73c, Pat76]. Schur [NPS77, PS77]. Score [Adi74]. Screening [Klo78, KKMP79, Yah79]. Search [Hal76, Hal77, SK75]. Second [Efr75, Hok75, Kha62, Kha73c, Kha73d, Woo77]. Second-Order [Hok75]. Sections [Bla79, Bon76a]. Seeger [Can79]. Segmented [Fed75b, Fed75a]. Selecting [BY77, GR77]. Selection [ARS76, Ber79c, DH79, Gho73, GPW75, Hal74, HS79, Mie79, RS74b, San75, Spr76a, Ton78]. Selections [Bla73, BP73, Dar74]. Semicontinuous [Sen75]. Semiregular [Joh76a].

Sense [Haj74]. Separating [Gor74a]. Sequence [Ade74, But77, Sac75, Sus74, Var78]. Sequence-Compound [Sus74]. Sequences [Atw73, Atw76, BF79, FM74, Lai76, Mac78, Sha79a, WB79]. Sequential [Alv77, Ber73b, Ber75d, Ber75e, Ber76f, Ber78b, BB78, BM73b, Bic73b, Blu76, Bra76, BCS79b, BCS79a, Bur73, Cab77, Car77, EGS76, Fab74a, Fab74b, For79, GK76, Hal76, Hal77, Hol73b, Hol77, KS77, KS83, Kel77, Kha73a, Lai73, Lai75b, Lai75a, Lai75c, LS77, Lai78, LS79a, Lai79, Le 79, Lor77, MB79, Pat76, PT74, PS75b, Rey75, SG74, SG76b, Sen77a, Sie76a, Sim76b, SK75, SS73, Ton73, Var79b, Var79a, WBG74, Wei78, Wei79, Wij60, Wij75a, Wij75b, Woo77, Woo78, Yoh73]. Serial [Han76a, Kar77, Lig77]. Series [CV76, Dun79, GP73, Par73, PK77, PK79, Pie79, Rob78b, Wah75b, Wal77, Que49]. Set [Fed77, Jos75c]. Sets [AM75, BDS76, EH77, Mey77b, Wij79, Wyn77]. Setting [BCS79b, SS79]. Several [Adi74, Adi75, Cau75a, Cau75b, FC74, NH77]. Shape [Bic73b, Jam79]. Sharp [Ski76]. Shift [Joh75b, RSL75]. Short [Moh77]. Sided [LJJ73, Ray74]. Siegmund [Anb76]. Sign [Hol73b, Kra73]. Signal [Dun79, Pag74, Pie79]. Signed [Ant76, Rey75]. Signed-Rank [Ant76, Rey75]. Significance [Dyk79, FM75, LS79b]. Similar [BV73]. Simple [Fin78, Hil75, Hoe73, Hus77, JT78, Pot74]. Simplest [Sni77]. Simultaneous [Goo64, Goo76b, O’R76, OD80, RS74b, Wij79]. Singular [Paz78]. Singularity [EP73]. Situation [MJ76]. Situations [Por77b, Por79]. Six [Kul73]. Size [ARS76, Ber79c, BCS79b, Jam79, Lai73, Lor76, MB79, Mey77b, PS75b, RS74c, Sen73]. Sizes [HL79, O’B76]. Slope [Spi79, Spr76b]. Slopes [Koz78]. Small [Epl79, Hab77a, Wei74]. Smallest [Joh74, Wij79]. Smirnov [Moh77, FR79, KS76b, Rag73a, RSL75]. Smooth [Sti74, Sti79]. Sobolev [M.75]. Solution [PT74, Rao76a, Rob76, Rob82, Ton78]. Solutions [Fox78, GHH76, Gol75a, GO78, Var78]. Some [AH75, Bar73, Ber73b, BM73b, Bic73a, BR73, BR75, Daw77, Deo73, EP74, Fer73, Fuj73, Gil77, Gol75a, Gol75b, Han76b, Hoe77, Hol73c, Hwa73, JM77, Kha73b, Mac74a, Mui78, Nag73, O’B76, Per77, Pie73, RS74c, RS74d, SG76a, SG77, Sen77a, Sen77b, Sen79c, Sin77, Sub75, Tak79, Uos74, Wah74, WBG74, WD79, Wu78, vZ76, MN79]. Space [Dos76, MP73]. Spaced [CV76, Wah75a]. Spaces [Fer74]. Spacing [Kul73]. Spacings [dP79]. Spanning [BDS76]. Sparse [SU74]. Special [Kag73b, Mee76, PR75, Rie77]. Species [Sta79]. Specification [Sen79a]. Specified [IT78, MBH75]. Spectra [Gre74b]. Spectral [Ber74b, But77, CD74]. Sphere [Bin74, Roy73]. Spheres [Lau74]. Spherical [KE77, Lau74]. Spherically [BS78c]. Spline [Hub74, Lii78, Wah75a, WK74]. Splines [CD74]. SPRT [Per78a, RS77b]. SPRT’S [Lor76]. Square [CS75, Dav75, Dav77a, Dut73, Hab75b, OQ73, Ruy75, Spr76b, Spr76a]. Squared [Fox78, MS75, Str74b, Sus74]. Squared-Distance [Sus74]. Squares [Alb76, AT76, AT79, BK78a, Fed77, HRS75, KN78, YM77]. Stabilizing [Hol73a]. Stable [DuM73, Fen76, KC78]. Stages [Zam73]. Standardized [Eic79, Jae79]. Starshaped [Sha79a]. State [Dos76, GHH76, Ing79, Var78]. Statement [Sim76b]. States [Koh74, Sti78]. Stationary [Ade74, Ade75, But77, Dun79, GP73, Ngu79, Ork74, Sha76]. Statist [Mar74a]. Statistic [And73a, DOY77, Dut73, Epl79, Har75, Hoe73, Jir76, Joh75b, KS76b, Maa73, MWS77, Par79, Sch77, Spr76a, Str73b, TW75]. Statistical [BK77, BF74a, Dok74, Efr75, Klo73, III75, Rob79a, Rob79b, Sch79b, Sch73b, Sen74a, Sed77]. Statistics [Ant76, AM75, AG79, AG80, BP77, Bha74, BJM77, Bic73a, Bic74, BL75a, BL75b, BL76, BV73, Bje77, Boo79, Bra76, BK78b, CJ78, EK76, Goo77, GS73, Gre77, Hab73, Hec76, Hip74, Hip75, Hir73, Hus77, HK75, Jen74, Jur73, Kad71, Kad73, KR77, Lai75b, Lai78, Lan74, MM76, MR75, Moh77, MS75, NPS77, III75, PS77, Rag73a, RSL75, Rus76a, Rus82, Ruy75, RvZ78, SG73, Sen74a, Sen74b, Sen74c, Sen76, SH79, Shi76a, Sho73, Ste76, Sti74, Sti78, Sti79, Sub75, Sza79, TG74, Wei74, Wel77b, Wel77c, Wel78a, Woo78, Yan77, DOG77]. Steck [Moh77]. Stein [MB78, Per77]. Step [WW78]. Step-Length [WW78]. Stepwise [Hsu79]. Stochastic [Abd73, Anb73, Bla76, Cai75, Fab73a, GR75a, HK74a, KN78, KG73, Kus74, LR79, Lju78, NPS77, Rup79, Sen75, SS79, Uos74, WK74]. Stochastically [RW74b, Sax76]. Stopping [BM73b, Fer76, Lai73, Lai75c, Per78a, Pol78, RS77b, Ros75b, WBG74, Wij75a, Wij77]. Straight [Spi79]. Strategies [KS77, KS83]. Strategy [Enn73, Jos79a, Ram75]. Strategy-1 [Ram75]. Stratified [Tho78]. Strength [Sen73, Shi76c]. Strict [BCS76]. Strong [AT76, AT79, CR78, DW77, DW78, HK74b, Jos75c, Lju78, RC79, Sil78, Wel77b, Wel78a, Win75]. Strongly [Hal77, Mic78]. Structure [And73b, And75a, She79]. Structures [Gil77]. Student [Rob82, BV73, JS78, Rob76]. Studentized [And73a, Bon76b]. Studies [PS75a]. Study [BC74a, Sie76a]. Sub [Adi78]. Sub-Hypotheses [Adi78]. Subfair [Klu77]. Subfield [HP74, HP75]. Subhypotheses [PS73]. Subintervals [Jae79]. Subject [GC78, RS75, Spi79]. Subsample [Har75]. Subsampling [Gor74a, Gor74b]. Subset [Ber79c, GR77, San75]. Substitution [PW75]. Success [Fer76, Ros75b]. Sufficiency [Bro75, Haj74, LR73, MJ76, PW73, TA75, Tor77b, Wei74]. Sufficient [BJM77, Bly74, Hab73, Har75, HP74, HP75, Hip74, Hip75, Lan74]. Sufficiently [Atw76]. Sum [Alb76, Fer76, Ser74]. Sums [Bha74, Kao78, Lai74, Lai77, Mac78, Mor75, Per77, SH79, Szy73]. Support [HL79]. Suprema [Eic79]. Supremum [Jae79]. Sure [Car78, Sen74a]. Survey [LR74, Sed77, Ste79]. Survival [GF79, SV78a, SV78b]. Susceptible [HJ74]. Symmetric [Bin74, BS78c, Cop74, Fen76, Gho73, Goo75, Goo76a, Joh75b, Kar74, KZ74, Saw73, SN73, SK74, TG74]. Symmetrical [Fuj76]. Symmetry [Arn76, Ber79b, Bur73, KE77, KS76b, RSL75, Rey75, Rus76b]. System [SY79]. Systematic [AG79, AG80]. Systems [BS79a, Rob79c].

Table [BC74a]. Tables [Goo76a, Hab74, Hab75a, Hab77a, Hab77c, Shi76b]. Tail [And73c, Ber75b, Ber76e, GH78, Hil75]. Termination [Bur73, Lai75c]. Terms [BP77, TA75]. Test [Bur73, FS77, HC75, HPL74, Kra73, LG77, Lig77, MBH75, Nag73, OW75, Par79, Pot74, Rey75, Ros75a, Ros82, SS78, Sch79a, Str73b, Uth73, Wij60, Wij75a, Wij75b]. Testing [Adi75, Alb78c, Arn77, Ber79b, CS77, JT78, Klo78, KS76b, Rie78, Rus76b, Sen76, Sub75, Sza79, Tho75, Woo78, Yak74]. Tests [Adi78, ABvZ76, ABvZ78, Alb78b, Alb79, AL75, Beh76, Ber75a, Ber75d, Ber75e, Ber76f, Ber78b, Bic73b, BvZ78, Bic78, BCS79b, BCS79a, CS75, Con73, Dyk79, EP74, EGS76, ES78, For79, FR79, Fuj73, Gre77, GG73, Hol73b, Hol77, Hsi79b, Hsi79a, HS73, HS74, JM75, JT74, JS78, Kar77, KE77, Koz78, Kre79, Lai73, Lai75b, Lai75c, Lai77, Lai78, Lai79, Lor73, Lor77, M.75, Mac74a, Mac78, MB79, McD75, MJ76, Mic79, Mie79, MS75, OQ73, PS75a, Pir74, Pir75, PS75b, Pre78, Pre79, PS73, Ray74, RGQO79, RS74c, RW78a, Rob73, Ruy74, Sch77, SG74, SS75, SG76b, Sen79c, Shi75, Shi76b, Sie76a, Sim76b, Spr76b, TH73, Tus77, WBG74, Wei74, Wel79, dWV73, dP79, vH76, Que49]. th [Mal74]. Their [HW78, HPS77, JM77, Mey77a, PS75a, RF73, Sen77a, Sha76, Goo76a, Pie75]. Theorem [Anb76, BN75, Bro75, BCS76, CJ78, CH79, Dun79, FM74, Fin78, Fu73, Jur73, KHS72, KHS74, Le 79, Paz74, Per77, PV76, PS77, SS79, Ton77, Tsa77, Wel77b, Wel78a, Wij77, Har76]. Theorems [Deo73, Eat78, Hol73c, Hol79, Lai78, Mor75, Sni77]. Theory [AH75, Arn73, BH79, Bro78, EH74, Fed75b, FH77b, FH79, GR75a, GT73, Gre77, Hab76, HRS75, Hol73c, JM75, Kie74, LS77, LS79a, MS75, Paz74, Pie73, RS74a, Sch79b, Sch73b, SY79, Sed77, SG76a, Sen76, SG77, SV78b, Tsa77, Wu78, Yan77]. Thompson [Ram73]. Three [Bar73, SRK76, Ski76]. Tied [Sen77b]. Tied-Down [Sen77b]. Ties [Beh76, Kra73]. Time [And77, Bha78, BH79, But77, CV76, Dos76, Dun79, GP73, Jaq73, Jaq75, Kao78, Lem74, Ngu79, PK77, PK79, Pie79, Rob75, Rob78b, Rob79c, Wij75a, Wij77, Que49]. Time-Invariant [Rob79c]. time-series [Que49]. Times [BM73b, Lai75a, Mac74a, Mac78, Per78a, RS77b, WBG74]. Tolerance [Blu74, Sax76]. Tournament [Mau75]. Trade [HL79]. Trade-Off [HL79]. Transfer [Pie75]. Transformation [Abd73, Hip75, OQ73]. Transformations [Hol73a, RGQO79]. Transition [Aal78a, Bai75, Yak79]. Translation [Blu74, CS73b, WW73, Woo74]. Transposition [HPS77]. Trend [Ber75a]. Trials [Cab77, Klo73, Wei79]. Triangular [Fab74a, Fab74b, Ple76]. Trimmed [Sti73]. Trivariate [AHS78]. Truncated [WW73, Woo74]. Truncation [Par73]. Two [Ber76d, BvZ78, Bla79, BF74b, CPL76, CS74, CS76, Con73, Dok74, Fab73c, Fab81, FM74, FR79, GHH76, Gre74b, Hol79, Ing79, JLL79, Jos75a, Jos85, Kel74, KR77, Kre79, LJJ73, Lin74, MB79, MJ76, Moh77, OSB76, PS75a, Pot74, RF75, Ray74, Rob76, Rob82, Ros75a, Ros82, Sch73c, SRK76, Sha73, Shi76a, Tho75, TDT76, Win75]. Two-Action [Ing79]. Two-Armed [Jos75a, Jos85, Kel74]. Two-Dimension [Ros82]. Two-Dimensional [Bla79, BF74b, Ros75a]. Two-Move [Lin74]. Two-Sample [BvZ78, Dok74, FM74, FR79, Kre79, Moh77, Shi76a]. Two-Sided [LJJ73, Ray74]. Two-State [Ing79]. Two-Way [RF75, Tho75]. Type [Bly74, Bro76, Epl79, Kag74, Rag73a, SKD73, SS73, Ste69, Ste73].

UMP [Kra73]. UMVU [Meh75]. Unbalanced [AL75, Mil74b, TH78, Tho75]. Unbiased [BK78a, Mic78, OSB76, Pat76, Pfa79, Por77a, Zid78]. Unbiasedness [CS75]. Unbounded [Fis75]. Unequally [CV76]. Unified [MS75]. Uniform [DW77, DW78, EH77, Han74, Hec76, KMR79, Lai78, Sco75, Sil78]. Uniformity [Car77, M.75, Pre78, Pre79]. Uniformly [Bon75, Sär76, ST75b, Sin77]. Unimodal [BS78c, LR77, Rei73, Wel78b]. Unimodality [DJ76]. Unit [HF79, PD78]. Units [Man77]. Univariate [Div79]. Universal [Ruk78]. Unknown [BB76, BBB^{+}77, BHS74, Bre74, FR78, Gle79, GCL75, LT73, Mac74a, Mac78, Par73, Ste76, Str73a, Tor77b, Zid73]. Unstable [Rao78, Rao80]. Upper [Cau75a, Cau75b, Col77, Fuj76, Kle75, Ski76, Wes79]. Urn [BM73c, Wei79]. Use [Adi75]. Used [Sub75]. Using [Anb73, Bic78, Cra74, Gor73, III75, Str74a, TG73, WB79, Zid78]. Utility [Ber79d, Fis75, Lin76, Sha79b].

V [DOG77]. Validity [BG78, BG80]. Value [Bro73b, DOY77, Kul73, RW74a, Zam73]. Valued [Ade75, AL77]. Values [Har75, Lor77, SK74]. Variable [Cov73, Enn70, Enn73, HL79, Wah75b]. Variables [Bar73, Bro73a, CH79, Eat74, Jam79, Mak73, Mar74b, RS75, Rec75, Sha77a, Spi79, Ton77, Mar74a]. Variance [Alb76, AL75, Blu76, Bon75, Bro76, Dav77b, Dio75, Dyk79, Joh75b, LJJ73, MBH75, Mil77b, OSB76, Por77a, Sär76, SZ76, Str74b, Tho75]. Variances [BB76, Col77, Str73a, YM77]. Variate [Sus74]. Variates [MM76, Mui78, RW74b]. Variational [JM77]. Varying [De 74, O’B76]. Vector [Ade75, AL77, BM73a, Ber76a, Ber76d, Ber76c, Ber76e, BBB^{+}77, Dun79, Gle79, Haj74, Kha62, Kha73c, Por75, Str73a]. Vector-Valued [Ade75, AL77]. Vectors [Ber75c, Kha73d, Neu75]. Version [Ant74a]. Versus [Bic73b, GHH76, Uth73]. Via [BC79, BM73c]. VII [Shi76d]. Vik [HF75]. Villegas [GT74b]. Violations [Pil75, PS75a]. Volume [Ano76]. vs [Fly74].

W [Hab76]. Wald [FR79, Str73b]. Way [Che78a, PF74, RF75, TH78, Tho75]. Weak [BK77, Bra76, Dur73, ES78, FM74, HL77, RS75, Rec75, Sen76, Sen79c, Sil78]. Weight [Sti74, Sti79]. Weighted [Lai74, Sch78a, SH79]. Weiss [Lor76]. Well [Gil77]. Well-Known [Gil77]. Whether [Pot74]. Which [Wie76, DHT73, Fin78, JLL79, Lai73, Wel78b]. White [Pag74]. Whitney [Epl79]. Wide [Ber75c]. Wiener [Sen77b]. Wilcoxon [Ant76, HPL74, Jur73, SS73]. Wilks [TW75]. Win [BHS74]. Wishart [Haf81, DH78, Haf79, Saw73, SK74]. Wishartness [Kha73c, Kha73d, Kha62]. Without [Con73, Kal74, Wij77, Ser74]. Wolfowitz [FR79, SS79]. Work [And79, Chu79, Leh79, Sti77]. Wrong [Str74a].

Y. [Hab76]. Youden [AH75, Kie75, RS74d, SW78]. Yu [DOG77].

Zero [Anb77, PD78]. Zonal [Zid78].

References

Aalen:1978:NEP

[Aal78a] Odd Aalen. Nonparametric estimation of partial transition probabilities in multiple decrement models. Annals of Statistics, 6(3): 534–545, May 1978. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176344198.

Aalen:1978:NIF

[Aal78b] Odd Aalen. Nonparametric inference for a family of counting processes. Annals of Statistics, 6(4):701–726, July 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176344247.

Agrawal:1976:NMC

[AB76] H. L. Agrawal and B. S. Boob. A note on method of construction of affine resolvable balanced incomplete block designs. Annals of Statistics, 4(5):954–955, September 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176343591.

Abdelhamid:1973:TOS

[Abd73] Sami Naguib Abdelhamid. Transformation of observations in stochastic approximation. Annals of Statistics, 1(6):1158–1174, November 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176342564.

Albers:1976:AEP

[ABvZ76] W. Albers, P. J. Bickel, and W. R. van Zwet. Asymptotic expansions for the power of distribution-free tests in the one-sample problem. Annals of Statistics, 4(1):108–156, January 1976. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343350. See

correction [ABvZ78].

Albers:1978:NCA

[ABvZ78] W. Albers, P. J. Bickel, and W. R. van Zwet. Note: Correction to “Asymptotic Expansions for the Power of Distribution-free Tests in the One-Sample Problem”. Annals of Statistics, 6(5):1170–1171, September 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/

1176344321. See [ABvZ76].

Anderson:1979:PLH

[ACL79] T. W. Anderson, K. L. Chung, and E. L. Lehmann. Pao-Lu Hsu 1909–1970. Annals of Statistics, 7(3):467–470, May 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176344666.

Adenstedt:1974:LSE

[Ade74] Rolf K. Adenstedt. On large-sample estimation for the mean of a stationary random sequence. Annals of Statistics, 2(6):1095–1107, November 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176342867.

Adenstedt:1975:AEE

[Ade75] Rolf K. Adenstedt. Asymptotically efficient estimators for a constant regression with vector-valued stationary residuals. Annals of Statistics, 3(5):1109–1121, September 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176343243.

Adichie:1974:RSC

[Adi74] J. N. Adichie. Rank score comparison of several regression parameters. Annals of Statistics, 2(2):396–402, March 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176342676.

Adichie:1975:URT

[Adi75] J. N. Adichie. On the use of ranks for testing the coincidence of several regression lines. Annals of Statistics, 3(2):521–527, March 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343083.

Adichie:1978:RTS

[Adi78] J. N. Adichie. Rank tests of sub-hypotheses in the general linear regression. Annals of Statistics, 6(5):1012–1026, September 1978.

CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176344307.

Arnold:1979:BEL

[AG79] Barry C. Arnold and Richard A. Groeneveld. Bounds on expectations of linear systematic statistics based on dependent samples. Annals of Statistics, 7(1):220–223, January 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176344567. See [AG80].

Arnold:1980:CBE

[AG80] Barry C. Arnold and Richard A. Groeneveld. Correction to “Bounds on Expectations of Linear Systematic Statistics Based on Dependent Samples”. Annals of Statistics, 8(6):1401, November 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345215. See

correction [AG79].

Afsarinejad:1975:SCT

[AH75] K. Afsarinejad and A. Hedayat. Some contributions to the theory of multistage Youden design. Annals of Statistics, 3(3):707– 711, May 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343133.

Ahsanullah:1977:CPE

[Ahs77] M. Ahsanullah. A characteristic property of the exponential distribution. Annals of Statistics, 5(3):580–582, May 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343860.

Abdel-Hameed:1978:PDB

[AHS78] M. Abdel-Hameed and Allan R. Sampson. Positive dependence of the bivariate and trivariate absolute normal, t, χ^{2}, and F distributions. Annals of Statistics, 6(6):1360–1368, November 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176344381.

Arvesen:1975:ART

[AL75] James N. Arvesen and Maxwell W. J. Layard. Asymptotically robust tests in unbalanced variance component models. Annals of Statistics, 3(5):1122–1134, September 1975. CODEN ASTSC7.

ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176343244.

Ahmad:1977:NEV

[AL77] Ibrahim A. Ahmad and Pi-Erh Lin. Nonparametric estimation of a vector-valued bivariate failure rate. Annals of Statistics, 5(5):1027– 1038, September 1977. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176343957.

Alam:1973:FAM

[Ala73a] Khursheed Alam. A family of admissible minimax estimators of the mean of a multivariate normal distribution. Annals of Statistics, 1 (3):517–525, May 1973. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176342417.

Alam:1973:MDR

[Ala73b] Khursheed Alam. On a multiple decision rule. Annals of Statistics, 1 (4):750–755, July 1973. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176342470.

Alam:1979:EMP

[Ala79] Khursheed Alam. Estimation of multinomial probabilities. Annals of Statistics, 7(2):282–283, March 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176344612.

Albert:1976:WSS

[Alb76] Arthur Albert. When is a sum of squares an analysis of variance? Annals of Statistics, 4(4):775–778, July 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343550.

Albers:1978:NEE

[Alb78a] W. Albers. A note on the Edgeworth expansion for the Kendall rank correlation coefficient. Annals of Statistics, 6(4):923–925, July 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176344267.

Albers:1978:OSR

[Alb78b] W. Albers. One-sample rank tests under autoregressive dependence. Annals of Statistics, 6(4):836–845, July 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176344257.

Albers:1978:TMN

[Alb78c] W. Albers. Testing the mean of a normal population under dependence. Annals of Statistics, 6(6):1337–1344, November 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176344378.

Albers:1979:ADO

[Alb79] W. Albers. Asymptotic deficiencies of one-sample rank tests under restricted adaptation. Annals of Statistics, 7(5):944–954, September 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176344780.

Alvo:1977:BSE

[Alv77] Mayer Alvo. Bayesian sequential estimation. Annals of Statistics, 5(5):955–968, September 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176343951.

Arnold:1975:CDS

[AM75] Barry C. Arnold and Glen Meeden. Characterization of distributions by sets of moments of order statistics. Annals of Statistics, 3(3): 754–758, May 1975. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176343141.

Anbar:1973:OEM

[Anb73] Dan Anbar. On optimal estimation methods using stochastic approximation procedures. Annals of Statistics, 1(6):1175–1184, November 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176342565.

Anbar:1976:ATR

[Anb76] Dan Anbar. An application of a theorem of Robbins and Siegmund. Annals of Statistics, 4(5):1018–1021, September 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343602.

Anbar:1977:MRM

[Anb77] Dan Anbar. A modified Robbins–Monro procedure approximating the zero of a regression function from below. Annals of Statistics, 5(1):229–234, January 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176343758.

Anderson:1973:AED

[And73a] T. W. Anderson. An asymptotic expansion of the distribution of the studentized classification statistic W^{1}. Annals of Statistics, 1 (5):964–972, September 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176342518.

Anderson:1973:AEE

[And73b] T. W. Anderson. Asymptotically efficient estimation of covariance matrices with linear structure. Annals of Statistics, 1(1):135–141, January 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1193342389.

Andrews:1973:GMA

[And73c] D. F. Andrews. A general method for the approximation of tail areas. Annals of Statistics, 1(2):367–372, March 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176342376.

Anderson:1975:MLE

[And75a] T. W. Anderson. Maximum likelihood estimation of parameters of autoregressive processes with moving average residuals and other covariance matrices with linear structure. Annals of Statistics, 3 (6):1283–1304, November 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176343285.

Andersson:1975:INM

[And75b] Steen Andersson. Invariant normal models. Annals of Statistics, 3(1):132–154, January 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176343004.

Anderson:1977:EAM

[And77] T. W. Anderson. Estimation for autoregressive moving average models in the time and frequency domains. Annals of Statistics, 5(5):842–865, September 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176343942.

Anderson:1979:HWM

[And79] T. W. Anderson. Hsu’s work in multivariate analysis. Annals of Statistics, 7(3):474–478, May 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176344668.

Anonymous:1976:CNC

[Ano76] Anonymous. Correction note: Correction to “Contents of Volume 2”. Annals of Statistics, 4(2):440, March 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343423.

Anonymous:1977:RDC

[Ano77] Anonymous. Reply to discussion: Consistent nonparametric regression. Annals of Statistics, 5(4):640–645, July 1977. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343888. See

[Sto77].

Antille:1974:LVH

[Ant74a] Andre Antille. A linearized version of the Hodges–Lehmann estimator. Annals of Statistics, 2(6):1308–1313, November 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176342883.

Antoniak:1974:MDP

[Ant74b] Charles E. Antoniak. Mixtures of Dirichlet processes with applications to Bayesian nonparametric problems. Annals of Statistics, 2 (6):1152–1174, November 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176342871.

Antille:1976:ALW

[Ant76] Andre Antille. Asymptotic linearity of Wilcoxon signed-rank statistics. Annals of Statistics, 4(1):175–186, January 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343352.

Arnold:1973:ATP

[Arn73] Steven F. Arnold. Application of the theory of products of problems to certain patterned covariance matrices. Annals of Statistics, 1(4): 682–699, July 1973. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176342463.

Arnold:1976:APG

[Arn76] Steven F. Arnold. Applications of products to the generalized compound symmetry problem. Annals of Statistics, 4(1):227–233, January 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343357.

Arnold:1977:GGT

[Arn77] Steven F. Arnold. Generalized group testing procedures. Annals of Statistics, 5(6):1170–1182, November 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176344002.

Arnold:1979:CFA

[Arn79] Steven F. Arnold. A coordinate-free approach to finding optimal procedures for repeated measures designs. Annals of Statistics, 7 (4):812–822, July 1979. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176344731.

Aroian:1941:CFI

[Aro41] Leo A. Aroian. Continued fractions for the incomplete beta function. Annals of Mathematical Statistics, 12(2):218–223, June 1941. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177731751. See

additions and corrections [Aro59, TW79].

Aroian:1959:CNC

[Aro59] Leo A. Aroian. Correction notes: Corrections to “Continued Fractions for the Incomplete Beta Function”. Annals of Mathematical Statistics, 30(4):1265, December 1959. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://

projecteuclid.org/euclid.aoms/1177706114. See [Aro41].

Alam:1976:SLM

[ARS76] Khursheed Alam, M. Haseeb Rizvi, and Herbert Solomon. Selection of largest multiple correlation coefficients: Exact sample size case. Annals of Statistics, 4(3):614–620, May 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343467.

Alalouf:1979:CEG

[AS79] I. S. Alalouf and G. P. H. Styan. Characterizations of estimability in the general linear model. Annals of Statistics, 7(1):194–200, January 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176344564.

Anderson:1976:SCL

[AT76] T. W. Anderson and John B. Taylor. Strong consistency of least squares estimates in normal linear regression. Annals of Statistics, 4 (4):788–790, July 1976. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176343552.

Anderson:1979:SCL

[AT79] T. W. Anderson and John B. Taylor. Strong consistency of least squares estimates in dynamic models. Annals of Statistics, 7(3): 484–489, May 1979. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176344670.

Atwood:1973:SCO

[Atw73] Corwin L. Atwood. Sequences converging to D-optimal designs of experiments. Annals of Statistics, 1(2):342–352, March 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176342371.

Atwood:1976:CDS

[Atw76] Corwin L. Atwood. Convergent design sequences, for sufficiently regular optimality criteria. Annals of Statistics, 4(6):1124–1138, November 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343647.

Aumann:1976:AD

[Aum76] Robert J. Aumann. Agreeing to disagree. Annals of Statistics, 4 (6):1236–1239, November 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176343654.

Bai:1975:EET

[Bai75] Do Sun Bai. Efficient estimation of transition probabilities in a Markov chain. Annals of Statistics, 3(6):1305–1317, November 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343286.

Baranchik:1973:IML

[Bar73] A. J. Baranchik. Inadmissibility of maximum likelihood estimators in some multiple regression problems with three or more independent variables. Annals of Statistics, 1(2):312–321, March 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176342368.

Basar:1977:OFI

[Bas77] Tamer Basar. Optimum Fisherian information for multivariate distributions. Annals of Statistics, 5(6):1240–1244, November 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176344009.

Berger:1976:CIN

[BB76] James O. Berger and M. E. Bock. Combining independent normal mean estimation problems with unknown variances. Annals of Statistics, 4(3):642–648, May 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176343472.

Berk:1978:SBE

[BB78] Robert H. Berk and L. D. Brown. Sequential Bahadur efficiency. Annals of Statistics, 6(3):567–581, May 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176344201.

Berger:1977:MEN

[BBB^{+}77] J. Berger, M. E. Bock, L. D. Brown, G. Casella, and L. Gleser. Minimax estimation of a normal mean vector for arbitrary quadratic loss and unknown covariance matrix. Annals of Statistics, 5(4):763– 771, July 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343898.

Breslow:1974:LSS

[BC74a] N. Breslow and J. Crowley. A large sample study of the life table and product limit estimates under random censorship. Annals of Statistics, 2(3):437–453, May 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176342705.

Brown:1974:PCE

[BC74b] L. D. Brown and Arthur Cohen. Point and confidence estimation of a common mean and recovery of interblock information. Annals of Statistics, 2(5):963–976, September 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176342817.

Berry:1979:EBE

[BC79] Donald A. Berry and Ronald Christensen. Empirical Bayes estimation of a binomial parameter via mixtures of Dirichlet processes. Annals of Statistics, 7(3):558–568, May 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176344677.

Brown:1976:CCT

[BCS76] L. D. Brown, Arthur Cohen, and W. E. Strawderman. A complete class theorem for strict monotone likelihood ratio with applications. Annals of Statistics, 4(4):712–722, July 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343543.

Brown:1979:MBS

[BCS79a] L. D. Brown, Arthur Cohen, and W. E. Strawderman. Monotonicity of Bayes sequential tests. Annals of Statistics, 7(6):1222–1230, November 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176344841.

Brown:1979:AIF

[BCS79b] L. D. Brown, Arthur Cohen, and W. E. Strawderman. On the admissibility or inadmissibility of fixed sample size tests in a sequential setting. Annals of Statistics, 7(3):569–578, May 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176344678.

Birkes:1976:SSE

[BDS76] David Birkes, Yadolah Dodge, and Justus Seely. Spanning sets for estimable contrasts in classification models. Annals of Statistics, 4(1):86–107, January 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176343349.

Behnen:1976:ACR

[Beh76] Konrad Behnen. Asymptotic comparison of rank tests for the regression problem when ties are present. Annals of Statistics, 4(1): 157–174, January 1976. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176343351.

Berk:1972:CAN

[Ber72] Robert H. Berk. Consistency and asymptotic normality of MLE’s for exponential models. Annals of Mathematical Statistics, 43(1):193– 204, February 1972. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.

aoms/1177692713. See acknowledgement of priority and correction [Ber73a].

Berk:1973:NAP

[Ber73a] Robert H. Berk. Notes: Acknowledgement of priority and correction to “Consistency and Asymptotic Normality of MLE’s for Exponential Models”. Annals of Statistics, 1(3):593, May 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176342432. See

[Ber72].

Berk:1973:SAA

[Ber73b] Robert H. Berk. Some asymptotic aspects of sequential analysis. Annals of Statistics, 1(6):1126–1138, November 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176342561.

Beran:1974:AEA

[Ber74a] Rudolf Beran. Asymptotically efficient adaptive rank estimates in location models. Annals of Statistics, 2(1):63–74, January 1974.

CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176342613.

Berk:1974:CAS

[Ber74b] Kenneth N. Berk. Consistent autoregressive spectral estimates. Annals of Statistics, 2(3):489–502, May 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176342709.

Beran:1975:LAP

[Ber75a] Rudolf Beran. Local asymptotic power of quadratic rank tests for trend. Annals of Statistics, 3(2):401–412, March 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343065.

Beran:1975:TPN

[Ber75b] Rudolf Beran. Tail probabilities of noncentral quadratic forms. Annals of Statistics, 3(4):969–974, July 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176343199.

Berger:1975:MEL

[Ber75c] James Berger. Minimax estimation of location vectors for a wide class of densities. Annals of Statistics, 3(6):1318–1328, November 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343287.

Berk:1975:CSN

[Ber75d] Robert H. Berk. Comparing sequential and non-sequential tests. Annals of Statistics, 3(4):991–998, July 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176343202.

Berk:1975:LMP

[Ber75e] Robert H. Berk. Locally most powerful sequential tests. Annals of Statistics, 3(2):373–381, March 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176343063.

Berger:1976:ARG

[Ber76a] James O. Berger. Admissibility results for generalized Bayes estimators of coordinates of a location vector. Annals of Statistics, 4(2):

334–356, March 1976. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176343410.

Berger:1976:AME

[Ber76b] James O. Berger. Admissible minimax estimation of a multivariate normal mean with arbitrary quadratic loss. Annals of Statistics, 4(1):223–226, January 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176343356.

Berger:1976:IRG

[Ber76c] James O. Berger. Inadmissibility results for generalized Bayes estimators of coordinates of a location vector. Annals of Statistics, 4(2): 302–333, March 1976. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176343409.

Berger:1976:IRB

[Ber76d] James O. Berger. Inadmissibility results for the best invariant estimator of two coordinates of a location vector. Annals of Statistics, 4 (6):1065–1076, November 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176343642.

Berger:1976:TML

[Ber76e] James O. Berger. Tail minimaxity in location vector problems and its applications. Annals of Statistics, 4(1):33–50, January 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343346.

Berk:1976:AES

[Ber76f] Robert H. Berk. Asymptotic efficiencies of sequential tests. Annals of Statistics, 4(5):891–911, September 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176343587.

Beran:1977:EDF

[Ber77a] Rudolf Beran. Estimating a distribution function. Annals of Statistics, 5(2):400–404, March 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176343806.

Beran:1977:MHD

[Ber77b] Rudolf Beran. Minimum Hellinger distance estimates for parametric models. Annals of Statistics, 5(3):445–463, May 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343842.

Beran:1977:RLE

[Ber77c] Rudolf Beran. Robust location estimates. Annals of Statistics, 5 (3):431–444, May 1977. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176343841.

Beran:1978:ERA

[Ber78a] Rudolf Beran. An efficient and robust adaptive estimator of location. Annals of Statistics, 6(2):292–313, March 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176344125.

Berk:1978:AES

[Ber78b] Robert H. Berk. Asymptotic efficiencies of sequential tests II. Annals of Statistics, 6(4):813–819, July 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176344254.

Beran:1979:EMD

[Ber79a] Rudolf Beran. Exponential models for directional data. Annals of Statistics, 7(6):1162–1178, November 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176344838.

Beran:1979:TES

[Ber79b] Rudolf Beran. Testing for ellipsoidal symmetry of a multivariate density. Annals of Statistics, 7(1):150–162, January 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176344561.

Berger:1979:MSS

[Ber79c] Roger L. Berger. Minimax subset selection for loss measured by subset size. Annals of Statistics, 7(6):1333–1338, November 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176344851.

Bernardo:1979:EIE

[Ber79d] Jose M. Bernardo. Expected information as expected utility. Annals of Statistics, 7(3):686–690, May 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176344689.

Brown:1974:ASP

[BF74a] Lawrence D. Brown and Martin Fox. Admissibility in statistical problems involving a location or scale parameter. Annals of Statistics, 2(4):807–814, July 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176342768.

Brown:1974:APT

[BF74b] Lawrence D. Brown and Martin Fox. Admissibility of procedures in two-dimensional location parameter problems. Annals of Statistics, 2(2):248–266, March 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176342661.

Berry:1979:BOA

[BF79] Donald A. Berry and Bert Fristedt. Bernoulli one-armed bandits — arbitrary discount sequences. Annals of Statistics, 7(5):1086– 1105, September 1979. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176344792.

Bhattacharya:1978:VFE

[BG78] R. N. Bhattacharya and J. K. Ghosh. On the validity of the formal Edgeworth expansion. Annals of Statistics, 6(2):434–451, March 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/

1176344134. See correction [BG80].

Bhattacharya:1980:CVF

[BG80] R. N. Bhattacharya and J. K. Ghosh. Correction to “On the Validity of the Formal Edgeworth Expansion”. Annals of Statistics, 8(6): 1399, November 1980. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345213. See [BG78].

Bickel:1979:RDA

[BH79] P. J. Bickel and Agnes M. Herzberg. Robustness of design against autocorrelation in time i: Asymptotic theory, optimality for location and linear regression. Annals of Statistics, 7(1):77–95, January 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176344556.

Bhattacharya:1974:CSP

[Bha74] P. K. Bhattacharya. Convergence of sample paths of normalized sums of induced order statistics. Annals of Statistics, 2(5):1034– 1039, September 1974. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176342823.

Bhattacharya:1976:IPR

[Bha76] P. K. Bhattacharya. An invariance principle in regression analysis. Annals of Statistics, 4(3):621–624, May 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176343468.

Bhansali:1978:LPA

[Bha78] R. J. Bhansali. Linear prediction by autoregressive model fitting in the time domain. Annals of Statistics, 6(1):224–231, January 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176344081.

Bickel:1977:DCN

[BHO^{+}77] Peter J. Bickel, Frank Hampel, Richard A. Olshen, Emanuel Parzen, M. Rosenblatt, Jerome Sacks, Grace Wahba, Leo Breiman, David R. Brillinger, H. D. Brunk, Donald A. Pierce, Herman Chernoff, Thomas M. Cover, D. R. Cox, and William F. Eddy. Discussion: Consistent nonparametric regression. Annals of Statistics, 5(4):620– 640, July 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/

1176343887. See [Sto77].

Berry:1974:RBU

[BHS74] Donald A. Berry, David C. Heath, and William D. Sudderth. Redand-black with unknown win probability. Annals of Statistics, 2 (3):602–608, May 1974. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176342724.

Bickel:1973:SAL

[Bic73a] P. J. Bickel. On some analogues to linear combinations of order statistics in the linear model. Annals of Statistics, 1(4):597– 616, July 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176342457.

Bickel:1973:ASB

[Bic73b] P. J. Bickel. On the asymptotic shape of Bayesian sequential tests of θ ≤ 0 versus θ > 0 for exponential families. Annals of Statistics, 1(2):231–240, March 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176342361.

Bickel:1974:EEN

[Bic74] P. J. Bickel. Edgeworth expansions in nonparametric statistics. Annals of Statistics, 2(1):1–20, January 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176342609.

Bickel:1978:URR

[Bic78] P. J. Bickel. Using residuals robustly i: Tests for heteroscedasticity, nonlinearity. Annals of Statistics, 6(2):266–291, March 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176344124.

Bingham:1974:ASD

[Bin74] Christopher Bingham. An antipodally symmetric distribution on the sphere. Annals of Statistics, 2(6):1201–1225, November 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176342874.

Bjerve:1977:EBL

[Bje77] Steinar Bjerve. Error bounds for linear combinations of order statistics. Annals of Statistics, 5(2):357–369, March 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343800.

Bhattacharyya:1977:CMS

[BJM77] G. K. Bhattacharyya, Richard A. Johnson, and K. G. Mehrotra. On the completeness of minimal sufficient statistics with censored observations. Annals of Statistics, 5(3):547–553, May 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343854.

Baksalary:1976:EME

[BK76] J. K. Baksalary and R. Kala. Extensions of Milliken’s estimability criterion. Annals of Statistics, 4(3):639–641, May 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343471.

Bonner:1977:NCW

[BK77] N. Bonner and H.-P. Kirschner. Note on conditions for weak convergence of von Mises’ differentiable statistical functions. Annals of Statistics, 5(2):405–407, March 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176343807.

Baksalary:1978:BEN

[BK78a] J. K. Baksalary and R. Kala. A bound for the Euclidean norm of the difference between the least squares and the best linear unbiased estimators. Annals of Statistics, 6(6):1390–1393, November 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176344383.

Brown:1978:RSH

[BK78b] B. M. Brown and D. G. Kildea. Reduced U-statistics and the Hodges–Lehmann estimator. Annals of Statistics, 6(4):828–835, July 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176344256.

Bevan:1979:ACM

[BKR79] Stuart Bevan, Richard Kullberg, and John Rice. An analysis of Cell membrane noise. Annals of Statistics, 7(2):237–257, March 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176344609.

Bickel:1975:DSNa

[BL75a] P. J. Bickel and E. L. Lehmann. Descriptive statistics for nonparametric models I. Introduction. Annals of Statistics, 3(5):1038–1044, September 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343239.

Bickel:1975:DSNb

[BL75b] P. J. Bickel and E. L. Lehmann. Descriptive statistics for nonparametric models II. location. Annals of Statistics, 3(5):1045–1069, September 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343240.

Bickel:1976:DSN

[BL76] P. J. Bickel and E. L. Lehmann. Descriptive statistics for nonparametric models. III. dispersion. Annals of Statistics, 4(6):1139–1158, November 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343648.

Brockett:1977:CR

[BL77] Patrick L. Brockett and Arnold Levine. On a characterization of Ridits. Annals of Statistics, 5(6):1245–1248, November 1977. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176344010.

Blackwell:1973:DFS

[Bla73] David Blackwell. Discreteness of Ferguson selections. Annals of Statistics, 1(2):356–358, March 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176342373.

Blackwell:1976:SPB

[Bla76] David Blackwell. The stochastic processes of Borel gambling and dynamic programming. Annals of Statistics, 4(2):370–374, March 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343412.

Blaesild:1979:CCS

[Bla79] P. Blaesild. Conditioning with conic sections in the two-dimensional normal distribution. Annals of Statistics, 7(3):659–670, May 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176344686.

Block:1977:CBE

[Blo77] Henry W. Block. A characterization of a bivariate exponential distribution. Annals of Statistics, 5(4):808–812, July 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343905.

Blumenthal:1974:ATI

[Blu74] Saul Blumenthal. Admissibility of translation invariant tolerance intervals in the location parameter case. Annals of Statistics, 2 (4):694–702, July 1974. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176342757.

Blumenthal:1976:SEL

[Blu76] Saul Blumenthal. Sequential estimation of the largest normal mean when the variance is known. Annals of Statistics, 4(6):1077–1087, November 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343643.

Blyth:1974:NSC

[Bly74] Colin R. Blyth. Necessary and sufficient conditions for inequalities of Cramér–Rao type. Annals of Statistics, 2(3):464–473, May 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176342707.

Basar:1973:MEM

[BM73a] Tamer Basar and Max Mintz. On a minimax estimate for the mean of a normal random vector under a generalized quadratic loss function. Annals of Statistics, 1(1):127–134, January 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1193342388.

Bhattacharya:1973:ANS

[BM73b] P. K. Bhattacharya and Ashim Mallik. Asymptotic normality of the stopping times of some sequential procedures. Annals of Statistics, 1 (6):1203–1211, November 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176342570.

Blackwell:1973:FDP

[BM73c] David Blackwell and James B. MacQueen. Ferguson distributions via Pólya urn schemes. Annals of Statistics, 1(2):353–355, March 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176342372.

Behnen:1975:CLT

[BN75] K. Behnen and G. Neuhaus. A central limit theorem under contiguous alternatives. Annals of Statistics, 3(6):1349–1353, November 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343289.

Bock:1975:MEM

[Boc75] M. E. Bock. Minimax estimators of the mean of a multivariate normal distribution. Annals of Statistics, 3(1):209–218, January 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343009.

Bohrer:1973:OPS

[Boh73] Robert Bohrer. An optimality property of scheffe bounds. Annals of Statistics, 1(4):766–772, July 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176342473.

Bondesson:1975:UMV

[Bon75] Lennart Bondesson. Uniformly minimum variance estimation in location parameter families. Annals of Statistics, 3(3):637–660, May 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343127.

Bondar:1976:BCS

[Bon76a] James V. Bondar. Borel cross-sections and maximal invariants. Annals of Statistics, 4(5):866–877, September 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176343585.

Bondesson:1976:WMS

[Bon76b] Lennart Bondesson. When are the mean and the studentized differences independent? Annals of Statistics, 4(3):668–672, May 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343477.

Bondar:1977:CCP

[Bon77] James V. Bondar. A conditional confidence principle. Annals of Statistics, 5(5):881–891, September 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176343944.

Boos:1979:DS

[Boo79] Dennis D. Boos. A differential for L-statistics. Annals of Statistics, 7(5):955–959, September 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176344781.

Brown:1973:MSE

[BP73] L. D. Brown and R. Purves. Measurable selections of extrema. Annals of Statistics, 1(5):902–912, September 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176342510.

Bergstrom:1977:CRT

[BP77] Harald Bergström and Madan L. Puri. Convergence and remainder terms in linear rank statistics. Annals of Statistics, 5(4):671– 680, July 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343891.

Bickel:1973:SGM

[BR73] P. J. Bickel and M. Rosenblatt. On some global measures of the deviations of density function estimates. Annals of Statistics, 1(6): 1071–1095, November 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/ euclid.aos/1176342558. See corrections [BR75].

Bickel:1975:NCS

[BR75] P. J. Bickel and M. Rosenblatt. Notes: Corrections to “On Some Global Measures of the Deviations of Density Function Estimates”. Annals of Statistics, 3(6):1370, November 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176343293. See [BR73].

Barr:1977:IPD

[BR77] D. R. Barr and F. R. Richards. Induced priors in decision problems. Annals of Statistics, 5(1):182–184, January 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343751.

Braun:1976:WCS

[Bra76] Henry I. Braun. Weak convergence of sequential linear rank statistics. Annals of Statistics, 4(3):554–575, May 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343461.

Brewster:1974:AES

[Bre74] J. F. Brewster. Alternative estimators for the scale parameter of the exponential distribution with unknown location. Annals of Statistics, 2(3):553–557, May 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176342715.

Breth:1978:BCB

[Bre78] M. Breth. Bayesian confidence bands for a distribution function. Annals of Statistics, 6(3):649–657, May 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176344209.

Brooks:1973:CVP

[Bro73a] R. J. Brooks. The choice of variables for prediction in curvilinear multiple regression. Annals of Statistics, 1(3):506–516, May 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176342416.

Brown:1973:NCA

[Bro73b] L. D. Brown. Notes: Correction to admissible estimators, recurrent diffusions, and insoluble boundary value problems. Annals of Statistics, 1(3):594–596, May 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176342433.

Brown:1975:TMC

[Bro75] L. Brown. On a theorem of Morimoto concerning sufficiency for discrete distributions. Annals of Statistics, 3(5):1180–1182, September 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343249.

Brown:1976:ABM

[Bro76] K. G. Brown. Asymptotic behavior of Minque-type estimators of variance components. Annals of Statistics, 4(4):746–754, July 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343546.

Brown:1978:CKT

[Bro78] Lawrence D. Brown. A contribution to Kiefer’s theory of conditional confidence procedures. Annals of Statistics, 6(1):59–71, January 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176344065.

Brown:1979:HMD

[Bro79] Lawrence D. Brown. A heuristic method for determining admissibility of estimators–with applications. Annals of Statistics, 7(5):960– 994, September 1979. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176344782.

Berger:1978:GBE

[BS78a] James O. Berger and C. Srinivasan. Generalized Bayes estimators in multivariate problems. Annals of Statistics, 6(4):783–801, July 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176344252.

Boyd:1978:LBN

[BS78b] David W. Boyd and J. Michael Steele. Lower bounds for nonparametric density estimation rates. Annals of Statistics, 6(4):932– 934, July 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176344269.

Brandwein:1978:MEL

[BS78c] Ann R. Cohen Brandwein and William E. Strawderman. Minimax estimation of location parameters for spherically symmetric unimodal distributions under quadratic loss. Annals of Statistics, 6(2): 377–416, March 1978. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176344131.

Block:1979:SEL

[BS79a] Henry W. Block and Thomas H. Savits. Systems with exponential life and IFRA component lives. Annals of Statistics, 7(4):911– 916, July 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176344740.

Brown:1979:CPN

[BS79b] Mark Brown and Herbert Solomon. On combining pseudorandom number generators. Annals of Statistics, 7(3):691–695, May 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176344690;   http://www.jstor.org/stable/2958754.

Buehler:1976:CP

[Bue76] Robert J. Buehler. Coherent preferences. Annals of Statistics, 4 (6):1051–1064, November 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176343641.

Burdick:1973:BST

[Bur73] David L. Burdick. A best sequential test for symmetry when the probability of termination is not one. Annals of Statistics, 1(6): 1195–1199, November 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176342568.

Butcher:1977:ESD

[But77] A. C. Butcher. An estimator for the spectral density of a stationary time sequence. Annals of Statistics, 5(3):541–543, May 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343852.

Birnbaum:1973:LDS

[BV73] Z. W. Birnbaum and I. Vincze. Limiting distributions of statistics similar to Student’s t. Annals of Statistics, 1(5):958–963, September 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176342517.

Bickel:1978:AEP

[BvZ78] P. J. Bickel and W. R. van Zwet. Asymptotic expansions for the power of distribution-free tests in the two-sample problem. Annals of Statistics, 6(5):937–1004, September 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176344305.

Berry:1977:NIS

[BY77] Donald A. Berry and D. H. Young. A note on inverse sampling procedures for selecting the best binomial population. Annals of Statistics, 5(1):235–236, January 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176343759.

Brewster:1974:IEE

[BZ74] J. F. Brewster and J. V. Zidek. Improving on equivariant estimators. Annals of Statistics, 2(1):21–38, January 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176342610.

Cabilio:1977:SEB

[Cab77] Paul Cabilio. Sequential estimation in Bernoulli trials. Annals of Statistics, 5(2):342–356, March 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176343799.

Caines:1975:NCM

[Cai75] P. E. Caines. A note on the consistency of maximum likelihood estimates for finite families of stochastic processes. Annals of Statistics, 3(2):539–546, March 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176343086.

Cane:1979:NRS

[Can79] Gwenda J. Cane. Note on a result of Seeger in partitioning normal populations. Annals of Statistics, 7(4):917–919, July 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176344741.

Carroll:1977:USP

[Car77] Raymond J. Carroll. On the uniformity of sequential procedures. Annals of Statistics, 5(5):1039–1046, September 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343958.

Carroll:1978:ASE

[Car78] Raymond J. Carroll. On almost sure expansions for M-estimates. Annals of Statistics, 6(2):314–318, March 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176344126.

Causey:1975:NCS

[Cau75a] B. Causey. Note: Correction to “Upper Confidence and Fiducial Limits to the Range of Several Means”. Annals of Statistics, 3(5): 1188, September 1975. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176343251. See [Cau75b].

Causey:1975:UCF

[Cau75b] B. Causey. Upper confidence and fiducial limits to the range of several means. Annals of Statistics, 3(2):473–477, March 1975. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343074. See

correction [Cau75a].

Casady:1976:MPR

[CC76] Robert J. Casady and Jonathan D. Cryer. Monotone percentile regression. Annals of Statistics, 4(3):532–541, May 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343459.

Cogburn:1974:PSS

[CD74] Robert Cogburn and Herbert T. Davis. Periodic splines and spectral estimation. Annals of Statistics, 2(6):1108–1126, November 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176342868.

Campbell:1978:ROE

[CH78] Gregory Campbell and Myles Hollander. Rank order estimation with the Dirichlet prior. Annals of Statistics, 6(1):142–153, January 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176344073.

Christopeit:1979:CTR

[CH79] N. Christopeit and K. Helmes. A convergence theorem for random linear combinations of independent normal random variables. Annals of Statistics, 7(4):795–800, July 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176344729.

Chew:1973:POD

[Che73] Milton C. Chew. On pairing observations from a distribution with monotone likelihood ratio. Annals of Statistics, 1(3):433– 445, May 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176342409.

Cheng:1978:ODE

[Che78a] Ching-Shui Cheng. Optimal designs for the elimination of multi-way heterogeneity. Annals of Statistics, 6(6):1262–1272, November 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176344372.

Cheng:1978:OCA

[Che78b] Ching-Shui Cheng. Optimality of certain asymmetrical experimental designs. Annals of Statistics, 6(6):1239–1261, November 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176344371.

Chung:1979:HWP

[Chu79] K. L. Chung. Hsu’s work in probability. Annals of Statistics, 7 (3):479–483, May 1979. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176344669.

Callaert:1978:BET

[CJ78] Herman Callaert and Paul Janssen. The Berry–Esseen theorem for U-statistics. Annals of Statistics, 6(2):417–421, March 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176344132.

Crain:1975:ANP

[CM75] Bradford R. Crain and Ronnie L. Morgan. Asymptotic normality of the posterior distribution for exponential models. Annals of Statistics, 3(1):223–227, January 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176343011.

Cobb:1975:NSR

[Cob75] E. Benton Cobb. A note on sampling with replacement. Annals of Statistics, 3(2):500–503, March 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176343079.

Collins:1976:REL

[Col76] John R. Collins. Robust estimation of a location parameter in the presence of asymmetry. Annals of Statistics, 4(1):68–85, January 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343348.

Collins:1977:UBA

[Col77] John R. Collins. Upper bounds on asymptotic variances of Mestimators of location. Annals of Statistics, 5(4):646–657, July 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343889.

Conover:1973:RTO

[Con73] W. J. Conover. Rank tests for one sample, two samples, and k samples without the assumption of a continuous distribution function. Annals of Statistics, 1(6):1105–1125, November 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176342560.

Copas:1974:SCD

[Cop74] J. B. Copas. On symmetric compound decision rules for dichotomies. Annals of Statistics, 2(1):199–204, January 1974. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176342626.

Cover:1973:DIM

[Cov73] Thomas M. Cover. On determining the irrationality of the mean of a random variable. Annals of Statistics, 1(5):862–871, September 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176342507.

Chattopadhyay:1976:MIM

[CPL76] A. K. Chattopadhyay, K. C. S. Pillai, and Hung C. Li. Maximization of an integral of a matrix function and asymptotic expansions of distributions of latent roots of two matrices. Annals of Statistics, 4 (4):796–806, July 1976. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176343554.

Csorgo:1978:SAQ

[CR78] Miklós Csörgő and Pal Revesz. Strong approximations of the quantile process. Annals of Statistics, 6(4):882–894, July 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176344261.

Crain:1974:EDU

[Cra74] Bradford R. Crain. Estimation of distributions using orthogonal expansions. Annals of Statistics, 2(3):454–463, May 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176342706.

Craig:1978:HCC

[Cra78] Cecil C. Craig. Harry C. Carver, 1890–1977. Annals of Statistics, 6 (1):1–4, January 1978. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176344062.

Cohen:1973:AID

[CS73a] Arthur Cohen and William E. Strawderman. Admissibility implications for different criteria in confidence estimation. Annals of Statistics, 1(2):363–366, March 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176342375.

Cohen:1973:ACI

[CS73b] Arthur Cohen and William E. Strawderman. Admissible confidence interval and point estimation for translation or scale parameters. Annals of Statistics, 1(3):545–550, May 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176342421.

Cohen:1974:ECM

[CS74] Arthur Cohen and Harold B. Sackrowitz. On estimating the common mean of two normal distributions. Annals of Statistics, 2(6): 1274–1282, November 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176342878. See correction [CS76].

Cohen:1975:UCS

[CS75] Arthur Cohen and H. B. Sackrowitz. Unbiasedness of the chisquare, likelihood ratio, and other goodness of fit tests for the equal Cell case. Annals of Statistics, 3(4):959–964, July 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343197.

Cohen:1976:NCE

[CS76] Arthur Cohen and Harold B. Sackrowitz. Notes: Correction to “On Estimating the Common Mean of Two Normal Distributions”. Annals of Statistics, 4(6):1294, November 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176343662. See [CS74].

Cohen:1977:HTC

[CS77] Arthur Cohen and H. B. Sackrowitz. Hypotheses testing for the common mean and for balanced incomplete blocks designs. Annals of Statistics, 5(6):1195–1211, November 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176344004.

Csorgo:1979:ERL

[Cso79] Sandor Csorgo. Erdős–Rényi laws. Annals of Statistics, 7(4):772– 787, July 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176344727.

Clinger:1976:UST

[CV76] William Clinger and John W. Van Ness. On unequally spaced time points in time series. Annals of Statistics, 4(4):736–745, July 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343545.

Chang:1980:CDO

[CW80] Der-Shin Chang and Chi Song Wong. Correction to “Design of Optimal Control for a Regression Problem”. Annals of Statistics, 8(6): 1402, November 1980. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176345216. See [sC79].

Darst:1974:RMS

[Dar74] R. B. Darst. Remarks on measurable selections. Annals of Statistics, 2(4):845–847, July 1974. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176342775.

Davis:1975:MSE

[Dav75] Kathryn Bullock Davis. Mean square error properties of density estimates. Annals of Statistics, 3(4):1025–1030, July 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343207.

Davis:1977:MIS

[Dav77a] Kathryn Bullock Davis. Mean integrated square error properties of density estimates. Annals of Statistics, 5(3):530–535, May 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343850.

Davis:1977:RIE

[Dav77b] William W. Davis. Robust interval estimation of the innovation variance of an ARMA model. Annals of Statistics, 5(4):700–708, July 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343893.

Dawid:1977:FCS

[Daw77] A. P. Dawid. Further comments on some comments on a paper by Bradley Efron. Annals of Statistics, 5(6):1249, November 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176344011.

DeHaan:1974:SQR

[De 74] Laurens De Haan. On sample quantiles from a regularly varying distribution function. Annals of Statistics, 2(4):815–818, July 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176342769.

Dion:1979:EPB

[DE79] J. P. Dion and W. W. Esty. Estimation problems in branching processes with random environments. Annals of Statistics, 7(3): 680–685, May 1979. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176344688.

Deo:1973:SLT

[Deo73] Chandrakant M. Deo. Some limit theorems for maxima of nonstationary Gaussian processes. Annals of Statistics, 1(5):981–984, September 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176342520.

Devore:1976:NEP

[Dev76] Jay L. Devore. A note on the estimation of parameters in a Bernoulli model with dependence. Annals of Statistics, 4(5):990– 992, September 1976. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176343597.

Dykstra:1978:PDR

[DH78] Richard L. Dykstra and John E. Hewett. Positive dependence of the roots of a Wishart matrix. Annals of Statistics, 6(1):235–238, January 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176344083.

Dalal:1979:MER

[DH79] S. R. Dalal and W. J. Hall. Most economical robust selection procedures for location parameters. Annals of Statistics, 7(6):1321–1328, November 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176344849.

Dykstra:1973:EWA

[DHT73] R. L. Dykstra, John E. Hewett, and W. A. Thompson. Events which are almost independent. Annals of Statistics, 1(4):674–681, July 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176342462.

Dion:1975:EVB

[Dio75] Jean-Pierre Dion. Estimation of the variance of a branching process. Annals of Statistics, 3(5):1183–1187, September 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343250.

Divgi:1979:CUB

[Div79] D. R. Divgi. Calculation of univariate and bivariate normal probability functions. Annals of Statistics, 7(4):903–910, July 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176344739.

Dharmadhikari:1976:MU

[DJ76] S. W. Dharmadhikari and Kumar Jogdeo. Multivariate unimodality. Annals of Statistics, 4(3):607–613, May 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176343466.

Daniel:1979:HS

[DL79] C. Daniel and E. L. Lehmann. Henry Scheffe 1907–1977. Annals of Statistics, 7(6):1149–1161, November 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176344837.

deMontricher:1975:NML

[dMTT75] G. F. de Montricher, R. A. Tapia, and J. R. Thompson. Nonparametric maximum likelihood estimation of probability densities by penalty function methods. Annals of Statistics, 3(6):1329–1348, November 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343288.

Diaconis:1977:RMK

[DOG77] P. Diaconis, I. Olkin, and S. G. Ghurye. Review: A. M. Kagan, Yu. V. Linnik, C. Radhakrishna Rao, B. Ramachandran, Characterization Problems in Mathematical Statistics. Annals of Statistics, 5 (3):583–592, May 1977. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176343861.

Doksum:1974:EPP

[Dok74] Kjell Doksum. Empirical probability plots and statistical inference for nonlinear models in the two-sample case. Annals of Statistics, 2(2):267–277, March 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176342662.

Doshi:1976:CTC

[Dos76] Bharat T. Doshi. Continuous time control of Markov processes on an arbitrary state space: Discounted rewards. Annals of Statistics, 4 (6):1219–1235, November 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176343653.

David:1977:DEV

[DOY77] H. A. David, M. J. O’Connell, and S. S. Yang. Distribution and expected value of the rank of a concomitant of an order statistic. Annals of Statistics, 5(1):216–223, January 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343756.

delPino:1979:ADS

[dP79] Guido E. del Pino. On the asymptotic distribution of k-spacings with applications to goodness-of-fit tests. Annals of Statistics, 7(5): 1058–1065, September 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176344789.

Drygas:1976:GME

[Dry76] Hilmar Drygas. Gauss–Markov estimation for multivariate linear models with missing observations. Annals of Statistics, 4(4):779– 787, July 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343551.

Dawid:1973:ECG

[DS73] A. P. Dawid and M. Stone. Expectation consistency and generalized Bayes inference. Annals of Statistics, 1(3):478–485, May 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176342413.

DuMouchel:1973:ANM

[DuM73] William H. DuMouchel. On the asymptotic normality of the maximum-likelihood estimate when sampling from a stable distribution. Annals of Statistics, 1(5):948–957, September 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176342516.

Dunsmuir:1979:CLT

[Dun79] W. Dunsmuir. A central limit theorem for parameter estimation in stationary vector time series and its application to models for a signal observed with noise. Annals of Statistics, 7(3):490–506, May 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176344671.

Dupac:1975:JH

[Dup75] Vaclav Dupac. Jaroslav Hajek, 1926–1974. Annals of Statistics, 3 (5):1031–1037, September 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176343238.

Durbin:1973:WCS

[Dur73] J. Durbin. Weak convergence of the sample distribution function when parameters are estimated. Annals of Statistics, 1(2):279–290, March 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176342365.

Duttweiler:1973:MSE

[Dut73] D. L. Duttweiler. The mean-square error of Bahadur’s order-statistic approximation. Annals of Statistics, 1(3):446–453, May 1973. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176342410.

Devroye:1977:SUC

[DW77] Luc P. Devroye and T. J. Wagner. The strong uniform consistency of nearest neighbor density estimates. Annals of Statistics, 5(3): 536–540, May 1977. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176343851.

Devroye:1978:NAS

[DW78] L. P. Devroye and T. J. Wagner. Note: Addendum to “The Strong Uniform Consistency of Nearest Neighbor Density Estimates”. Annals of Statistics, 6(4):935, July 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176344270.

Devroye:1979:CKD

[DW79] L. P. Devroye and T. J. Wagner. The L_{1} convergence of kernel density estimates. Annals of Statistics, 7(5):1136–1139, September 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176344796.

deWet:1973:ADQ

[dWV73] T. de Wet and J. H. Venter. Asymptotic distributions for quadratic forms with applications to tests of fit. Annals of Statistics, 1(2): 380–387, March 1973. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176342378.

Diaconis:1979:CPE

[DY79] Persi Diaconis and Donald Ylvisaker. Conjugate priors for exponential families. Annals of Statistics, 7(2):269–281, March 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176344611.

Dykstra:1979:DTS

[Dyk79] Richard L. Dykstra. On dependent tests of significance in the multivariate analysis of variance. Annals of Statistics, 7(2):459–461, March 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176344628.

Deely:1976:AOE

[DZ76] J. J. Deely and W. J. Zimmer. Asymptotic optimality of the empirical Bayes procedure. Annals of Statistics, 4(3):576–580, May 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343462.

Eaton:1974:PIL

[Eat74] Morris L. Eaton. A probability inequality for linear combinations of bounded random variables. Annals of Statistics, 2(3):609–614, May 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176342725.

Eaton:1978:CCT

[Eat78] Morris L. Eaton. Complete class theorems derived from conditional complete class theorems. Annals of Statistics, 6(4):820– 827, July 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176344255.

Efron:1975:DCS

[Efr75] Bradley Efron. Defining the curvature of a statistical problem (with applications to second order efficiency). Annals of Statistics, 3(6): 1189–1242, November 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176343282.

Efron:1978:GEF

[Efr78] Bradley Efron. The geometry of exponential families. Annals of Statistics, 6(2):362–376, March 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176344130.

Efron:1979:RLB

[Efr79] B. Efron. The 1977 Rietz Lecture: Bootstrap methods: Another look at the jackknife. Annals of Statistics, 7(1):1–26, January 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (elec-

tronic). URL http://www.jstor.org/stable/2958830;   http:
//projecteuclid.org/euclid.aos/1176344552. See [Que49,

Que56, Tuk58, Mil74a].

Enis:1974:OPL

[EG74] Peter Enis and Seymour Geisser. Optimal predictive linear discriminants. Annals of Statistics, 2(2):403–410, March 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176342677.

Eisenberg:1976:PGS

[EGS76] Bennett Eisenberg, B. K. Ghosh, and Gordon Simons. Properties of generalized sequential probability ratio tests. Annals of Statistics, 4(2):237–251, March 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176343404.

Eccleston:1974:TCD

[EH74] J. A. Eccleston and A. Hedayat. On the theory of connected designs: Characterization and optimality. Annals of Statistics, 2(6):1238– 1255, November 1974. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176342876.

Eddy:1977:UCE

[EH77] W. F. Eddy and J. A. Hartigan. Uniform convergence of the empirical distribution function over convex sets. Annals of Statistics, 5(2): 370–374, March 1977. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176343801.

Eicker:1979:ADS

[Eic79] F. Eicker. The asymptotic distribution of the suprema of the standardized empirical processes. Annals of Statistics, 7(1):116–138, January 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176344559.

Erickson:1976:RCL

[EK76] R. V. Erickson and H. L. Koul. L_{1} rates of convergence for linear rank statistics. Annals of Statistics, 4(4):771–774, July 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343549.

Esary:1974:MDE

[EM74] James D. Esary and Albert W. Marshall. Multivariate distributions with exponential minimums. Annals of Statistics, 2(1):84–98, January 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176342615.

Efron:1976:FME

[EM76a] Bradley Efron and Carl Morris. Families of minimax estimators of the mean of a multivariate normal distribution. Annals of Statistics, 4(1):11–21, January 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176343344.

Efron:1976:MEB

[EM76b] Bradley Efron and Carl Morris. Multivariate empirical Bayes and estimation of covariance matrices. Annals of Statistics, 4(1):22–32, January 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343345.

Enns:1970:DPM

[Enn70] E. G. Enns. The decision process for maximizing the probability of obtaining the random variable nearest to an arbitrary real number. Annals of Mathematical Statistics, 41(5):1466–1471, October 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177696792. See

correction [Enn73].

Enns:1973:CNC

[Enn73] E. G. Enns. Correction notes: Correction to “The Optimum Strategy for Maximising the Probability of Obtaining the Random Variable Nearest to an Arbitrary Real Number”. Annals of Statistics, 1(2):400, March 1973. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176342382. See [Enn70].

Efron:1978:HBC

[EO78] Bradley Efron and Richard A. Olshen. How broad is the class of normal scale mixtures? Annals of Statistics, 6(5):1159–1164, September 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176344318.

Eaton:1973:NSG

[EP73] Morris L. Eaton and Michael D. Perlman. The non-singularity of generalized sample covariance matrices. Annals of Statistics, 1(4): 710–717, July 1973. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176342465.

Eaton:1974:MPP

[EP74] Morris L. Eaton and Michael D. Perlman. A monotonicity property of the power functions of some invariant tests for MANOVA. Annals of Statistics, 2(5):1022–1028, September 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176342821.

Eplett:1979:SSD

[Epl79] W. J. R. Eplett. The small sample distribution of a Mann–Whitney type statistic for circular data. Annals of Statistics, 7(2):446–453, March 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176344626.

Eisenberg:1978:WAT

[ES78] Bennett Eisenberg and Gordon Simons. On weak admissibility of tests. Annals of Statistics, 6(2):319–332, March 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176344127.

Fabian:1973:AES

[Fab73a] Vaclav Fabian. Asymptotically efficient stochastic approximation; the RM case. Annals of Statistics, 1(3):486–495, May 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176342414.

Fabian:1973:EDL

[Fab73b] Vaclav Fabian. Estimation of the derivative of the logarithm of a density. Annals of Statistics, 1(3):557–561, May 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176342424.

Fabius:1973:TCD

[Fab73c] J. Fabius. Two characterizations of the Dirichlet distribution. Annals of Statistics, 1(3):583–587, May 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http:

//projecteuclid.org/euclid.aos/1176342429. See correction

[Fab81].

Fabian:1974:NAP

[Fab74a] Vaclav Fabian. Note: Acknowledgment of priority to “Note on Anderson’s Sequential Procedures with Triangular Boundary”. Annals of Statistics, 2(5):1063, September 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176342828.

Fabian:1974:NAS

[Fab74b] Vaclav Fabian. Note on Anderson’s sequential procedures with triangular boundary. Annals of Statistics, 2(1):170–176, January 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176342622.

Fabian:1978:AER

[Fab78] Vaclav Fabian. On asymptotically efficient recursive estimation. Annals of Statistics, 6(4):854–866, July 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176344259.

Fabius:1981:CTC

[Fab81] J. Fabius. Correction to “Two Characterizations of the Dirichlet Distribution”. Annals of Statistics, 9(1):234, January 1981. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345356. See

[Fab73c].

Francis:1974:CES

[FC74] Ivor Francis and Samprit Chatterjee. Classification and estimation of several multiple regressions. Annals of Statistics, 2(3):558– 561, May 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176342716.

Feder:1975:LLR

[Fed75a] Paul I. Feder. The log likelihood ratio in segmented regression. Annals of Statistics, 3(1):84–97, January 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176343000.

Feder:1975:ADT

[Fed75b] Paul I. Feder. On asymptotic distribution theory in segmented regression problems– identified case. Annals of Statistics, 3(1):49–83, January 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176342999.

Federer:1977:ECC

[Fed77] Walter T. Federer. On the existence and construction of a complete set of orthogonal F (4t; 2t, 2t)-squares design. Annals of Statistics, 5 (3):561–564, May 1977. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176343856.

Fenech:1976:AEE

[Fen76] Alan Paul Fenech. Asymptotically efficient estimation of location for a symmetric stable law. Annals of Statistics, 4(6):1088–1100, November 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343644.

Ferguson:1973:BAS

[Fer73] Thomas S. Ferguson. A Bayesian analysis of some nonparametric problems. Annals of Statistics, 1(2):209–230, March 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176342360.

Ferguson:1974:PDS

[Fer74] Thomas S. Ferguson. Prior distributions on spaces of probability measures. Annals of Statistics, 2(4):615–629, July 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176342752.

Ferguson:1976:SSD

[Fer76] Thomas S. Ferguson. Stopping a sum during a success run. Annals of Statistics, 4(2):252–264, March 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176343405.

Fabian:1977:CRI

[FH77a] Vaclav Fabian and James Hannan. On the Cramér–Rao inequality. Annals of Statistics, 5(1):197–205, January 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343754.

Foody:1977:TAB

[FH77b] W. Foody and A. Hedayat. On theory and applications of BIB designs with repeated blocks. Annals of Statistics, 5(5):932–945, September 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/

1176343949. See correction [FH79].

Foody:1979:NCT

[FH79] W. Foody and A. Hedayat. Note: Correction to “On Theory and Application of BIB Designs with Repeated Blocks”. Annals of Statistics, 7(4):925, July 1979. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176344743. See [FH77b].

Findeisen:1978:SPC

[Fin78] Peter Findeisen. A simple proof of a classical theorem which characterizes the gamma distribution. Annals of Statistics, 6(5):1165– 1167, September 1978. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176344319.

Fishburn:1975:UEU

[Fis75] Peter C. Fishburn. Unbounded expected utility. Annals of Statistics, 3(4):884–896, July 1975. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176343189.

Flueck:1975:GAM

[FK75] John A. Flueck and James F. Korsh. A generalized approach to maximum likelihood paired comparison ranking. Annals of Statistics, 3(4):846–861, July 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176343186.

Fleming:1978:ADR

[Fle78a] Thomas R. Fleming. Asymptotic distribution results in competing risks estimation. Annals of Statistics, 6(5):1071–1079, September 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176344311.

Fleming:1978:NEN

[Fle78b] Thomas R. Fleming. Nonparametric estimation for nonhomogeneous Markov processes in the problem of competing risks. Annals of Statistics, 6(5):1057–1070, September 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176344310.

Flynn:1974:AVD

[Fly74] James Flynn. Averaging vs. discounting in dynamic programming: a counterexample. Annals of Statistics, 2(2):411–413, March 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176342678.

Flynn:1976:CEO

[Fly76] James Flynn. Conditions for the equivalence of optimality criteria in dynamic programming. Annals of Statistics, 4(5):936–953, September 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343590.

Fears:1974:WCT

[FM74] Thomas R. Fears and K. L. Mehra. Weak convergence of a twosample empirical process and a Chernoff–Savage theorem for ϕmixing sequences. Annals of Statistics, 2(3):586–596, May 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176342721.

Fraser:1975:PFI

[FM75] D. A. S. Fraser and Jock MacKay. Parameter factorization and inference based on significance, likelihood, and objective posterior. Annals of Statistics, 3(3):559–572, May 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176343122.

Feuerverger:1977:ECF

[FM77] Andrey Feuerverger and Roman A. Mureika. The empirical characteristic function and its applications. Annals of Statistics, 5(1): 88–97, January 1977. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176343742.

Fortus:1979:ABS

[For79] Robert Fortus. Approximations to Bayesian sequential tests of composite hypotheses. Annals of Statistics, 7(3):579–591, May 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176344679.

Fox:1978:SEB

[Fox78] Richard J. Fox. Solutions to empirical Bayes squared error loss estimation problems. Annals of Statistics, 6(4):846–853, July 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176344258.

Ferguson:1979:BNE

[FP79] Thomas S. Ferguson and Eswar G. Phadia. Bayesian nonparametric estimation based on censored data. Annals of Statistics, 7(1):163– 186, January 1979. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176344562.

Fox:1973:APM

[FR73] Bennett L. Fox and John E. Rolph. Adaptive policies for Markov renewal programs. Annals of Statistics, 1(2):334–341, March 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176342370.

Fuller:1978:ELR

[FR78] Wayne A. Fuller and J. N. K. Rao. Estimation for a linear regression model with unknown diagonal covariance matrix. Annals of Statistics, 6(5):1149–1158, September 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176344317.

Friedman:1979:MGW

[FR79] Jerome H. Friedman and Lawrence C. Rafsky. Multivariate generalizations of the Wald–Wolfowitz and Smirnov two-sample tests.

Annals of Statistics, 7(4):697–717, July 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176344722.

Fraser:1977:CPP

[Fra77] D. A. S. Fraser. Confidence, posterior probability, and the Buehler example. Annals of Statistics, 5(5):892–898, September 1977. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343945.

Foutz:1977:PLR

[FS77] Robert V. Foutz and R. C. Srivastava. The performance of the likelihood ratio test when the model is incorrect. Annals of Statistics, 5 (6):1183–1194, November 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176344003.

Fu:1973:TBR

[Fu73] James C. Fu. On a theorem of Bahadur on the rate of convergence of point estimators. Annals of Statistics, 1(4):745–749, July 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176342469.

Fu:1975:RCC

[Fu75] James C. Fu. The rate of convergence of consistent point estimators. Annals of Statistics, 3(1):234–240, January 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343013.

Fujikoshi:1973:MPF

[Fuj73] Yasunori Fujikoshi. Monotonicity of the power functions of some tests in general manova models. Annals of Statistics, 1(2):388–391, March 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176342379.

Fujii:1976:UBR

[Fuj76] Yoshio Fujii. An upper bound of resolution in symmetrical fractional factorial designs. Annals of Statistics, 4(3):662–667, May 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343476.

Guess:1978:MLE

[GC78] H. A. Guess and K. S. Crump. Maximum likelihood estimation of dose-response functions subject to absolutely monotonic constraints. Annals of Statistics, 6(1):101–111, January 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176344069.

Gray:1975:EEU

[GCL75] H. L. Gray, W. A. Coberly, and T. O. Lewis. On Edgeworth expansions with unknown cumulants. Annals of Statistics, 3(3):741– 746, May 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343139.

Goel:1979:CEI

[GD79] Prem K. Goel and Morris H. DeGroot. Comparison of experiments and information measures. Annals of Statistics, 7(5):1066–1077, September 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176344790.

Gillespie:1979:CBK

[GF79] Mary Jo Gillespie and Lloyd Fisher. Confidence bands for the Kaplan–Meier survival curve estimate. Annals of Statistics, 7(4): 920–924, July 1979. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176344742.

Gupta:1973:PTC

[GG73] S. Das Gupta and N. Giri. Properties of tests concerning covariance matrices of normal distributions. Annals of Statistics, 1(6):1222– 1224, November 1973. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176342572.

Gilliland:1977:IRE

[GH77a] Dennis C. Gilliland and James Hannan. Improved rates in the empirical Bayes monotone multiple decision problem MLR family. Annals of Statistics, 5(3):516–521, May 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176343848.

Gordon:1977:CMD

[GH77b] Louis Gordon and Malcolm Hudson. A characterization of the von Mises distribution. Annals of Statistics, 5(4):813–814, July 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343906.

Gross:1978:ATA

[GH78] Alan J. Gross and David W. Hosmer. Approximating tail areas of probability distributions. Annals of Statistics, 6(6):1352–1359, November 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176344380.

Gordon:1979:APC

[GH79] Louis Gordon and Malcolm Hudson. Acknowledgment of priority to “A Characterization of the von Mises Distribution”. Annals of Statistics, 7(4):927, July 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176344745.

Gilliland:1976:AST

[GHH76] Dennis C. Gilliland, James Hannan, and J. S. Huang. Asymptotic solutions to the two state component compound decision problem, Bayes versus diffuse priors on proportions. Annals of Statistics, 4 (6):1101–1112, November 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176343645.

Ghosh:1973:NSP

[Gho73] Malay Ghosh. Nonparametric selection procedures for symmetric location parameter populations. Annals of Statistics, 1(4):773– 779, July 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176342474.

Gilliland:1977:NOP

[Gil77] Dennis C. Gilliland. A note on orthogonal partitions and some wellknown structures in design of experiments. Annals of Statistics, 5 (3):565–570, May 1977. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176343857.

Gleser:1976:AOS

[GK76] Leon Jay Gleser and Sudhakar Kunte. On asymptotically optimal sequential Bayes interval estimation procedures. Annals of Statistics, 4(4):685–711, July 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176343542.

Gleser:1979:MEN

[Gle79] Leon Jay Gleser. Minimax estimation of a normal mean vector when the covariance matrix is unknown. Annals of Statistics, 7 (4):838–846, July 1979. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176344733.

Ghosh:1977:ALE

[GM77] Malay Ghosh and Glen Meeden. Admissibility of linear estimators in the one parameter exponential family. Annals of Statistics, 5 (4):772–778, July 1977. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176343899.

Gordon:1978:AES

[GO78] Louis Gordon and Richard A. Olshen. Asymptotically efficient solutions to the classification problem. Annals of Statistics, 6(3):515– 533, May 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176344197.

Goel:1975:RPO

[Goe75] Prem K. Goel. On re-pairing observations in a broken random sample. Annals of Statistics, 3(6):1364–1369, November 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343292.

Goldstein:1975:ABS

[Gol75a] M. Goldstein. Approximate Bayes solutions to some nonparametric problems. Annals of Statistics, 3(2):512–517, March 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343081.

Goldstein:1975:NSB

[Gol75b] M. Goldstein. A note on some Bayesian nonparametric estimates. Annals of Statistics, 3(3):736–740, May 1975. CODEN ASTSC7.

ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176343138.

Goldstein:1977:CBE

[Gol77] Michael Goldstein. On contractions of Bayes estimators for exponential family distributions. Annals of Statistics, 5(6):1235–1239, November 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176344008.

Goodman:1964:SCI

[Goo64] Leo A. Goodman. Simultaneous confidence intervals for contrasts among multinomial populations. Annals of Mathematical Statistics, 35(2):716–725, June 1964. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http://projecteuclid.org/ euclid.aoms/1177703569. See correction [Goo76b].

Good:1973:JPG

[Goo73] I. J. Good. The joint probability generating function for run-lengths in regenerative binary Markov chains, with applications. Annals of Statistics, 1(5):933–939, September 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176342513.

Good:1975:BFA

[Goo75] I. J. Good. The Bayes factor against equiprobability of a multinomial population assuming a symmetric Dirichlet prior. Annals of Statistics, 3(1):246–250, January 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176343015.

Good:1976:ASD

[Goo76a] I. J. Good. On the application of symmetric Dirichlet distributions and their mixtures to contingency tables. Annals of Statistics, 4 (6):1159–1189, November 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176343649.

Goodman:1976:NCS

[Goo76b] Leo A. Goodman. Notes: Correction to “Simultaneous Confidence Intervals for Contrasts Among Multinomial Populations”. Annals of Statistics, 4(4):821, July 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176343558. See [Goo64].

Good:1977:NFS

[Goo77] I. J. Good. A new formula for k-statistics. Annals of Statistics, 5(1):224–228, January 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176343757.

Gordon:1973:CPU

[Gor73] F. S. Gordon. Characterizations of populations using regression properties. Annals of Statistics, 1(1):114–126, January 1973. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1193342387.

Gordon:1974:CSG

[Gor74a] Louis Gordon. Completely separating groups in subsampling. Annals of Statistics, 2(3):572–578, May 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176342719.

Gordon:1974:ES

[Gor74b] Louis Gordon. Efficiency in subsampling. Annals of Statistics, 2 (4):739–750, July 1974. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176342761.

Gleser:1973:ACQ

[GP73] Leon Jay Gleser and Marcello Pagano. Approximating circulant quadratic forms in jointly stationary Gaussian time series. Annals of Statistics, 1(2):322–333, March 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176342369.

Guseman:1975:MPM

[GPW75] L. F. Guseman, B. Charles Peters, and Homer F. Walker. On minimizing the probability of misclassification for linear feature selection. Annals of Statistics, 3(3):661–668, May 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343128.

Gastwirth:1975:ADT

[GR75a] Joseph L. Gastwirth and Herman Rubin. The asymptotic distribution theory of the empiric CDF for mixing stochastic processes. Annals of Statistics, 3(4):809–824, July 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343184.

Gastwirth:1975:BRE

[GR75b] Joseph L. Gastwirth and Herman Rubin. The behavior of robust estimators on dependent data. Annals of Statistics, 3(5):1070–1100, September 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343241.

Goel:1977:SSC

[GR77] Prem K. Goel and Herman Rubin. On selecting a subset containing the best population — a Bayesian approach. Annals of Statistics, 5(5):969–983, September 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176343952.

Green:1974:NOP

[Gre74a] Richard F. Green. A note on outlier-prone families of distributions. Annals of Statistics, 2(6):1293–1295, November 1974. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176342880.

Grenander:1974:LSD

[Gre74b] Ulf Grenander. Large sample discrimination between two Gaussian processes with different spectra. Annals of Statistics, 2(2):347–352, March 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176342668.

Gregory:1977:LST

[Gre77] Gavin G. Gregory. Large sample theory for U-statistics and tests of fit. Annals of Statistics, 5(1):110–123, January 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343744.

Grams:1973:CRS

[GS73] William F. Grams and R. J. Serfling. Convergence rates for Ustatistics and related statistics. Annals of Statistics, 1(1):153–160, January 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1193342392.

Godambe:1973:EST

[GT73] V. P. Godambe and M. E. Thompson. Estimation in sampling theory with exchangeable prior distributions. Annals of Statistics, 1 (6):1212–1221, November 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176342571.

Godambe:1974:EEP

[GT74a] V. P. Godambe and M. E. Thompson. Estimating equations in the presence of a nuisance parameter. Annals of Statistics, 2(3): 568–571, May 1974. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176342718.

Guttman:1974:IVM

[GT74b] Irwin Guttman and W. Y. Tan. Inconsistencies in the Villegas method of determining a prior distribution. Annals of Statistics, 2(2):383–386, March 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176342674.

Gupta:1974:PIE

[Gup74] Somesh Das Gupta. Probability inequalities and errors in classification. Annals of Statistics, 2(4):751–762, July 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176342762.

Hedayat:1978:RMD

[HA78] A. Hedayat and K. Afsarinejad. Repeated measurements designs, II. Annals of Statistics, 6(3):619–628, May 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176344206.

Haberman:1973:LLM

[Hab73] Shelby J. Haberman. Log-linear models for frequency data: Sufficient statistics and likelihood equations. Annals of Statistics, 1 (4):617–632, July 1973. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176342458.

Haberman:1974:LLM

[Hab74] Shelby J. Haberman. Log-linear models for frequency tables derived by indirect observation: Maximum likelihood equations. Annals of Statistics, 2(5):911–924, September 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176342813.

Haberman:1975:DPL

[Hab75a] Shelby J. Haberman. Direct products and linear models for complete factorial tables. Annals of Statistics, 3(2):314–333, March 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343059.

Haberman:1975:HMD

[Hab75b] Shelby J. Haberman. How much do Gauss–Markov and least square estimates differ? A coordinate-free approach. Annals of Statistics, 3 (4):982–990, July 1975. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176343201.

Haberman:1976:RMM

[Hab76] Shelby J. Haberman. Review: Y. M. M. Bishop, S. E. Fienberg, P. W. Holland, Discrete Multivariate Analysis: Theory and Practice. Annals of Statistics, 4(4):817–820, July 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343556.

Haberman:1977:LLM

[Hab77a] Shelby J. Haberman. Log-linear models and frequency tables with small expected Cell counts. Annals of Statistics, 5(6):1148–1169, November 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176344001.

Haberman:1977:MLE

[Hab77b] Shelby J. Haberman. Maximum likelihood estimates in exponential response models. Annals of Statistics, 5(5):815–841, September 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343941.

Haberman:1977:PMF

[Hab77c] Shelby J. Haberman. Product models for frequency tables involving indirect observation. Annals of Statistics, 5(6):1124–1147, November 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176344000.

Haff:1979:EIC

[Haf79] L. R. Haff. Estimation of the inverse covariance matrix: Random mixtures of the inverse Wishart matrix and the identity. Annals of Statistics, 7(6):1264–1276, November 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176344845. See cor-

rection [Haf81].

Haff:1981:CEI

[Haf81] L. R. Haff. Corrections to “Estimation of the Inverse Covariance Matrix: Random Mixtures of the Inverse Wishart Matrix and the Identity”. Annals of Statistics, 9(5):1132, September 1981. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345596. See

[Haf79].

Hajek:1974:ASV

[Haj74] Jaroslav Hajek. Asymptotic sufficiency of the vector of ranks in the Bahadur sense. Annals of Statistics, 2(1):75–83, January 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176342614.

Halpern:1974:PCC

[Hal74] Elkan F. Halpern. Posterior consistency for coefficient estimation and model selection in the general linear hypothesis. Annals of Statistics, 2(4):703–712, July 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176342758.

Hall:1976:SSR

[Hal76] Gaineford J. Hall. Sequential search with random overlook probabilities. Annals of Statistics, 4(4):807–816, July 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343555.

Hall:1977:SOP

[Hal77] Gaineford J. Hall. Strongly optimal policies in sequential search with random overlook probabilities. Annals of Statistics, 5(1):124– 135, January 1977. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176343745.

Hamada:1974:GMC

[Ham74] Noboru Hamada. On a geometrical method of construction of group divisible incomplete block designs. Annals of Statistics, 2(6):1351– 1356, November 1974. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176342890.

Hannan:1974:UCA

[Han74] E. J. Hannan. The uniform convergence of autocovariances. Annals of Statistics, 2(4):803–806, July 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176342767.

Hannan:1975:EAM

[Han75] E. J. Hannan. The estimation of ARMA models. Annals of Statistics, 3(4):975–981, July 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/ euclid.aos/1176343200. See correction [Han81].

Hannan:1976:ADS

[Han76a] E. J. Hannan. The asymptotic distribution of serial covariances. Annals of Statistics, 4(2):396–399, March 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176343415.

Hannan:1976:CSR

[Han76b] E. J. Hannan. The convergence of some recursions. Annals of Statistics, 4(6):1258–1270, November 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176343658.

Hannan:1981:CEA

[Han81] E. J. Hannan. Correction to “The Estimation of ARMA Models”. Annals of Statistics, 9(1):233, January 1981. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345355. See [Han75].

Hartigan:1975:NSC

[Har75] J. A. Hartigan. Necessary and sufficient conditions for asymptotic joint normality of a statistic and its subsample values. Annals of Statistics, 3(3):573–580, May 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176343123.

Harville:1976:EGM

[Har76] David Harville. Extension of the Gauss–Markov Theorem to include the estimation of random effects. Annals of Statistics, 4(2):384–395, March 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343414.

Hartigan:1978:ADC

[Har78] J. A. Hartigan. Asymptotic distributions for clustering criteria. Annals of Statistics, 6(1):117–131, January 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176344071.

Hirschler:1975:FMT

[HC75] Patrick Hirschler and Thomas M. Cover. A finite memory test of the irrationality of the parameter of a coin. Annals of Statistics, 3 (4):939–946, July 1975. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176343194.

Hecker:1976:CAN

[Hec76] H. Hecker. A characterization of the asymptotic normality of linear combinations of order statistics from the uniform distribution. Annals of Statistics, 4(6):1244–1246, November 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343656.

Hedayat:1975:NKV

[HF75] A. Hedayat and W. T. Federer. On the nonexistence of Knut Vik designs for all even orders. Annals of Statistics, 3(2):445–447, March 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343068.

Hasza:1979:EAP

[HF79] David P. Hasza and Wayne A. Fuller. Estimation for autoregressive processes with unit roots. Annals of Statistics, 7(5):1106–1120, September 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176344793.

Hasminskii:1975:LBM

[HI75] R. Z. Hasminskii and I. A. Ibragimov. On a lower bound for moments of point estimators. Annals of Statistics, 3(1):228–233, January 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343012.

Hill:1975:SGA

[Hil75] Bruce M. Hill. A simple general approach to inference about the tail of a distribution. Annals of Statistics, 3(5):1163–1174, September 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343247.

Hinkley:1979:PL

[Hin79] David Hinkley. Predictive likelihood. Annals of Statistics, 7(4): 718–728, July 1979. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.

aos/1176344723. See correction [Hin80].

Hinkley:1980:NCP

[Hin80] David V. Hinkley. Note: Correction to “Predictive Likelihood”. Annals of Statistics, 8(3):694, May 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345022. See [Hin79].

Hipp:1974:SSE

[Hip74] Christian Hipp. Sufficient statistics and exponential families. Annals of Statistics, 2(6):1283–1292, November 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176342879.

Hipp:1975:NPT

[Hip75] C. Hipp. Note on the paper “Transformation Groups and Sufficient Statistics” by J. Pfanzagl. Annals of Statistics, 3(2):478–482, March 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343075.

Hirakawa:1973:MOS

[Hir73] Kosaburo Hirakawa. Moments of order statistics. Annals of Statistics, 1(2):392–394, March 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176342380.

Hedayat:1974:RSB

[HJ74] A. Hedayat and P. W. M. John. Resistant and susceptible BIB designs. Annals of Statistics, 2(1):148–158, January 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176342620.

Hajek:1974:RDA

[HK74a] Jaroslav Hajek and George Kimeldorf. Regression designs in autoregressive stochastic processes. Annals of Statistics, 2(3):520– 527, May 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176342711.

Hettmansperger:1974:NSC

[HK74b] Thomas P. Hettmansperger and Lawrence A. Klimko. A note on the strong convergence of distributions. Annals of Statistics, 2(3): 597–598, May 1974. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176342722.

Hwang:1975:BEL

[HK75] T. Y. Hwang and J. H. Klotz. Bahadur efficiency of linear rank statistics for scale alternatives. Annals of Statistics, 3(4):947–954, July 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343195.

Hall:1977:WCP

[HL77] W. J. Hall and R. M. Loynes. Weak convergence of processes related to likelihood ratios. Annals of Statistics, 5(2):330–341, March 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343798.

Hedayat:1979:TMC

[HL79] A. Hedayat and Shuo-Yen Robert Li. The trade-off method in the construction of BIB designs with variable support sizes. Annals of Statistics, 7(6):1277–1287, November 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176344846.

Hipp:1976:BMA

[HM76] C. Hipp and R. Michel. On the Bernstein–von Mises approximation of posterior distributions. Annals of Statistics, 4(5):972–980, September 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343595.

Hoeffding:1973:CSL

[Hoe73] Wassily Hoeffding. On the centering of a simple linear rank statistic. Annals of Statistics, 1(1):54–66, January 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1193342381.

Hoeffding:1977:SIB

[Hoe77] Wassily Hoeffding. Some incomplete and boundedly complete families of distributions. Annals of Statistics, 5(2):278–291, March 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343795.

Hoke:1975:CPI

[Hok75] Albert T. Hoke. The characteristic polynomial of the information matrix for second-order models. Annals of Statistics, 3(3): 780–786, May 1975. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176343145.

Holland:1973:CST

[Hol73a] Paul W. Holland. Covariance stabilizing transformations. Annals of Statistics, 1(1):84–92, January 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1193342384.

Holm:1973:AMC

[Hol73b] Sture Holm. The asymptotic minimax character of sequential binomial and sign tests. Annals of Statistics, 1(6):1139–1148, November 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176342562.

Holst:1973:SLT

[Hol73c] Lars Holst. Some limit theorems with applications in sampling theory. Annals of Statistics, 1(4):644–658, July 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176342460.

Holm:1977:CAL

[Hol77] Sture Holm. On a conjecture about the limiting minimal efficiency of sequential tests. Annals of Statistics, 5(2):375–378, March 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343802.

Holst:1979:TCL

[Hol79] Lars Holst. Two conditional limit theorems with applications. Annals of Statistics, 7(3):551–557, May 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176344676.

Horn:1975:OCC

[Hor75] Susan D. Horn. On the optimality criterion in compound decision problems. Annals of Statistics, 3(3):767–772, May 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343143.

Hosoya:1979:HOE

[Hos79] Yuzo Hosoya. High-order efficiency in the estimation of linear processes. Annals of Statistics, 7(3):516–530, May 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176344673.

Hoyle:1975:EGF

[Hoy75] M. H. Hoyle. Estimating generating functions. Annals of Statistics, 3(6):1361–1363, November 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176343291.

Hasegawa:1974:EMS

[HP74] Minoru Hasegawa and Michael D. Perlman. On the existence of a minimal sufficient subfield. Annals of Statistics, 2(5):1049–1055, September 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/

1176342826. See correction [HP75].

Hasegawa:1975:NCE

[HP75] Minoru Hasegawa and Michael D. Perlman. Notes: Correction to “On the Existence of a Minimal Sufficient Subfield”. Annals of Statistics, 3(6):1371–1372, November 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176343294. See [HP74].

Hanson:1976:CCR

[HP76] D. L. Hanson and Gordon Pledger. Consistency in concave regression. Annals of Statistics, 4(6):1038–1050, November 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343640.

Hollander:1974:RWT

[HPL74] Myles Hollander, Gordon Pledger, and Pi-Erh Lin. Robustness of the Wilcoxon test to a certain dependency between samples. Annals of Statistics, 2(1):177–181, January 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176342623.

Hollander:1977:FDT

[HPS77] Myles Hollander, Frank Proschan, and Jayaram Sethuraman. Functions decreasing in transposition and their applications in ranking problems. Annals of Statistics, 5(4):722–733, July 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343895.

Hanson:1973:CMR

[HPW73] D. L. Hanson, Gordon Pledger, and F. T. Wright. On consistency in monotonic regression. Annals of Statistics, 1(3):401–421, May 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176342407.

Hedayat:1974:MAD

[HRF74] A. Hedayat, B. L. Raktoe, and W. T. Federer. On a measure of aliasing due to fitting an incomplete model. Annals of Statistics, 2 (4):650–660, July 1974. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176342754.

Hedayat:1975:FCT

[HRS75] A. Hedayat, D. Raghavarao, and E. Seiden. Further contributions to the theory of F -squares design. Annals of Statistics, 3(3):712– 716, May 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343134.

Huber:1973:MTN

[HS73] Peter J. Huber and Volker Strassen. Minimax tests and the Neyman–Pearson lemma for capacities. Annals of Statistics, 1(2): 251–263, March 1973. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.

aos/1176342363. See correction [HS74].

Huber:1974:NCM

[HS74] Peter J. Huber and Volker Strassen. Note: Correction to minimax tests and the Neyman–Pearson lemma for capacities. Annals of Statistics, 2(1):223–224, January 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176342630. See [HS73].

Heath:1978:FAP

[HS78] David Heath and William Sudderth. On finitely additive priors, coherence, and extended admissibility. Annals of Statistics, 6(2): 333–345, March 1978. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176344128.

Hooper:1979:DES

[HS79] Jeffrey H. Hooper and Thomas J. Santner. Design of experiments for selection from ordered families of distributions. Annals of Statistics, 7(3):615–643, May 1979. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176344683.

Hsieh:1979:EBE

[Hsi79a] H. K. Hsieh. Exact Bahadur efficiencies for tests of the multivariate linear hypothesis. Annals of Statistics, 7(6):1231–1245, November 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176344842.

Hsieh:1979:AOL

[Hsi79b] H. K. Hsieh. On asymptotic optimality of likelihood ratio tests for multivariate normal distributions. Annals of Statistics, 7(3):592– 598, May 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176344680.

Hsuan:1979:SBP

[Hsu79] Francis C. Hsuan. A stepwise Bayesian procedure. Annals of Statistics, 7(4):860–868, July 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176344735.

Hansen:1974:CLN

[HT74] Ole Havard Hansen and Erik N. Torgersen. Comparison of linear normal experiments. Annals of Statistics, 2(2):367–373, March 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176342672.

Hamada:1975:DBI

[HT75a] Noboru Hamada and Fumikazu Tamari. Duals of balanced incomplete block designs derived from an affine geometry. Annals of Statistics, 3(4):926–938, July 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176343193.

Hordijk:1975:MFI

[HT75b] Arie Hordijk and Henk Tijms. A modified form of the iterative method of dynamic programming. Annals of Statistics, 3(1):203– 208, January 1975. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176343008.

Huber:1973:RRA

[Hub73] Peter J. Huber. Robust regression: Asymptotics, conjectures and Monte Carlo. Annals of Statistics, 1(5):799–821, September 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176342503.

Huber:1974:FIS

[Hub74] Peter J. Huber. Fisher information and spline interpolation. Annals of Statistics, 2(5):1029–1033, September 1974. CODEN ASTSC7.

ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176342822.

Hudson:1978:NIE

[Hud78] H. M. Hudson. A natural identity for exponential families with applications in multiparameter estimation. Annals of Statistics, 6 (3):473–484, May 1978. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176344194.

Huskova:1977:RCS

[Hus77] Marie Huskova. The rate of convergence of simple linear rank statistics under hypothesis and alternatives. Annals of Statistics, 5(4): 658–670, July 1977. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176343890.

Hedayat:1978:HMT

[HW78] A. Hedayat and W. D. Wallis. Hadamard matrices and their applications. Annals of Statistics, 6(6):1184–1238, November 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176344370.

Hwang:1973:CSC

[Hwa73] F. K. Hwang. Constructions for some classes of neighbor designs. Annals of Statistics, 1(4):786–790, July 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176342476.

Pickands:1975:SIU

[III75] James Pickands III. Statistical inference using extreme order statistics. Annals of Statistics, 3(1):119–131, January 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343003.

Inglis:1979:ADR

[Ing79] James Inglis. Admissible decision rules for the compound decision problem: The two-action two-state case. Annals of Statistics, 7(5): 1127–1135, September 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176344795.

Isii:1978:BCE

[IT78] K. Isii and Y. Taga. Bound on the classification error for discriminating between multivariate populations with specified means and covariance matrices. Annals of Statistics, 6(1):132–141, January 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176344072.

Jacroux:1978:PPO

[Jac78] Michael A. Jacroux. On the properties of proper (M, S) optimal block designs. Annals of Statistics, 6(6):1302–1309, November 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176344375.

Jaeschke:1979:ADS

[Jae79] D. Jaeschke. The asymptotic distribution of the supremum of the standardized empirical distribution function on subintervals. Annals of Statistics, 7(1):108–115, January 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176344558.

James:1979:CFD

[Jam79] Ian R. James. Characterization of a family of distributions by the independence of size and shape variables. Annals of Statistics, 7 (4):869–881, July 1979. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176344736.

Jaquette:1973:MDP

[Jaq73] Stratton C. Jaquette. Markov decision processes with a new optimality criterion: Discrete time. Annals of Statistics, 1(3):496– 505, May 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176342415.

Jaquette:1975:MDP

[Jaq75] Stratton C. Jaquette. Markov decision processes with a new optimality criterion: Continuous time. Annals of Statistics, 3(2):547– 553, March 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343087.

Jensen:1974:JDF

[Jen74] D. R. Jensen. On the joint distribution of Friedman’s χ_{r}^{2} statistics. Annals of Statistics, 2(2):311–322, March 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176342665.

Jirina:1976:ANK

[Jir76] Miloslav Jirina. On the asymptotic normality of Kendall’s rank correlation statistic. Annals of Statistics, 4(1):214–215, January 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343354.

Johnson:1979:DRO

[JLL79] Richard A. Johnson, J. Ladalla, and S. T. Liu. Differential relations, in the original parameters, which determine the first two moments of the multiparameter exponential family. Annals of Statistics, 7(1): 232–235, January 1979. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176344569.

Jensen:1975:NTA

[JM75] D. R. Jensen and L. S. Mayer. Normal-theory approximations to tests for linear hypotheses. Annals of Statistics, 3(2):429–444, March 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343067.

Jensen:1977:SVR

[JM77] D. R. Jensen and L. S. Mayer. Some variational results and their applications in multiple inference. Annals of Statistics, 5(5):922– 931, September 1977. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176343948.

Jupp:1979:MLE

[JM79] P. E. Jupp and K. V. Mardia. Maximum likelihood estimators for the matrix von Mises–Fisher and Bingham distributions. Annals of Statistics, 7(3):599–606, May 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176344681.

Jogdeo:1977:API

[Jog77] Kumar Jogdeo. Association and probability inequalities. Annals of Statistics, 5(3):495–504, May 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176343846.

Jogdeo:1978:PBM

[Jog78] Kumar Jogdeo. On a probability bound of Marshall and Olkin. Annals of Statistics, 6(1):232–234, January 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176344082.

John:1973:FEC

[Joh73] J. A. John. Factorial experiments in cyclic designs. Annals of Statistics, 1(1):188–194, January 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1193342399.

Johnson:1974:ARI

[Joh74] Richard A. Johnson. Asymptotic results for inference procedures based on the r smallest observations. Annals of Statistics, 2(6): 1138–1151, November 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176342870.

John:1975:ID

[Joh75a] Peter W. M. John. Isomorphic L_{2}(7) designs. Annals of Statistics, 3 (3):728–729, May 1975. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176343136.

Johnson:1975:BVU

[Joh75b] Bruce McK. Johnson. Bounds on the variance of the U-statistic for symmetric distributions with shift alternatives. Annals of Statistics, 3(4):955–958, July 1975. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176343196.

John:1976:ISG

[Joh76a] Peter W. M. John. Inequalities for semiregular group divisible designs. Annals of Statistics, 4(5):956–959, September 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343592.

John:1976:RBI

[Joh76b] Peter W. M. John. Robustness of balanced incomplete block designs. Annals of Statistics, 4(5):960–962, September 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343593.

John:1978:NFP

[Joh78] Peter W. M. John. A note on a finite population plan. Annals of Statistics, 6(3):697–699, May 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176344213.

Joshi:1975:CBR

[Jos75a] V. M. Joshi. A conjecture of Berry regarding a Bernoulli twoarmed bandit. Annals of Statistics, 3(1):189–202, January 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343007. See

correction [Jos85].

Joshi:1975:MER

[Jos75b] V. M. Joshi. On the minimax estimation of a random probability with known first N moments. Annals of Statistics, 3(3):680– 687, May 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343130.

Joshi:1975:SAS

[Jos75c] V. M. Joshi. Strong admissibility of a set of confidence intervals for the mean of a finite population. Annals of Statistics, 3(2):483–488, March 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343076.

Joshi:1976:CIL

[Jos76a] V. M. Joshi. Confidence intervals for linear functions of the normal parameters. Annals of Statistics, 4(2):413–418, March 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343419.

Joshi:1976:ACR

[Jos76b] V. M. Joshi. On the attainment of the Cramér–Rao lower bound. Annals of Statistics, 4(5):998–1002, September 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343599.

Joshi:1977:NEF

[Jos77] V. M. Joshi. A note on estimators in finite populations. Annals of Statistics, 5(5):1051–1053, September 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176343960.

Joshi:1979:BSE

[Jos79a] V. M. Joshi. The best strategy for estimating the mean of a finite population. Annals of Statistics, 7(3):531–536, May 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176344674.

Joshi:1979:JAS

[Jos79b] V. M. Joshi. Joint admissibility of the sample means as estimators of the means of finite populations. Annals of Statistics, 7(5):995– 1002, September 1979. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176344783.

Joshi:1985:CCB

[Jos85] V. M. Joshi. Correction: “A Conjecture of Berry Regarding a Bernoulli Two-Armed Bandit”. Annals of Statistics, 13(3):1249, September 1985. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/

1176349670. See [Jos75a].

Jones:1978:BES

[JS78] Douglas H. Jones and J. Sethuraman. Bahadur efficiencies of the Student’s t-tests. Annals of Statistics, 6(3):559–566, May 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176344200.

Johnson:1974:ABB

[JT74] B. R. Johnson and D. R. Truax. Asymptotic behavior of Bayes tests and Bayes risk. Annals of Statistics, 2(2):278–294, March 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176342663.

Johnson:1978:ABB

[JT78] B. R. Johnson and D. R. Traux. Asymptotic behavior of Bayes procedures for testing simple hypotheses in multiparameter exponential families. Annals of Statistics, 6(2):346–361, March 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176344129.

Jureckova:1973:CLT

[Jur73] Jana Jureckova. Central limit theorem for Wilcoxon rank statistics process. Annals of Statistics, 1(6):1046–1060, November 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176342556.

Jureckova:1977:ARE

[Jur77] Jana Jureckova. Asymptotic relations of M-estimates and Restimates in linear regression model. Annals of Statistics, 5(3): 464–472, May 1977. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176343843.

Kadane:1971:MPO

[Kad71] Joseph B. Kadane. A moment problem for order statistics. Annals of Mathematical Statistics, 42(2):745–751, April 1971. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http:

//projecteuclid.org/euclid.aoms/1177693423. See correction

[Kad73].

Kadane:1973:CNC

[Kad73] Joseph B. Kadane. Correction note: Correction to “A Moment Problem for Order Statistics”. Annals of Statistics, 1(4):797, July 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/

1176342478. See [Kad71].

Kageyama:1973:RAR

[Kag73a] Sanpei Kageyama. On μ-resolvable and affine μ-resolvable balanced incomplete block designs. Annals of Statistics, 1(1):195–203, January 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1193342400.

Kageyama:1973:IBS

[Kag73b] Sanpei Kageyama. On the inequality for BIBDs with special parameters. Annals of Statistics, 1(1):204–207, January 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1193342401.

Kageyama:1974:RAC

[Kag74] Sanpei Kageyama. On the reduction of associate classes for the PBIB design of a certain generalized type. Annals of Statistics, 2 (6):1346–1350, November 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176342889.

Kageyama:1976:RBD

[Kag76] Sanpei Kageyama. Resolvability of block designs. Annals of Statistics, 4(3):655–661, May 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176343475.

Kallenberg:1974:NAE

[Kal74] Olav Kallenberg. A note on the asymptotic equivalence of sampling with and without replacement. Annals of Statistics, 2(4): 819–821, July 1974. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176342770.

Kao:1978:TEL

[Kao78] Chung-Siung Kao. On the time and the excess of linear boundary crossings of sample sums. Annals of Statistics, 6(1):191–199, January 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176344078.

Karlin:1974:ISS

[Kar74] Samuel Karlin. Inequalities for symmetric sampling plans I. Annals of Statistics, 2(6):1065–1094, November 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176342866.

Kariya:1977:RPT

[Kar77] Takeaki Kariya. A robustness property of the tests for serial correlation. Annals of Statistics, 5(6):1212–1220, November 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176344005.

Kariya:1978:GMP

[Kar78] Takeaki Kariya. The general MANOVA problem. Annals of Statistics, 6(1):200–214, January 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176344079.

Kadane:1978:SDP

[KC78] Joseph B. Kadane and David T. Chuang. Stable decision problems. Annals of Statistics, 6(5):1095–1110, September 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176344313.

Kariya:1977:RTS

[KE77] Takeaki Kariya and Morris L. Eaton. Robust tests for spherical symmetry. Annals of Statistics, 5(1):206–215, January 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343755.

Keiding:1975:MLE

[Kei75] Niels Keiding. Maximum likelihood estimation in the birth-and-death process. Annals of Statistics, 3(2):363–372, March 1975. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343062. See

correction [Kei78].

Keiding:1978:NCM

[Kei78] Niels Keiding. Note: Correction to “Maximum Likelihood Estimation in the Birth-and-Death Process”. Annals of Statistics, 6 (2):472, March 1978. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176344139. See [Kei75].

Kelley:1974:NBT

[Kel74] Thomas A. Kelley. A note on the Bernoulli two-armed bandit problem. Annals of Statistics, 2(5):1056–1062, September 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176342827.

Kelley:1977:SBE

[Kel77] Thomas A. Kelley. Sequential Bayes estimation of the difference between means. Annals of Statistics, 5(2):379–384, March 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343803.

Kushner:1973:EKA

[KG73] H. J. Kushner and T. Gavin. Extensions of Kesten’s adaptive stochastic approximation method. Annals of Statistics, 1(5):851– 861, September 1973. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176342506.

Korwar:1976:EBE

[KH76] Ramesh M. Korwar and Myles Hollander. Empirical Bayes estimation of a distribution function. Annals of Statistics, 4(3):581– 588, May 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343463.

Khatri:1962:CWI

[Kha62] C. G. Khatri. Conditions for Wishartness and independence of second degree polynomials in a normal vector. Annals of Mathematical Statistics, 33(3):1002–1007, September 1962. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http:

//projecteuclid.org/euclid.aoms/1177704467. See corrections

[Kha73c].

Khan:1973:SD

[Kha73a] Rasul A. Khan. On sequential distinguishability. Annals of Statistics, 1(5):838–850, September 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176342505.

Khan:1973:SPH

[Kha73b] Rasul A. Khan. On some properties of Hammersley’s estimator of an integer mean. Annals of Statistics, 1(4):756–762, July 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176342471.

Khatri:1973:NCC

[Kha73c] C. G. Khatri. Note: Corrections to “Conditions for Wishartness and Independence of Second Degree Polynomials in Normal Vector”.

Annals of Statistics, 1(5):991, September 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176342522. See [Kha62].

Khatri:1973:NCF

[Kha73d] C. G. Khatri. Note: Corrections to “Further Contributions to Wishartness and Independence of Second Degree Polynomials in Normal Vectors”. Annals of Statistics, 1(5):991, September 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176342523.

Killeen:1972:ETP

[KHS72] Timothy J. Killeen, Thomas P. Hettmansperger, and Gerald L. Sievers. An elementary theorem on the probability of large deviations. Annals of Mathematical Statistics, 43(1):181–192, February 1972. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic).

URL http://projecteuclid.org/euclid.aoms/1177692712. See

correction [KHS74].

Killeen:1974:NCE

[KHS74] Timothy J. Killeen, Thomas P. Hettmansperger, and Gerald Sievers. Notes: Correction to “An Elementary Theorem on the Probability of Large Deviations”. Annals of Statistics, 2(6):1357, November 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/

1176342891. See [KHS72].

Kiefer:1974:GET

[Kie74] J. Kiefer. General equivalence theory for optimum designs (approximate theory). Annals of Statistics, 2(5):849–879, September 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176342810.

Kiefer:1975:BBD

[Kie75] J. Kiefer. Balanced block designs and generalized Youden designs, I. Construction (patchwork). Annals of Statistics, 3(1):109–118, January 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343002.

Kiefer:1976:ACC

[Kie76] J. Kiefer. Admissibility of conditional confidence procedures. Annals of Statistics, 4(5):836–865, September 1976. CODEN ASTSC7.

ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176343584.

Kimelfeld:1974:EKN

[Kim74] Benjamin Kimelfeld. Estimating the kernels of nonlinear orthogonal polynomial functionals. Annals of Statistics, 2(2):353–358, March 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176342669.

Kowalczyk:1979:SMD

[KKMP79] T. Kowalczyk, A. Kowalski, A. Matuszewski, and E. Pleszczynska. Screening and monotonic dependence functions in the multivariate case. Annals of Statistics, 7(3):607–614, May 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176344682.

Kleyle:1975:ULP

[Kle75] Robert Kleyle. Upper and lower posterior probabilities for discrete distributions. Annals of Statistics, 3(2):504–511, March 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343080.

Klotz:1973:SIB

[Klo73] Jerome Klotz. Statistical inference in Bernoulli trials with dependence. Annals of Statistics, 1(2):373–379, March 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176342377.

Klotz:1978:TLC

[Klo78] Jerome H. Klotz. Testing a linear constraint for multinomial Cell frequencies and disease screening. Annals of Statistics, 6(4):904– 909, July 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176344263.

Klugman:1977:DRS

[Klu77] Stuart Klugman. Discounted and rapid subfair red-and-black. Annals of Statistics, 5(4):734–745, July 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176343896.

Kent:1979:CUD

[KMR79] J. T. Kent, K. V. Mardia, and J. S. Rao. A characterization of the uniform distribution on the circle. Annals of Statistics, 7(4): 882–889, July 1979. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176344737.

Klimko:1978:CLS

[KN78] Lawrence A. Klimko and Paul I. Nelson. On conditional least squares estimation for stochastic processes. Annals of Statistics, 6 (3):629–642, May 1978. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176344207.

Kohlberg:1974:RGA

[Koh74] Elon Kohlberg. Repeated games with absorbing states. Annals of Statistics, 2(4):724–738, July 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176342760.

Koul:1977:BRE

[Kou77] Hira L. Koul. Behavior of robust estimators in the regression model with dependent errors. Annals of Statistics, 5(4):681–699, July 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343892.

Koziol:1978:ESC

[Koz78] James A. Koziol. Exact slopes of certain multivariate tests of hypotheses. Annals of Statistics, 6(3):546–558, May 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176344199.

Kowalczyk:1977:MDF

[KP77] T. Kowalczyk and E. Pleszczynska. Monotonic dependence functions of bivariate distributions. Annals of Statistics, 5(6):1221–1227, November 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176344006.

Koziol:1977:AET

[KR77] James A. Koziol and Nancy Reid. On the asymptotic equivalence of two ranking methods for K-sample linear rank statistics. Annals of Statistics, 5(6):1099–1106, November 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176343998.

Krauth:1973:AUS

[Kra73] J. Krauth. An asymptotic UMP sign test in the presence of ties. Annals of Statistics, 1(1):166–169, January 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1193342394.

Kremer:1979:ALB

[Kre79] Erhard Kremer. Approximate and local Bahadur efficiency of linear rank tests in the two-sample problem. Annals of Statistics, 7(6): 1246–1255, November 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176344843.

Khatri:1974:AEN

[KS74] C. G. Khatri and M. S. Srivastava. Asymptotic expansions of the non-null distributions of likelihood ratio criteria for covariance matrices. Annals of Statistics, 2(1):109–117, January 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176342617.

Kiefer:1976:ODL

[KS76a] J. Kiefer and W. J. Studden. Optimal designs for large degree polynomial regression. Annals of Statistics, 4(6):1113–1123, November 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343646.

Koul:1976:PBS

[KS76b] Hira Lal Koul and R. G. Staudte. Power bounds for a Smirnov statistic in testing the hypothesis of symmetry. Annals of Statistics, 4(5):924–935, September 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176343589.

Kadane:1977:OSC

[KS77] Joseph B. Kadane and Herbert A. Simon. Optimal strategies for a class of constrained sequential problems. Annals of Statistics, 5(2): 237–255, March 1977. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.

aos/1176343791. See correction [KS83].

Kimeldorf:1978:MD

[KS78] George Kimeldorf and Allan R. Sampson. Monotone dependence. Annals of Statistics, 6(4):895–903, July 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176344262.

Kadane:1983:COS

[KS83] Joseph B. Kadane and Herbert A. Simon. Corrections: “Optimal Strategies for a Class of Constrained Sequential Problems”. Annals of Statistics, 11(1):346, March 1983. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176346087. See [KS77].

Kulldorff:1973:NOS

[Kul73] Gunnar Kulldorff. A note on the optimum spacing of sample quantiles from the six extreme value distributions. Annals of Statistics, 1 (3):562–567, May 1973. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176342425.

Kushner:1974:SAA

[Kus74] Harold J. Kushner. Stochastic approximation algorithms for constrained optimization problems. Annals of Statistics, 2(4):713– 723, July 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176342759.

Kohlberg:1974:RGI

[KZ74] Elon Kohlberg and Shmuel Zamir. Repeated games of incomplete information: The symmetric case. Annals of Statistics, 2(5):1040– 1041, September 1974. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176342824.

Lai:1973:OSS

[Lai73] Tze Leung Lai. Optimal stopping and sequential tests which minimize the maximum expected sample size. Annals of Statistics, 1 (4):659–673, July 1973. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176342461.

Lai:1974:CCB

[Lai74] Tze Leung Lai. Control charts based on weighted sums. Annals of Statistics, 2(1):134–147, January 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176342619.

Lai:1975:NFE

[Lai75a] Tze Leung Lai. A note on first exit times with applications to sequential analysis. Annals of Statistics, 3(4):999–1005, July 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343203.

Lai:1975:CSS

[Lai75b] Tze Leung Lai. On Chernoff–Savage statistics and sequential rank tests. Annals of Statistics, 3(4):825–845, July 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343185.

Lai:1975:TME

[Lai75c] Tze Leung Lai. Termination, moments and exponential boundedness of the stopping rule for certain invariant sequential probability ratio tests. Annals of Statistics, 3(3):581–598, May 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343124.

Lai:1976:CS

[Lai76] Tze Leung Lai. On confidence sequences. Annals of Statistics, 4(2): 265–280, March 1976. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176343406.

Lai:1977:POT

[Lai77] Tze Leung Lai. Power-one tests based on sample sums. Annals of Statistics, 5(5):866–880, September 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176343943.

Lai:1978:PES

[Lai78] Tze Leung Lai. Pitman efficiencies of sequential tests and uniform limit theorems in nonparametric statistics. Annals of Statistics, 6 (5):1027–1047, September 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176344308.

Lai:1979:STH

[Lai79] Tze Leung Lai. Sequential tests for hypergeometric distributions and finite populations. Annals of Statistics, 7(1):46–59, January 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176344554.

LaMotte:1977:CFG

[LaM77] Lynn Roy LaMotte. A canonical form for the general linear model. Annals of Statistics, 5(4):787–789, July 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176343901.

Landers:1974:MSS

[Lan74] D. Landers. Minimal sufficient statistics for families of product measures. Annals of Statistics, 2(6):1335–1339, November 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176342887.

Latta:1979:CRP

[Lat79] Robert B. Latta. Composition rules for probabilities from paired comparisons. Annals of Statistics, 7(2):349–371, March 1979. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176344619.

Laurent:1974:IRS

[Lau74] Andre G. Laurent. The intersection of random spheres and the noncentral radial error distribution for spherical models. Annals of Statistics, 2(1):182–189, January 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176342624.

Lawrence:1975:IOD

[Law75] M. J. Lawrence. Inequalities of s-ordered distributions. Annals of Statistics, 3(2):413–428, March 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176343066.

LeCam:1974:ICA

[Le 74] Lucien Le Cam. On the information contained in additional observations. Annals of Statistics, 2(4):630–649, July 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176342753.

LeCam:1979:RTC

[Le 79] L. Le Cam. A reduction theorem for certain sequential experiments. II. Annals of Statistics, 7(4):847–859, July 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176344734.

LeCam:1973:CEU

[LeC73] L. LeCam. Convergence of estimates under dimensionality restrictions. Annals of Statistics, 1(1):38–53, January 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1193342380.

Lehmann:1979:HWI

[Leh79] E. L. Lehmann. Hsu’s work on inference. Annals of Statistics, 7 (3):471–473, May 1979. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176344667.

Lembersky:1974:MRO

[Lem74] Mark R. Lembersky. On maximal rewards and ε-optimal policies in continuous time Markov decision chains. Annals of Statistics, 2(1): 159–169, January 1974. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176342621.

Lee:1977:OST

[LG77] Austin F. S. Lee and John Gurland. One-sample t-test when sampling from a mixture of normal distributions. Annals of Statistics, 5 (4):803–807, July 1977. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176343904.

Liggett:1977:TSC

[Lig77] Walter S. Liggett. A test for serial correlation in multivariate data. Annals of Statistics, 5(2):408–413, March 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343808.

Lii:1978:GMS

[Lii78] Keh-Shin Lii. A global measure of a spline density estimate. Annals of Statistics, 6(5):1138–1148, September 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176344316.

Ling:1974:TMP

[Lin74] Kang Ling. On two-move prediction games. Annals of Statistics, 2 (5):988–999, September 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176342819.

Lin:1975:RCE

[Lin75] Pi-Erh Lin. Rates of convergence in empirical Bayes estimation problems: Continuous case. Annals of Statistics, 3(1):155–164, January 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343005.

Lindley:1976:CUF

[Lin76] D. V. Lindley. A class of utility functions. Annals of Statistics, 4 (1):1–10, January 1976. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176343343.

Land:1973:NTS

[LJJ73] C. E. Land, B. R. Johnson, and V. M. Joshi. A note on twosided confidence intervals for linear functions of the normal mean and variance. Annals of Statistics, 1(5):940–943, September 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176342514.

Ljung:1978:SCS

[Lju78] Lennart Ljung. Strong convergence of a stochastic approximation algorithm. Annals of Statistics, 6(3):680–696, May 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176344212.

Lecam:1974:JNO

[LL74] L. Lecam and E. L. Lehmann. J. Neyman: On the occasion of his 80th birthday. Annals of Statistics, 2(3):??, May 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176342703.

Lorden:1973:OET

[Lor73] Gary Lorden. Open-ended tests for Koopman–Darmois families. Annals of Statistics, 1(4):633–643, July 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176342459.

Lorden:1976:SMK

[Lor76] Gary Lorden. 2-SPRT’S and the modified Kiefer–Weiss problem of minimizing an expected sample size. Annals of Statistics, 4(2): 281–291, March 1976. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176343407.

Lorden:1977:NOS

[Lor77] Gary Lorden. Nearly-optimal sequential tests for finitely many parameter values. Annals of Statistics, 5(1):1–21, January 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343737.

Landers:1973:SI

[LR73] D. Landers and L. Rogge. On sufficiency and invariance. Annals of Statistics, 1(3):543–544, May 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176342420.

Lanke:1974:HSS

[LR74] Jan Lanke and M. K. Ramakrishnan. Hyperadmissibility in survey sampling. Annals of Statistics, 2(1):205–215, January 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176342627.

Levin:1977:CML

[LR77] Bruce Levin and James Reeds. Compound multinomial likelihood functions are unimodal: Proof of a conjecture of I. J. Good. Annals of Statistics, 5(1):79–87, January 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176343741.

Lai:1979:ADS

[LR79] T. L. Lai and Herbert Robbins. Adaptive design and stochastic approximation. Annals of Statistics, 7(6):1196–1221, November 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176344840.

Lai:1977:NRT

[LS77] T. L. Lai and D. Siegmund. A nonlinear renewal theory with applications to sequential analysis I. Annals of Statistics, 5(5):946–954, September 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343950.

Lane:1978:DMS

[LS78] David A. Lane and William D. Sudderth. Diffuse models for sampling and predictive inference. Annals of Statistics, 6(6):1318–1336, November 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176344377.

Lai:1979:NRT

[LS79a] T. L. Lai and D. Siegmund. A nonlinear renewal theory with applications to sequential analysis II. Annals of Statistics, 7(1):60–76, January 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176344555.

Lehmann:1979:OSL

[LS79b] E. L. Lehmann and Juliet Popper Shaffer. Optimum significance levels for multistage comparison procedures. Annals of Statistics, 7(1): 27–45, January 1979. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176344553.

Lin:1973:GBM

[LT73] Pi-Erh Lin and Hui-Liang Tsai. Generalized Bayes minimax estimators of the multivariate normal mean with unknown covariance matrix. Annals of Statistics, 1(1):142–145, January 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1193342390.

M:1975:ITU

[M.75] Evarist Gine M. Invariant tests for uniformity on compact Riemannian manifolds based on Sobolev norms. Annals of Statistics, 3(6): 1243–1266, November 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176343283.

Malik:1973:MLD

[MA73] Henrick J. Malik and Bovas Abraham. Multivariate logistic distributions. Annals of Statistics, 1(3):588–590, May 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176342430.

Maag:1973:AED

[Maa73] Urs R. Maag. The asymptotic expansion of the distribution of the goodness of fit statistic V_{NN}. Annals of Statistics, 1(6):1185–1188, November 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176342566.

MacNeill:1974:TCP

[Mac74a] Ian B. MacNeill. Tests for change of parameter at unknown times and distributions of some related functionals on Brownian motion. Annals of Statistics, 2(5):950–962, September 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176342816.

MacRae:1974:MDA

[Mac74b] Elizabeth Chase MacRae. Matrix derivatives with an application to an adaptive linear decision problem. Annals of Statistics, 2(2): 337–346, March 1974. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176342667.

MacNeill:1978:PSP

[Mac78] Ian B. MacNeill. Properties of sequences of partial sums of polynomial regression residuals with applications to tests for change of regression at unknown times. Annals of Statistics, 6(2):422–433, March 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176344133.

Makowski:1973:LIL

[Mak73] Gary G. Makowski. Laws of the iterated logarithm for permuted random variables and regression applications. Annals of Statistics, 1(5):872–887, September 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176342508.

Makowski:1974:RCD

[Mak74] Gary Makowski. A rate of convergence of a distribution connected with integral regression function estimation. Annals of Statistics, 2 (4):829–832, July 1974. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176342772.

Makani:1977:PA

[Mak77] Sherali Mavjibhai Makani. A paradox in admissibility. Annals of Statistics, 5(3):544–546, May 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176343853.

Maltz:1974:EDD

[Mal74] Carl Maltz. Estimation of the k th derivative of a distribution function. Annals of Statistics, 2(2):359–361, March 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176342670.

Manoussakis:1977:RSP

[Man77] E. Manoussakis. Repeated sampling with partial replacement of units. Annals of Statistics, 5(4):795–802, July 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343903.

Marsaglia:1974:APRb

[Mar74a] George Marsaglia. Acknowledgement of priority to: “Random variables with independent binary digits” (Ann. Math. Statist. 42 (1971), 1922–1929). Annals of Statistics, 2(4):848, 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176342776.

Marsaglia:1974:NAP

[Mar74b] George Marsaglia. Note: Acknowledgment of priority to “Random Variables with Independent Binary Digits”. Annals of Statistics, 2(4):848, July 1974. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176342776.

Maronna:1976:REM

[Mar76] Ricardo Antonio Maronna. Robust M-estimators of multivariate location and scatter. Annals of Statistics, 4(1):51–67, January 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343347.

Maurer:1975:MET

[Mau75] Willi Maurer. On most effective tournament plans with fewer games than competitors. Annals of Statistics, 3(3):717–727, May 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343135.

Moore:1978:REO

[MB78] Terry Moore and Richard J. Brook. Risk estimate optimality of James–Stein estimators. Annals of Statistics, 6(4):917–919, July 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176344265.

Mann:1979:BST

[MB79] J. E. Mann and T. L. Bratcher. On Bayes sequential tests regarding two Koopman–Darmois distributions: Bounds on the maximum sample size. Annals of Statistics, 7(6):1339–1344, November 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176344852.

Mead:1975:PAV

[MBH75] Ronald Mead, T. A. Bancroft, and Chien-Pai Han. Power of analysis of variance test procedures for incompletely specified fixed models. Annals of Statistics, 3(4):797–808, July 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343183.

Muirhead:1975:AEJ

[MC75] R. J. Muirhead and Y. Chikuse. Asymptotic expansions for the joint and marginal distributions of the latent roots of the covariance matrix. Annals of Statistics, 3(4):1011–1017, July 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343205.

McDonald:1975:TGL

[McD75] Lyman McDonald. Tests for the general linear hypothesis under the multiple design multivariate linear model. Annals of Statistics, 3(2): 461–466, March 1975. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176343072.

Meeden:1976:SPL

[Mee76] Glen Meeden. A special property of linear estimates of the normal mean. Annals of Statistics, 4(3):649–650, May 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343473.

Mehran:1975:RBU

[Meh75] Farhad Mehran. Relationships between the UMVU estimators of the mean and median of a function of a normal distribution. Annals of Statistics, 3(2):457–460, March 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176343071.

Mentz:1977:EFO

[Men77] Raul Pedro Mentz. Estimation in the first order moving average model based on sample autocorrelations. Annals of Statistics, 5(6): 1250–1257, November 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176344012.

Meyer:1977:BED

[Mey77a] Terry G. Meyer. Bounds for estimation of density functions and their derivatives. Annals of Statistics, 5(1):136–142, January 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343746.

Meyer:1977:FSR

[Mey77b] Terry G. Meyer. On fixed or scaled radii confidence sets: The fixed sample size case. Annals of Statistics, 5(1):65–78, January 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343740.

Meeden:1977:ABP

[MI77] Glen Meeden and Dean Isaacson. Approximate behavior of the posterior distribution for a large observation. Annals of Statistics, 5(5):899–908, September 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176343946.

Michel:1978:AES

[Mic78] R. Michel. On the asymptotic efficiency of strongly asymptotically median unbiased estimators. Annals of Statistics, 6(4):920– 922, July 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176344266.

Michel:1979:AEC

[Mic79] R. Michel. On the asymptotic efficiency of conditional tests for exponential families. Annals of Statistics, 7(6):1256–1263, November 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176344844.

Miescke:1979:ISP

[Mie79] Klaus J. Miescke. Identification and selection procedures based on tests. Annals of Statistics, 7(1):207–219, January 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176344566.

Miller:1974:JR

[Mil74a] Rupert G. Miller. The jackknife — a review. Biometrika, 61(1):1–15, April 1974. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510

(electronic). URL http://www.jstor.org/stable/2334280. See

[Que49, Que56, Efr79].

Miller:1974:UJ

[Mil74b] Rupert G. Miller. An unbalanced jackknife. Annals of Statistics, 2 (5):880–891, September 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176342811.

Miller:1977:NIM

[Mil77a] Douglas R. Miller. A note on independence of multivariate lifetimes in competing risks models. Annals of Statistics, 5(3):576– 579, May 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343859.

Miller:1977:APM

[Mil77b] John J. Miller. Asymptotic properties of maximum likelihood estimates in the mixed model of the analysis of variance. Annals of Statistics, 5(4):746–762, July 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176343897.

Mehrotra:1976:ASA

[MJ76] K. G. Mehrotra and Richard A. Johnson. Asymptotic sufficiency and asymptotically most powerful tests for the two sample censored situation. Annals of Statistics, 4(3):589–596, May 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343464.

Maurer:1976:MIE

[MM76] Willi Maurer and Barry H. Margolin. The multivariate inclusionexclusion formula and order statistics from dependent variates. Annals of Statistics, 4(6):1190–1199, November 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343650.

Magnus:1979:CMS

[MN79] Jan R. Magnus and Heinz Neudecker. The commutation matrix: Some properties and applications. Annals of Statistics, 7(2):381– 394, March 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176344621.

Marshall:1974:MMD

[MO74] Albert W. Marshall and Ingram Olkin. Majorization in multivariate distributions. Annals of Statistics, 2(6):1189–1200, November 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176342873.

Mohanty:1977:NAP

[Moh77] S. G. Mohanty. Note: Acknowledgement of priority to “A Short Proof of Steck’s Result on Two-Sample Smirnov Statistics”. Annals of Statistics, 5(2):429, March 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176343811.

Moore:1974:BRP

[Moo74] Marc Moore. Bayesian reconstructions of m, n-patterns. Annals of Statistics, 2(6):1226–1237, November 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176342875.

Morris:1975:CLT

[Mor75] Carl Morris. Central limit theorems for multinomial sums. Annals of Statistics, 3(1):165–188, January 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176343006.

Most:1975:RBI

[Mos75] Bernard M. Most. Resistance of balanced incomplete block designs. Annals of Statistics, 3(5):1149–1162, September 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343246.

McEliece:1973:HCC

[MP73] Robert J. McEliece and Edward C. Posner. Hiding and covering in a compact metric space. Annals of Statistics, 1(4):729–739, July 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176342467.

Mehra:1975:FOS

[MR75] K. L. Mehra and M. Sudhakara Rao. On functions of order statistics for mixing processes. Annals of Statistics, 3(4):874–883, July 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343188.

Moore:1975:ULS

[MS75] David S. Moore and M. C. Spruill. Unified large-sample theory of general chi-squared statistics for tests of fit. Annals of Statistics, 3 (3):599–616, May 1975. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176343125.

Mucci:1973:DEO

[Muc73] Anthony G. Mucci. Differential equations and optimal choice problems. Annals of Statistics, 1(1):104–113, January 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1193342386.

Muirhead:1978:LRM

[Mui78] Robb J. Muirhead. Latent roots and matrix variates: A review of some asymptotic results. Annals of Statistics, 6(1):5–33, January 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176344063.

Morf:1978:CCP

[MVK78] M. Morf, A. Vieira, and T. Kailath. Covariance characterization by partial autocorrelation matrices. Annals of Statistics, 6(3):643– 648, May 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176344208.

Maritz:1977:LEB

[MWS77] J. S. Maritz, Margaret Wu, and R. G. Stuadte. A location estimator based on a U-statistic. Annals of Statistics, 5(4):779–786, July 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343900.

Moore:1977:CPN

[MY77] David S. Moore and James W. Yackel. Consistency properties of nearest neighbor density function estimators. Annals of Statistics, 5(1):143–154, January 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176343747.

Nagao:1973:STC

[Nag73] Hisao Nagao. On some test criteria for covariance matrix. Annals of Statistics, 1(4):700–709, July 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176342464.

Neuhaus:1975:CRE

[Neu75] Georg Neuhaus. Convergence of the reduced empirical process for non-I.I.D. random vectors. Annals of Statistics, 3(2):528–531, March 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343084.

Nguyen:1979:DEC

[Ngu79] Hung T. Nguyen. Density estimation in a continuous-time stationary Markov process. Annals of Statistics, 7(2):341–348, March 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176344618.

Norwood:1977:ECM

[NH77] Thomas E. Norwood and Klaus Hinkelmann. Estimating the common mean of several normal populations. Annals of Statistics, 5(5): 1047–1050, September 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176343959.

Nigam:1970:BDM

[Nig70] A. K. Nigam. Block designs for mixture experiments. Annals of Mathematical Statistics, 41(6):1861–1869, December 1970. CODEN AASTAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL

http://projecteuclid.org/euclid.aoms/1177696687. See cor-

rection [Nig76].

Nigam:1976:NCB

[Nig76] A. K. Nigam. Notes: Correction to “Block Designs for Mixture Experiments”. Annals of Statistics, 4(6):1294–1295, November 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343663. See

[Nig70].

Nordbrock:1974:RDC

[Nor74] Earl Nordbrock. A rational decision criterion, the iterated minimax regret criterion. Annals of Statistics, 2(4):775–786, July 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176342764.

Nevius:1977:SFS

[NPS77] S. E. Nevius, F. Proschan, and J. Sethuraman. Schur functions in statistics II. stochastic majorization. Annals of Statistics, 5(2): 263–273, March 1977. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176343793.

OBryan:1976:SEB

[O’B76] Thomas E. O’Bryan. Some empirical Bayes results in the case of component problems with varying sample sizes for discrete exponential families. Annals of Statistics, 4(6):1290–1293, November 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343661.

Obenchain:1978:GOR

[Obe78] R. L. Obenchain. Good and optimal ridge estimators. Annals of Statistics, 6(5):1111–1121, September 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176344314.

OReilly:1980:CCS

[OD80] Federico J. O’Reilly and Joaquin Diaz. Correction to “On a Criterion for Simultaneous Extrapolation in Nonfull Rank Normal Regression”. Annals of Statistics, 8(5):1177–1178, September 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176345158. See

[O’R76].

Okamoto:1973:DEQ

[Oka73] Masashi Okamoto. Distinctness of the eigenvalues of a quadratic form in a multivariate sample. Annals of Statistics, 1(4):763– 765, July 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176342472.

OReilly:1973:CPI

[OQ73] Federico J. O’Reilly and C. P. Quesenberry. The conditional probability integral transformation and applications to obtain composite chi-square goodness-of-fit tests. Annals of Statistics, 1(1):74–83, January 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1193342383.

OReilly:1975:CEN

[O’R75] Federico J. O’Reilly. On a criterion for extrapolation in normal regression. Annals of Statistics, 3(1):219–222, January 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343010.

OReilly:1976:CSE

[O’R76] Federico J. O’Reilly. On a criterion for simultaneous extrapolation in nonfull rank normal regression. Annals of Statistics, 4(3):625– 628, May 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/

1176343469. See correction [OD80].

Orkin:1974:SPG

[Ork74] Michael Orkin. On stationary policies — the general case. Annals of Statistics, 2(1):219–222, January 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176342629.

Olkin:1979:AME

[OS79] Ingram Olkin and Milton Sobel. Admissible and minimax estimation for the multinomial distribution and for K independent binomial distributions. Annals of Statistics, 7(2):284–290, March 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176344613.

Olsen:1976:IQU

[OSB76] Anthony Olsen, Justus Seely, and David Birkes. Invariant quadratic unbiased estimation for two variance components. Annals of Statistics, 4(5):878–890, September 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176343586.

Oden:1975:AFP

[OW75] Anders Oden and Hans Wedel. Arguments for Fisher’s permutation test. Annals of Statistics, 3(2):518–520, March 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343082.

Pagano:1974:EMA

[Pag74] Marcello Pagano. Estimation of models of autoregressive signal plus white noise. Annals of Statistics, 2(1):99–108, January 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176342616.

Pagano:1978:PMA

[Pag78] Marcello Pagano. On periodic and multiple autoregressions. Annals of Statistics, 6(6):1310–1317, November 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176344376.

Park:1973:PSD

[Par73] C. J. Park. The power series distribution with unknown truncation parameter. Annals of Statistics, 1(2):395–399, March 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176342381.

Parsons:1979:NED

[Par79] Van L. Parsons. A note on the exact distribution of a nonparametric test statistic for ordered alternatives. Annals of Statistics, 7(2): 454–458, March 1979. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176344627.

Pathak:1976:UEF

[Pat76] P. K. Pathak. Unbiased estimation in fixed cost sequential sampling schemes. Annals of Statistics, 4(5):1012–1017, September 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343601.

Pazman:1974:CTT

[Paz74] A. Pazman. A convergence theorem in the theory of D-optimum experimental designs. Annals of Statistics, 2(1):216–218, January 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176342628.

Pazman:1978:COD

[Paz78] Andrej Pazman. Computation of the optimum designs under singular information matrices. Annals of Statistics, 6(2):465–467, March 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176344137.

Pham-Dinh:1978:EPA

[PD78] Tuan Pham-Dinh. Estimation of parameters in the ARMA model when the characteristic polynomial of the MA operator has a unit zero. Annals of Statistics, 6(6):1369–1389, November 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176344382.

Perng:1977:SET

[Per77] S.-S. Perng. Some extensions of a theorem of Stein on cumulative sums. Annals of Statistics, 5(1):98–109, January 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343743.

Perng:1978:EBS

[Per78a] S.-S. Perng. Exponentially bounded stopping times of invariant SPRT’s in general linear models: Finite mgf case. Annals of Statistics, 6(1):85–91, January 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176344067.

Perng:1978:RCE

[Per78b] S.-S. Perng. Rate of convergence of estimators based on sample mean. Annals of Statistics, 6(5):1048–1056, September 1978. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176344309.

Paik:1974:ANW

[PF74] U. B. Paik and W. T. Federer. Analysis of nonorthogonal n-way classifications. Annals of Statistics, 2(5):1000–1021, September 1974.

CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176342820.

Pfanzagl:1973:AER

[Pfa73] J. Pfanzagl. Asymptotic expansions related to minimum contrast estimators. Annals of Statistics, 1(6):993–1026, November 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176342554. See

corrections [Pfa74].

Pfanzagl:1974:NCA

[Pfa74] J. Pfanzagl. Notes: Corrections to “Asymptotic Expansions Related to Minimum Contrast Estimators”. Annals of Statistics, 2(6):1357– 1358, November 1974. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176342892. See [Pfa73].

Pfanzagl:1979:OMU

[Pfa79] J. Pfanzagl. On optimal median unbiased estimators in the presence of nuisance parameters. Annals of Statistics, 7(1):187–193, January 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176344563.

Phadia:1973:MEC

[Pha73] E. G. Phadia. Minimax estimation of a cumulative distribution function. Annals of Statistics, 1(6):1149–1157, November 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176342563.

Pierce:1973:SDF

[Pie73] Donald A. Pierce. On some difficulties in a frequency theory of inference. Annals of Statistics, 1(2):241–250, March 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176342362.

Pierce:1975:NTF

[Pie75] David A. Pierce. Noninvertible transfer functions and their forecasts. Annals of Statistics, 3(6):1354–1360, November 1975. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343290.

Pierce:1979:SEE

[Pie79] David A. Pierce. Signal extraction error in nonstationary time series. Annals of Statistics, 7(6):1303–1320, November 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176344848.

Pillai:1975:DCR

[Pil75] K. C. S. Pillai. The distribution of the characteristic roots of S_{1}S_{2}^{−1} under violations. Annals of Statistics, 3(3):773–779, May 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343144.

Pirie:1974:CRT

[Pir74] Walter R. Pirie. Comparing rank tests for ordered alternatives in randomized blocks. Annals of Statistics, 2(2):374–382, March 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176342673. See

correction [Pir75].

Pirie:1975:NCC

[Pir75] Walter R. Pirie. Notes: Correction to “Comparing Rank Tests for Ordered Alternatives in Randomized Blocks”. Annals of Statistics, 3(3):796, May 1975. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176343149. See [Pir74].

Peele:1977:PFM

[PK77] Lawrence Peele and George Kimeldorf. Prediction functions and mean-estimation functions for a time series. Annals of Statistics, 5 (4):709–721, July 1977. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176343894.

Peele:1979:TSP

[PK79] Lawrence Peele and George Kimeldorf. Time series prediction functions based on imprecise observations. Annals of Statistics, 7(4): 801–811, July 1979. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176344730.

Pledger:1976:CIR

[Ple76] Gordon Pledger. Consistency in integral regression estimation with a triangular array of observation points. Annals of Statistics, 4(1):

234–236, January 1976. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176343358.

Pollak:1973:ED

[Pol73] Moshe Pollak. On equal distributions. Annals of Statistics, 1(1): 180–182, January 1973. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1193342397.

Pollak:1978:OAO

[Pol78] Moshe Pollak. Optimality and almost optimality of mixture stopping rules. Annals of Statistics, 6(4):910–916, July 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176344264.

Portnoy:1975:ABI

[Por75] Stephen L. Portnoy. Admissibility of the best invariant estimator of one co-ordinate of a location vector. Annals of Statistics, 3(2): 448–450, March 1975. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176343069.

Portnoy:1977:AEM

[Por77a] Stephen Portnoy. Asymptotic efficiency of minimum variance unbiased estimators. Annals of Statistics, 5(3):522–529, May 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343849.

Portnoy:1977:RED

[Por77b] Stephen L. Portnoy. Robust estimation in dependent situations. Annals of Statistics, 5(1):22–43, January 1977. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343738. See

remarks [Por79].

Portnoy:1979:FRR

[Por79] Stephen L. Portnoy. Further remarks on robust estimation in dependent situations. Annals of Statistics, 7(1):224–231, January 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176344568. See

[Por77b].

Potthoff:1974:NPT

[Pot74] Richard F. Potthoff. A non-parametric test of whether two simple regression lines are parallel. Annals of Statistics, 2(2):295–310, March 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176342664.

Philippou:1973:ADL

[PR73] A. N. Philippou and G. G. Roussas. Asymptotic distribution of the likelihood function in the independent not identically distributed case. Annals of Statistics, 1(3):454–471, May 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176342411.

Paranjape:1975:SCD

[PR75] S. R. Paranjape and Herman Rubin. Special case of the distribution of the median. Annals of Statistics, 3(1):251–256, January 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343016.

Pratt:1976:FER

[Pra76] John W. Pratt. F. Y. Edgeworth and R. A. Fisher on the efficiency of maximum likelihood estimation. Annals of Statistics, 4(3):501– 514, May 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343457.

Prentice:1978:ITU

[Pre78] M. J. Prentice. On invariant tests of uniformity for directions and orientations. Annals of Statistics, 6(1):169–176, January 1978. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176344075. See

correction [Pre79].

Prentice:1979:NCI

[Pre79] M. J. Prentice. Note: Correction to “On Invariant Tests of Uniformity for Directions and Orientations”. Annals of Statistics, 7(4): 926, July 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/

1176344744. See [Pre78].

Puri:1973:NAD

[PS73] M. L. Puri and P. K. Sen. A note on asymptotically distribution free tests for subhypotheses in multiple linear regression. Annals of Statistics, 1(3):553–556, May 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176342423.

Port:1974:FIP

[PS74] Sidney C. Port and Charles J. Stone. Fisher information and the Pitman estimator of a location parameter. Annals of Statistics, 2(2): 225–247, March 1974. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176342660.

Pillai:1975:ERS

[PS75a] K. C. S. Pillai and Sudjana. Exact robustness studies of tests of two multivariate hypotheses based on four criteria and their distribution problems under violations. Annals of Statistics, 3(3):617– 636, May 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343126.

Pollak:1975:AES

[PS75b] M. Pollak and D. Siegmund. Approximations to the expected sample size of certain sequential tests. Annals of Statistics, 3(6):1267–1282, November 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343284.

Proschan:1977:SFS

[PS77] F. Proschan and J. Sethuraman. Schur functions in statistics I. The preservation theorem. Annals of Statistics, 5(2):256–262, March 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343792.

Perng:1974:SSI

[PT74] S. K. Perng and Yung Liang Tong. A sequential solution to the inverse linear regression problem. Annals of Statistics, 2(3):535– 539, May 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176342713.

Perng:1977:OAO

[PT77] S. K. Perng and Y. L. Tong. Optimal allocation of observations in inverse linear regression. Annals of Statistics, 5(1):191–196, January 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343753.

Praskova-Vizkova:1976:AEL

[PV76] Zuzana Praskova-Vizkova. Asymptotic expansion and a local limit theorem for a function of the Kendall rank correlation coefficient. Annals of Statistics, 4(3):597–606, May 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176343465.

Pfanzagl:1973:NOS

[PW73] J. Pfanzagl and J. Weber. Nonlinear operators and sufficiency. Annals of Statistics, 1(1):183–187, January 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1193342398.

Perlman:1975:NSC

[PW75] Michael D. Perlman and Michael J. Wichura. A note on substitution in conditional distribution. Annals of Statistics, 3(5):1175–1179, September 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343248.

Quenouille:1949:ATC

[Que49] M. H. Quenouille. Approximate tests of correlation in time-series. Journal of the Royal Statistical Society. Series B (Methodological), 11(1):68–84, ???? 1949. CODEN JSTBAJ. ISSN 0035-9246. URL

http://www.jstor.org/stable/2983696. See extension in [Que56,

Tuk58], and reviews [Mil74a, Efr79].

Quenouille:1956:NBE

[Que56] M. H. Quenouille. Notes on bias in estimation. Biometrika, 43 (3–4):353–360, December 1956. CODEN BIOKAX. ISSN 0006-3444 (print), 1464-3510 (electronic). URL http://www.jstor.org/ stable/2332914. This extends earlier work [Que49, Tuk58] on what is now known as the Quenouille–Tukey jacknife technique. See reviews [Mil74a, Efr79].

Raghavachari:1973:LDK

[Rag73a] M. Raghavachari. Limiting distributions of Kolmogorov–Smirnov type statistics under the alternative. Annals of Statistics, 1(1):67– 73, January 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1193342382.

Raghavarao:1973:MDC

[Rag73b] Damaraju Raghavarao. Method of differences in the construction of L_{2}(s) designs. Annals of Statistics, 1(3):591–592, May 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176342431.

Ramakrishnan:1973:APA

[Ram73] M. K. Ramakrishnan. An alternative proof of the admissibility of the Horvitz–Thompson estimator. Annals of Statistics, 1(3):577– 579, May 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176342427.

Ramsey:1974:CPA

[Ram74] Fred L. Ramsey. Characterization of the partial autocorrelation function. Annals of Statistics, 2(6):1296–1301, November 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176342881.

Ramakrishnan:1975:COS

[Ram75] M. K. Ramakrishnan. Choice of an optimum sampling strategy-1. Annals of Statistics, 3(3):669–679, May 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176343129.

Rao:1976:CPD

[Rao76a] C. Radhakrishna Rao. Characterization of prior distributions and solution to a compound decision problem. Annals of Statistics, 4 (5):823–835, September 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176343583.

Rao:1976:EPL

[Rao76b] C. Radhakrishna Rao. Estimation of parameters in a linear model. Annals of Statistics, 4(6):1023–1037, November 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343639. See cor-

rection [Rao79].

Rao:1978:ADE

[Rao78] M. M. Rao. Asymptotic distribution of an estimator of the boundary parameter of an unstable process. Annals of Statistics, 6(1):185– 190, January 1978. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.

aos/1176344077. See correction [Rao80].

Rao:1979:CEP

[Rao79] C. Radhakrishna Rao. Correction to “Estimation of Parameters in a Linear Model”. Annals of Statistics, 7(3):696, May 1979. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176344691. See

[Rao76b].

Rao:1980:CAD

[Rao80] M. M. Rao. Correction to “Asymptotic Distribution of an Estimator of the Boundary Parameter of an Unstable Process”. Annals of Statistics, 8(6):1403, November 1980. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345217. See [Rao78].

Ray:1974:MTA

[Ray74] Rose M. Ray. Maxmin C(α) tests against two-sided alternatives. Annals of Statistics, 2(6):1175–1188, November 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176342872.

Rissanen:1979:SCM

[RC79] J. Rissanen and P. E. Caines. The strong consistency of maximum likelihood estimators for ARMA processes. Annals of Statistics, 7(2):297–315, March 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176344615.

Rechtschaffen:1975:WCE

[Rec75] R. Rechtschaffen. Weak convergence of the empiric process for independent random variables. Annals of Statistics, 3(3):787–792, May 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343146.

Reeds:1978:JML

[Ree78] James A. Reeds. Jackknifing maximum likelihood estimates. Annals of Statistics, 6(4):727–739, July 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176344248.

Reiss:1973:MCM

[Rei73] Rolf-Dieter Reiss. On the measurability and consistency of maximum likelihood estimates for unimodal densities. Annals of Statistics, 1(5):888–901, September 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176342509.

Reynolds:1975:SSR

[Rey75] Marion R. Reynolds. A sequential signed-rank test for symmetry. Annals of Statistics, 3(2):382–400, March 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343064.

Raktoe:1973:BOS

[RF73] B. L. Raktoe and W. T. Federer. Balanced optimal saturated main effect plans of the 2^{n} factorial and their relation to (v, k, λ) configurations. Annals of Statistics, 1(5):924–932, September 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176342512.

Raghavarao:1975:CTW

[RF75] D. Raghavarao and W. T. Federer. On connectedness in two-way elimination of heterogeneity designs. Annals of Statistics, 3(3):730– 735, May 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343137.

Rincon-Gallardo:1979:CPI

[RGQO79] S. Rincon-Gallardo, C. P. Quesenberry, and Federico J. O’Reilly. Conditional probability integral transformations and goodness-of-fit tests for multivariate normal distributions. Annals of Statistics, 7 (5):1052–1057, September 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176344788.

Rhenius:1974:IIM

[Rhe74] Detlef Rhenius. Incomplete information in Markovian decision models. Annals of Statistics, 2(6):1327–1334, November 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176342886.

Revankar:1974:CPD

[RHP74] Nagesh S. Revankar, Michael J. Hartley, and Marcello Pagano. A characterization of the Pareto distribution. Annals of Statistics, 2 (3):599–601, May 1974. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176342723.

Rieder:1977:LFP

[Rie77] Helmut Rieder. Least favorable pairs for special capacities. Annals of Statistics, 5(5):909–921, September 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176343947.

Rieder:1978:RAT

[Rie78] Helmut Rieder. A robust asymptotic testing model. Annals of Statistics, 6(5):1080–1094, September 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176344312.

Robinson:1973:LSP

[Rob73] J. Robinson. The large-sample power of permutation tests for randomization models. Annals of Statistics, 1(2):291–296, March 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176342366.

Robinson:1975:CTR

[Rob75] P. M. Robinson. Continuous time regressions with discrete data. Annals of Statistics, 3(3):688–697, May 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176343131.

Robinson:1976:PSB

[Rob76] G. K. Robinson. Properties of Student’s t and of the Behrens–Fisher solution to the two means problem. Annals of Statistics, 4(5):963– 971, September 1976. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.

aos/1176343594. See correction [Rob82].

Robinson:1978:AES

[Rob78a] J. Robinson. An asymptotic expansion for samples from a finite population. Annals of Statistics, 6(5):1005–1011, September 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176344306.

Robinson:1978:CTS

[Rob78b] P. M. Robinson. On consistency in time series analysis. Annals of Statistics, 6(1):215–223, January 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176344080.

Robinson:1979:CPSa

[Rob79a] G. K. Robinson. Conditional properties of statistical procedures. Annals of Statistics, 7(4):742–755, July 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176344725.

Robinson:1979:CPSb

[Rob79b] G. K. Robinson. Conditional properties of statistical procedures for location and scale parameters. Annals of Statistics, 7(4):756– 771, July 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176344726.

Robinson:1979:DLA

[Rob79c] P. M. Robinson. Distributed lag approximation to linear timeinvariant systems. Annals of Statistics, 7(3):507–515, May 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176344672.

Robinson:1982:CPS

[Rob82] G. K. Robinson. Corrections: “Properties of Student’s t and of the Behrens–Fisher Solution to the Two Means Problem”. Annals of Statistics, 10(1):321, March 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176345718. See [Rob76].

Rosenblatt:1975:QMD

[Ros75a] M. Rosenblatt. A quadratic measure of deviation of two-dimensional density estimates and a test of independence. Annals of Statistics, 3 (1):1–14, January 1975. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.

aos/1176342996. See correction [Ros82].

Ross:1975:NOS

[Ros75b] Sheldon M. Ross. A note on optimal stopping for success runs. Annals of Statistics, 3(3):793–795, May 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176343147.

Rosenblatt:1982:CQM

[Ros82] M. Rosenblatt. Corrections: “A Quadratic Measure of Deviation of Two-Dimension Density Estimates and a Test of Independence”. Annals of Statistics, 10(2):646, June 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176345809. See [Ros75a].

Roy:1973:ECF

[Roy73] Roch Roy. Estimation of the covariance function of a homogeneous process on the sphere. Annals of Statistics, 1(4):780–785, July 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176342475.

Rafter:1974:CTC

[RS74a] J. A. Rafter and E. Seiden. Contributions to the theory and construction of balanced arrays. Annals of Statistics, 2(6):1256–1273, November 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176342877.

Rizvi:1974:IES

[RS74b] M. Haseeb Rizvi and K. M. Lal Saxena. On interval estimation and simultaneous selection of ordered location or scale parameters. Annals of Statistics, 2(6):1340–1345, November 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176342888.

Robbins:1974:ESS

[RS74c] H. Robbins and D. Siegmund. The expected sample size of some tests of power one. Annals of Statistics, 2(3):415–436, May 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176342704.

Ruiz:1974:CSF

[RS74d] Felipe Ruiz and Esther Seiden. On construction of some families of generalized Youden designs. Annals of Statistics, 2(3):503– 519, May 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176342710.

Rao:1975:WCE

[RS75] J. S. Rao and J. Sethuraman. Weak convergence of empirical distribution functions of random variables subject to perturbations and scale factors. Annals of Statistics, 3(2):299–313, March 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343058.

Rinott:1977:IMN

[RS77a] Yosef Rinott and Thomas J. Santner. An inequality for multivariate normal probabilities with application to a design problem. Annals of Statistics, 5(6):1228–1234, November 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176344007.

Rootzen:1977:EBS

[RS77b] Holger Rootzen and Gordon Simons. On the exponential boundedness of stopping times of invariant SPRT’s. Annals of Statistics, 5 (3):571–575, May 1977. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176343858.

Rao:1975:ESC

[RSL75] P. V. Rao, Eugene F. Schuster, and Ramon C. Littell. Estimation of shift and center of symmetry based on Kolmogorov–Smirnov statistics. Annals of Statistics, 3(4):862–873, July 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343187.

Rubin:1978:BIC

[Rub78] Donald B. Rubin. Bayesian inference for causal effects: The role of randomization. Annals of Statistics, 6(1):34–58, January 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176344064.

Rukhin:1978:UBE

[Ruk78] A. L. Rukhin. Universal Bayes estimators. Annals of Statistics, 6 (6):1345–1351, November 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176344379.

Ruppert:1979:NDS

[Rup79] David Ruppert. A new dynamic stochastic approximation procedure. Annals of Statistics, 7(6):1179–1195, November 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176344839.

Ruschendorf:1976:ADM

[Rus76a] Ludger Ruschendorf. Asymptotic distributions of multivariate rank order statistics. Annals of Statistics, 4(5):912–923, September 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343588. See

correction [Rus82].

Ruschendorf:1976:HGG

[Rus76b] Ludger Ruschendorf. Hypotheses generating groups for testing multivariate symmetry. Annals of Statistics, 4(4):791–795, July 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343553.

Ruschendorf:1982:CAD

[Rus82] L. Ruschendorf. Corrections: “Asymptotic Distributions of Multivariate Rank Order Statistics”. Annals of Statistics, 10(4):1311, December 1982. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/

1176345998. See [Rus76a].

Ruymgaart:1974:ANN

[Ruy74] F. H. Ruymgaart. Asymptotic normality of nonparametric tests for independence. Annals of Statistics, 2(5):892–910, September 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176342812.

Ruymgaart:1975:NCS

[Ruy75] F. H. Ruymgaart. A note on chi-square statistics with random Cell boundaries. Annals of Statistics, 3(4):965–968, July 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343198.

Ruymgaart:1978:ANM

[RvZ78] F. H. Ruymgaart and M. C. A. van Zuijlen. Asymptotic normality of multivariate linear rank statistics in the non-I.I.D. case. Annals of Statistics, 6(3):588–602, May 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176344203.

Robertson:1973:MIM

[RW73] Tim Robertson and F. T. Wright. Multiple isotonic median regression. Annals of Statistics, 1(3):422–432, May 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176342408.

Robertson:1974:NRP

[RW74a] Tim Robertson and F. T. Wright. A norm reducing property for isotonized Cauchy mean value functions. Annals of Statistics, 2(6): 1302–1307, November 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176342882.

Robertson:1974:MLE

[RW74b] Tim Robertson and F. T. Wright. On the maximum likelihood estimation of stochastically ordered random variates. Annals of Statistics, 2(3):528–534, May 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176342712.

Robertson:1975:CGI

[RW75] Tim Robertson and F. T. Wright. Consistency in generalized isotonic regression. Annals of Statistics, 3(2):350–362, March 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343061.

Robertson:1978:LRT

[RW78a] Tim Robertson and Edward J. Wegman. Likelihood ratio tests for order restrictions in exponential families. Annals of Statistics, 6 (3):485–505, May 1978. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176344195.

Rogers:1978:FSD

[RW78b] W. H. Rogers and T. J. Wagner. A finite sample distribution-free performance bound for local discrimination rules. Annals of Statistics, 6(3):506–514, May 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176344196.

Sacks:1975:AES

[Sac75] Jerome Sacks. An asymptotically efficient sequence of estimators of a location parameter. Annals of Statistics, 3(2):285–298, March 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343057.

Sager:1975:CNE

[Sag75] Thomas W. Sager. Consistency in nonparametric estimation of the mode. Annals of Statistics, 3(3):698–706, May 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343132.

Sager:1978:EMM

[Sag78] Thomas W. Sager. Estimation of a multivariate mode. Annals of Statistics, 6(4):802–812, July 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176344253.

Sampson:1975:CEF

[Sam75] Allan R. Sampson. Characterizing exponential family distributions by moment generating functions. Annals of Statistics, 3(3):747– 753, May 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343140.

Santner:1975:RSS

[San75] Thomas J. Santner. A restricted subset selection approach to ranking and selection problems. Annals of Statistics, 3(2):334–349, March 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343060.

Sarndal:1976:UMV

[Sär76] Carl Erik Särndal. On uniformly minimum variance estimation in finite populations. Annals of Statistics, 4(5):993–997, September 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343598.

Savage:1976:RRF

[Sav76] Leonard J. Savage. On rereading R. A. Fisher. Annals of Statistics, 4(3):441–500, May 1976. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176343456.

Saw:1973:EES

[Saw73] John G. Saw. Expectation of elementary symmetric functions of a Wishart matrix. Annals of Statistics, 1(3):580–582, May 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176342428.

Saxena:1976:DFT

[Sax76] K. M. Lal Saxena. Distribution-free tolerance intervals for stochastically ordered distributions. Annals of Statistics, 4(6):1210–1218, November 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343652.

Sen:1973:LBE

[SBS73] P. K. Sen, B. B. Bhattacharyya, and M. W. Suh. Limiting behavior of the extremum of certain sample functions. Annals of Statistics, 1(2):297–311, March 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176342367.

Chang:1979:DOC

[sC79] Der shin Chang. Design of optimal control for a regression problem. Annals of Statistics, 7(5):1078–1085, September 1979. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176344791. See

correction [CW80].

Schaafsma:1973:PCO

[Sch73a] Willem Schaafsma. Paired comparisons with order-effects. Annals of Statistics, 1(6):1027–1045, November 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176342555.

Scheffe:1973:STC

[Sch73b] Henry Scheffe. A statistical theory of calibration. Annals of Statistics, 1(1):1–37, January 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1193342379.

Schwartz:1973:TEI

[Sch73c] Richard Schwartz. Two examples of invariant Bayes procedures. Annals of Statistics, 1(3):551–552, May 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176342422.

Scholz:1974:CEL

[Sch74] Friedrich-Wilhelm Scholz. A comparison of efficient location estimators. Annals of Statistics, 2(6):1323–1326, November 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176342885.

Schoenfeld:1977:APT

[Sch77] David A. Schoenfeld. Asymptotic properties of tests based on linear combinations of the orthogonal components of the Cramér– von Mises statistic. Annals of Statistics, 5(5):1017–1026, September 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343956.

Scholz:1978:WMR

[Sch78a] Friedrich-Wilhelm Scholz. Weighted median regression estimates. Annals of Statistics, 6(3):603–609, May 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176344204.

Schwarz:1978:EDM

[Sch78b] Gideon Schwarz. Estimating the dimension of a model. Annals of Statistics, 6(2):461–464, March 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176344136.

Schach:1979:AFT

[Sch79a] S. Schach. An alternative to the Friedman test with certain optimality properties. Annals of Statistics, 7(3):537–550, May 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176344675.

Schal:1979:DPS

[Sch79b] Manfred Schal. On dynamic programming and statistical decision theory. Annals of Statistics, 7(2):432–445, March 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176344625.

Scott:1975:AUA

[Sco75] A. J. Scott. On admissibility and uniform admissibility in finite population sampling. Annals of Statistics, 3(2):489–491, March 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343077.

Sedransk:1977:RHK

[Sed77] J. Sedransk. Review: H. S. Konijn, Statistical Theory of Sample Survey Design and Analysis. Annals of Statistics, 5(2):419–428, March 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343810.

Sen:1973:FSC

[Sen73] Pranab Kumar Sen. On fixed size confidence bands for the bundle strength of filaments. Annals of Statistics, 1(3):526–537, May 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176342418.

Sen:1974:ASB

[Sen74a] Pranab Kumar Sen. Almost sure behaviour of U-statistics and von Mises’ differentiable statistical functions. Annals of Statistics, 2(2): 387–395, March 1974. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176342675.

Sen:1974:IPO

[Sen74b] Pranab Kumar Sen. The invariance principle for one-sample rankorder statistics. Annals of Statistics, 2(1):49–62, January 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176342612. See

correction [Sen74c].

Sen:1974:NCI

[Sen74c] Pranab Kumar Sen. Notes: Correction to “The Invariance Principle for One-Sample Rank-Order Statistics”. Annals of Statistics, 2(6):

1358, November 1974. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176342893. See [Sen74b].

Sengupta:1975:LSS

[Sen75] Sailes K. Sengupta. Lower semicontinuous stochastic games with imperfect information. Annals of Statistics, 3(2):554–558, March 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343088.

Sen:1976:WCP

[Sen76] Pranab Kumar Sen. Weak convergence of progressively censored likelihood ratio statistics and its role in asymptotic theory of life testing. Annals of Statistics, 4(6):1247–1257, November 1976. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343657.

Sen:1977:SIP

[Sen77a] Pranab Kumar Sen. Some invariance principles relating to jackknifing and their role in sequential analysis. Annals of Statistics, 5(2): 316–329, March 1977. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176343797.

Sen:1977:TWP

[Sen77b] Pranab Kumar Sen. Tied-down Wiener process approximations for aligned rank order processes and some applications. Annals of Statistics, 5(6):1107–1123, November 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176343999.

Sen:1979:APM

[Sen79a] Pranab Kumar Sen. Asymptotic properties of maximum likelihood estimators based on conditional specification. Annals of Statistics, 7 (5):1019–1033, September 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176344785.

Sen:1979:IPC

[Sen79b] Pranab Kumar Sen. Invariance principles for the coupon Collector’s problem: A martingale approach. Annals of Statistics, 7(2):372– 380, March 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176344620.

Sen:1979:WCS

[Sen79c] Pranab Kumar Sen. Weak convergence of some quantile processes arising in progressively censored tests. Annals of Statistics, 7(2): 414–431, March 1979. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176344624.

Serfling:1974:PIS

[Ser74] R. J. Serfling. Probability inequalities for the sum in sampling without replacement. Annals of Statistics, 2(1):39–48, January 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176342611.

Sen:1973:LIL

[SG73] Pranab Kumar Sen and Malay Ghosh. A law of iterated logarithm for one-sample rank order statistics and an application. Annals of Statistics, 1(3):568–576, May 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176342426.

Sen:1974:SRT

[SG74] Pranab Kumar Sen and Malay Ghosh. Sequential rank tests for location. Annals of Statistics, 2(3):540–552, May 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176342714. See cor-

rection [SG76b].

Sen:1976:CSB

[SG76a] P. K. Sen and B. K. Ghosh. Comparison of some bounds in estimation theory. Annals of Statistics, 4(4):755–765, July 1976. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343547. See

correction [SG77].

Sen:1976:NCS

[SG76b] Pranab Kumar Sen and Malay Ghosh. Notes: Correction to “Sequential Rank Tests for Location”. Annals of Statistics, 4(4):821, July 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/

1176343557. See [SG74].

Sen:1977:NCC

[SG77] P. K. Sen and B. K. Ghosh. Note: Correction to “Comparison of Some Bounds in Estimation Theory”. Annals of Statistics, 5(3): 593, May 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/

1176343862. See [SG76a].

Shapiro:1979:ANP

[SH79] C. P. Shapiro and Lawrence Hubert. Asymptotic normality of permutation statistics derived from weighted sums of bivariate functions. Annals of Statistics, 7(4):788–794, July 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176344728.

Sharma:1973:AET

[Sha73] Divakar Sharma. Asymptotic equivalence of two estimators for an exponential family. Annals of Statistics, 1(5):973–980, September 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176342519.

Shapiro:1974:BCA

[Sha74] C. P. Shapiro. Bayesian classification: Asymptotic results. Annals of Statistics, 2(4):763–774, July 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176342763.

Shaman:1975:AIC

[Sha75] Paul Shaman. An approximate inverse for the covariance matrix of moving average and autoregressive processes. Annals of Statistics, 3(2):532–538, March 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176343085.

Shaman:1976:ASC

[Sha76] Paul Shaman. Approximations for stationary covariance matrices and their inverses with application to ARMA models. Annals of Statistics, 4(2):292–301, March 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176343408.

Shaked:1977:CPD

[Sha77a] Moshe Shaked. A concept of positive dependence for exchangeable random variables. Annals of Statistics, 5(3):505–515, May 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343847.

Shapiro:1977:CMP

[Sha77b] C. P. Shapiro. Classification by maximum posterior probability. Annals of Statistics, 5(1):185–190, January 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176343752.

Shah:1978:RAR

[Sha78] S. M. Shah. On α-resolvability and affine α-resolvability of incomplete block designs. Annals of Statistics, 6(2):468–471, March 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176344138.

Shaked:1979:ESS

[Sha79a] Moshe Shaked. Estimation of starshaped sequences of Poisson and normal means. Annals of Statistics, 7(4):729–741, July 1979. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176344724.

Shapiro:1979:CEU

[Sha79b] Leonard Shapiro. Conditions for expected utility maximization: The finite case. Annals of Statistics, 7(6):1288–1302, November 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176344847.

Shea:1979:MRC

[She79] Gerald Shea. Monotone regression and covariance structure. Annals of Statistics, 7(5):1121–1126, September 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176344794.

Shirahata:1975:LMP

[Shi75] Shingo Shirahata. Locally most powerful rank tests for independence with censored data. Annals of Statistics, 3(1):241–245, January 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343014.

Shirahata:1976:ANR

[Shi76a] Shingo Shirahata. On the asymptotic normality of rank statistics for the two-sample problem. Annals of Statistics, 4(2):400–405, March 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343416.

Shirahata:1976:TII

[Shi76b] Shingo Shirahata. Tests for independence in infinite contingency tables. Annals of Statistics, 4(3):542–553, May 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343460.

Shirakura:1976:BFF

[Shi76c] Teruhiro Shirakura. Balanced fractional 2^{m} factorial designs of even resolution obtained from balanced arrays of strength 2l with index μ_{l} = 0. Annals of Statistics, 4(4):723–735, July 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343544.

Shirakura:1976:OBF

[Shi76d] Teruhiro Shirakura. Optimal balanced fractional 2^{m} factorial designs of resolution VII, 6 ≤ m ≤ 8. Annals of Statistics, 4(3): 515–531, May 1976. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176343458.

Shorack:1973:CRE

[Sho73] Galen R. Shorack. Convergence of reduced empirical and quantile processes with application to functions of order statistics in the non-I.I.D. case. Annals of Statistics, 1(1):146–152, January 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1193342391.

Shorack:1974:RM

[Sho74] Galen R. Shorack. Random means. Annals of Statistics, 2(4):661– 675, July 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176342755.

Sievers:1975:MPL

[Sie75] Gerald L. Sievers. Multivariate probabilities of large deviations. Annals of Statistics, 3(4):897–905, July 1975. CODEN ASTSC7.

ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176343190.

Siegmund:1976:ISM

[Sie76a] D. Siegmund. Importance sampling in the Monte Carlo study of sequential tests. Annals of Statistics, 4(4):673–684, July 1976. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343541.

Sievers:1976:PLD

[Sie76b] Gerald L. Sievers. Probabilities of large deviations for empirical measures. Annals of Statistics, 4(4):766–770, July 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343548.

Sievers:1978:ELL

[Sie78] Gerald L. Sievers. Estimates of location: A large deviation comparison. Annals of Statistics, 6(3):610–618, May 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176344205.

Silverman:1978:WSU

[Sil78] Bernard W. Silverman. Weak and strong uniform consistency of the kernel estimate of a density and its derivatives. Annals of Statistics, 6(1):177–184, January 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176344076.

Simar:1976:MLE

[Sim76a] Leopold Simar. Maximum likelihood estimation of a compound Poisson process. Annals of Statistics, 4(6):1200–1209, November 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343651.

Simons:1976:ISO

[Sim76b] Gordon Simons. An improved statement of optimality for sequential probability ratio tests. Annals of Statistics, 4(6):1240–1243, November 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343655.

Singh:1976:NPC

[Sin76a] Jagbir Singh. A note on paired comparison rankings. Annals of Statistics, 4(3):651–654, May 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176343474.

Singh:1976:EBE

[Sin76b] R. S. Singh. Empirical Bayes estimation with convergence rates in noncontinuous Lebesgue exponential families. Annals of Statistics, 4(2):431–439, March 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176343422.

Singh:1977:ISK

[Sin77] R. S. Singh. Improvement on some known nonparametric uniformly consistent estimators of derivatives of a density. Annals of Statistics, 5(2):394–399, March 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176343805.

Singh:1979:EBE

[Sin79] R. S. Singh. Empirical Bayes estimation in Lebesgue-exponential families with rates near the best possible rate. Annals of Statistics, 7(4):890–902, July 1979. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176344738.

Shah:1974:PCA

[SK74] B. K. Shah and C. G. Khatri. Proof of conjectures about the expected values of the elementary symmetric functions of a noncentral Wishart matrix. Annals of Statistics, 2(4):833–836, July 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176342773.

Smith:1975:DSS

[SK75] Furman H. Smith and George Kimeldorf. Discrete sequential search for one of many objects. Annals of Statistics, 3(4):906– 915, July 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343191.

Saha:1973:NTC

[SKD73] G. M. Saha, A. C. Kulshreshtha, and A. Dey. On a new type of mclass cyclic association scheme and designs based on the scheme. Annals of Statistics, 1(5):985–990, September 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176342521.

Skibinsky:1976:SUB

[Ski76] Morris Skibinsky. Sharp upper bounds for probability on an interval when the first three moments are known. Annals of Statistics, 4(1): 187–213, January 1976. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176343353.

Smith:1978:HH

[Smi78] Walter L. Smith. Harold Hotelling 1895–1973. Annals of Statistics, 6 (6):1173–1183, November 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176344369.

Schuster:1973:NNE

[SN73] E. F. Schuster and J. A. Narvarte. A new nonparametric estimator of the center of a symmetric distribution. Annals of Statistics, 1 (6):1096–1104, November 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176342559.

Snijders:1977:CCT

[Sni77] Tom Snijders. Complete class theorems for the simplest empirical Bayes decision problems. Annals of Statistics, 5(1):164–171, January 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343749.

Spiegelman:1979:ESS

[Spi79] Clifford Spiegelman. On estimating the slope of a straight line when both variables are subject to error. Annals of Statistics, 7(1):201– 206, January 1979. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176344565.

Spruill:1976:CSC

[Spr76a] M. C. Spruill. Cell selection in the Chernoff–Lehmann chi-square statistic. Annals of Statistics, 4(2):375–383, March 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343413.

Spruill:1976:CCS

[Spr76b] M. C. Spruill. A comparison of chi-square goodness-of-fit tests based on approximate Bahadur slope. Annals of Statistics, 4(2):409–412, March 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343418.

Shah:1976:OTT

[SRK76] K. R. Shah, D. Raghavarao, and C. G. Khatri. Optimality of two and three factor designs. Annals of Statistics, 4(2):419–422, March 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343420.

Srivastava:1976:IRF

[SRP76] J. N. Srivastava, B. L. Raktoe, and H. Pesotan. On invariance and randomization in fractional replication. Annals of Statistics, 4(2): 423–430, March 1976. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176343421.

Srivastava:1973:SCI

[SS73] M. S. Srivastava and A. K. Sen. On sequential confidence intervals based on Wilcoxon type estimates. Annals of Statistics, 1(6):1200– 1202, November 1973. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176342569.

Sackrowitz:1974:AMO

[SS74] H. Sackrowitz and W. Strawderman. On the admissibility of the M.L.E. for ordered binomial parameters. Annals of Statistics, 2 (4):822–828, July 1974. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176342771.

Sen:1975:TDC

[SS75] Ashish Sen and Muni S. Srivastava. On tests for detecting change in mean. Annals of Statistics, 3(1):98–108, January 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343001.

Saleh:1978:NEL

[SS78] A. K. Md. Ehsanes Saleh and Pranab Kumar Sen. Nonparametric estimation of location parameter after a preliminary test on regression. Annals of Statistics, 6(1):154–168, January 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176344074.

Spruill:1979:KWT

[SS79] M. C. Spruill and W. J. Studden. A Kiefer–Wolfowitz theorem in a stochastic process setting. Annals of Statistics, 7(6):1329–1332, November 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176344850.

Sakai:1979:RBF

[SST79] Hideaki Sakai, Takashi Soeda, and Hidekatsu Tokumaru. On the relation between fitting autoregression and periodogram with applications. Annals of Statistics, 7(1):96–107, January 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176344557.

Schwaller:1975:MCR

[ST75a] Richard L. Schwaller and Kenneth J. Tiahrt. A method of constrained randomization for p_{1}^{n_{1}}p_{2}^{n_{2}} ⋯ p_{k}^{n_{k}} factorials. Annals of Statistics, 3(5):1135–1148, September 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176343245.

Sekkappan:1975:CUA

[ST75b] Rm. Sekkappan and M. E. Thompson. On a class of uniformly admissible estimators for finite populations. Annals of Statistics, 3(2): 492–499, March 1975. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176343078.

Studden:1976:RPF

[ST76] William J. Studden and Jia-Yeong Tsay. Remez’s procedure for finding optimal designs. Annals of Statistics, 4(6):1271–1279, November 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343659.

Starr:1979:LEP

[Sta79] Norman Starr. Linear estimation of the probability of discovering a new species. Annals of Statistics, 7(3):644–652, May 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176344684.

Steiger:1969:BPK

[Ste69] W. L. Steiger. A best possible Kolmogoroff-type inequality for martingales and a characteristic property. Annals of Mathematical Statistics, 40(3):764–769, June 1969. CODEN AAS-TAD. ISSN 0003-4851 (print), 2168-8990 (electronic). URL http:

//projecteuclid.org/euclid.aoms/1177697586. See erratum

[Ste73].

Steiger:1973:CNC

[Ste73] W. L. Steiger. Correction notes: Correction to “A Best Possible Kolmogoroff-Type Inequality for Martingales and a Characteristic Property”. Annals of Statistics, 1(2):400, March 1973. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176342383. See

[Ste69].

Stephens:1976:ARG

[Ste76] M. A. Stephens. Asymptotic results for goodness-of-fit statistics with unknown parameters. Annals of Statistics, 4(2):357–369, March 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343411.

Stenger:1979:MAR

[Ste79] H. Stenger. A minimax approach to randomization and estimation in survey sampling. Annals of Statistics, 7(2):395–399, March 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176344622.

Stigler:1973:ADT

[Sti73] Stephen M. Stigler. The asymptotic distribution of the trimmed mean. Annals of Statistics, 1(3):472–477, May 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176342412.

Stigler:1974:LFO

[Sti74] Stephen M. Stigler. Linear functions of order statistics with smooth weight functions. Annals of Statistics, 2(4):676–693, July 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176342756. See

correction [Sti79].

Stigler:1976:CDR

[Sti76] Stephen M. Stigler. Contribution to discussion of “On rereading R. A. Fisher. Annals of Statistics, 4(3):498–500, May 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343456.

Stigler:1977:DRE

[Sti77] Stephen M. Stigler. Do robust estimators work with real data? Annals of Statistics, 5(6):1055–1098, November 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343997.

Stigler:1978:MSE

[Sti78] Stephen M. Stigler. Mathematical statistics in the early states. Annals of Statistics, 6(2):239–265, March 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176344123.

Stigler:1979:NCL

[Sti79] Stephen M. Stigler. Note: Correction to linear functions of order statistics with smooth weight functions. Annals of Statistics, 7(2): 466, March 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/

1176344630. See [Sti74].

Stone:1974:APE

[Sto74a] Charles J. Stone. Asymptotic properties of estimators of a location parameter. Annals of Statistics, 2(6):1127–1137, November 1974.

CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176342869.

Stone:1974:LDE

[Sto74b] M. Stone. Large deviations of empirical probability measures. Annals of Statistics, 2(2):362–366, March 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176342671.

Stone:1975:AML

[Sto75] Charles J. Stone. Adaptive maximum likelihood estimators of a location parameter. Annals of Statistics, 3(2):267–284, March 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343056.

Stone:1977:CNR

[Sto77] Charles J. Stone. Consistent nonparametric regression. Annals of Statistics, 5(4):595–620, July 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid. org/euclid.aos/1176343886. See discussion [BHO^{+}77, Ano77].

Strawderman:1973:PBM

[Str73a] William E. Strawderman. Proper Bayes minimax estimators of the multivariate normal mean vector for the case of common unknown variances. Annals of Statistics, 1(6):1189–1194, November 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176342567.

Stroud:1973:NCW

[Str73b] T. W. F. Stroud. Noncentral convergence of Wald’s large-sample test statistic in exponential families. Annals of Statistics, 1(1):161– 165, January 1973. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1193342393.

Strand:1974:CEC

[Str74a] Otto Neall Strand. Coefficient errors caused by using the wrong covariance matrix in the general linear model. Annals of Statistics, 2(5):935–949, September 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176342815.

Strawderman:1974:MEP

[Str74b] William E. Strawderman. Minimax estimation of powers of the variance of a normal population under squared error loss. Annals of Statistics, 2(1):190–198, January 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176342625.

Strasser:1978:ARA

[Str78] H. Strasser. Admissible representation of asymptotically optimal estimates. Annals of Statistics, 6(4):867–881, July 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176344260.

Sobel:1974:SCC

[SU74] Milton Sobel and V. R. R. Uppuluri. Sparse and crowded cells and Dirichlet distributions. Annals of Statistics, 2(5):977–987, September 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176342818.

Subrahmaniam:1975:ADS

[Sub75] Kocherlakota Subrahmaniam. On the asymptotic distributions of some statistics used for testing Σ_{1} = Σ_{2}. Annals of Statistics, 3 (4):916–925, July 1975. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176343192.

Sugiura:1973:ANN

[Sug73] Nariaki Sugiura. Asymptotic non-null distributions of the likelihood ratio criteria for covariance matrix under local alternatives. Annals of Statistics, 1(4):718–728, July 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176342466.

Susarla:1974:RCS

[Sus74] V. Susarla. Rate of convergence in the sequence-compound squareddistance loss estimation problem for a family of m-variate normal distributions. Annals of Statistics, 2(1):118–133, January 1974. CO-DEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176342618.

Susarla:1978:EBE

[SV78a] V. Susarla and J. Van Ryzin. Empirical Bayes estimation of a distribution (survival) function from right censored observations. Annals of Statistics, 6(4):740–754, July 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176344249.

Susarla:1978:LST

[SV78b] V. Susarla and J. Van Ryzin. Large sample theory for a Bayesian nonparametric survival curve estimator based on censored samples. Annals of Statistics, 6(4):755–768, July 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176344250.

Seiden:1978:GCG

[SW78] Esther Seiden and Ching-Jung Wu. A geometric construction of generalized Youden designs for ν a power of a prime. Annals of Statistics, 6(2):452–460, March 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176344135.

Sacks:1978:LEA

[SY78] Jerome Sacks and Donald Ylvisaker. Linear estimation for approximately linear models. Annals of Statistics, 6(5):1122–1137, September 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176344315.

Schuster:1979:CTN

[SY79] E. Schuster and S. Yakowitz. Contributions to the theory of nonparametric regression, with application to system identification. Annals of Statistics, 7(1):139–149, January 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176344560.

Shorrock:1976:IEG

[SZ76] R. W. Shorrock and J. V. Zidek. An improved estimator of the generalized variance. Annals of Statistics, 4(3):629–638, May 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343470.

Szatrowski:1979:AND

[Sza79] Ted H. Szatrowski. Asymptotic nonnull distributions for likelihood ratio statistics in the multivariate normal patterned mean and covariance matrix testing problem. Annals of Statistics, 7(4):823– 837, July 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176344732.

Szynal:1973:EHR

[Szy73] Dominik Szynal. An extension of the Hajek–Rényi inequality for one maximum of partial sums. Annals of Statistics, 1(4):740– 744, July 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176342468.

Takeuchi:1975:CPS

[TA75] Kei Takeuchi and Masafumi Akahira. Characterizations of prediction sufficiency (adequacy) in terms of risk functions. Annals of Statistics, 3(4):1018–1024, July 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176343206.

Takada:1979:FME

[Tak79] Y. Takada. A family of minimax estimators in some multiple regression problems. Annals of Statistics, 7(5):1144–1147, September 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176344798.

Tyan:1976:TNC

[TDT76] Shu-Gwei Tyan, Haluk Derin, and John B. Thomas. Two necessary conditions on the representation of bivariate distributions by polynomials. Annals of Statistics, 4(1):216–222, January 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343355.

Tracy:1973:MPP

[TG73] D. S. Tracy and B. C. Gupta. Multiple products of polykays using ordered partitions. Annals of Statistics, 1(5):913–923, September 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176342511.

Tracy:1974:GSO

[TG74] D. S. Tracy and B. C. Gupta. Generalized h-statistics and other symmetric functions. Annals of Statistics, 2(4):837–844, July 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176342774.

Tryon:1973:CNP

[TH73] Peter V. Tryon and Thomas P. Hettmansperger. A class of non-parametric tests for homogeneity against ordered alternatives. Annals of Statistics, 1(6):1061–1070, November 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176342557.

Thomas:1978:IEU

[TH78] James D. Thomas and Robert A. Hultquist. Interval estimation for the unbalanced case of the one-way random effects model. Annals of Statistics, 6(3):582–587, May 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176344202.

Thomsen:1975:THU

[Tho75] Ib Thomsen. Testing hypotheses in unbalanced variance components models for two-way layouts. Annals of Statistics, 3(1):257– 265, January 1975. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176343017.

Thompson:1978:SSE

[Tho78] M. E. Thompson. Stratified sampling with exchangeable prior distributions. Annals of Statistics, 6(5):1168–1169, September 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176344320.

Tong:1973:AOS

[Ton73] Yung Liang Tong. An asymptotically optimal sequential procedure for the estimation of the largest mean. Annals of Statistics, 1(1): 175–179, January 1973. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1193342396.

Tong:1977:OTC

[Ton77] Y. L. Tong. An ordering theorem for conditionally independent and identically distributed random variables. Annals of Statistics, 5(2):

274–277, March 1977. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176343794.

Tong:1978:ASR

[Ton78] Y. L. Tong. An adaptive solution to ranking and selection problems. Annals of Statistics, 6(3):658–672, May 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176344210.

Torgersen:1977:MPD

[Tor77a] Erik N. Torgersen. Mixtures and products of dominated experiments. Annals of Statistics, 5(1):44–64, January 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343739.

Torgersen:1977:PSW

[Tor77b] Erik N. Torgersen. Prediction sufficiency when the loss function does not depend on the unknown parameter. Annals of Statistics, 5(1):155–163, January 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176343748.

Tsay:1977:CTO

[Tsa77] Jia-Yeong Tsay. A convergence theorem in L-optimal design theory. Annals of Statistics, 5(4):790–794, July 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176343902.

Tukey:1958:BCQ

[Tuk58] John W. Tukey. Bias and confidence in not-quite large samples (abstract). Annals of Mathematical Statistics, 29(2):614, June 1958. CODEN AASTAD. ISSN 0003-4851 (print), 2168-

8990 (electronic). URL http://projecteuclid.org/euclid.

aoms/1177706647. This is the other half of the Quenouille–Tukey jackknife technique; see [Que49, Que56, Mil74a, Efr79].

Tusnady:1977:AOT

[Tus77] G. Tusnady. On asymptotically optimal tests. Annals of Statistics, 5(2):385–393, March 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176343804.

Tretter:1975:CND

[TW75] Marietta J. Tretter and G. William Walster. Central and noncentral distributions of Wilks’ statistic in manova as mixtures of incomplete beta functions. Annals of Statistics, 3(2):467–472, March 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343073.

Tretter:1979:CFI

[TW79] Marietta J. Tretter and G. William Walster. Continued fractions for the incomplete beta function: Additions and corrections. Annals of Statistics, 7(2):462–465, March 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176344629. See [Aro41].

Uosaki:1974:SGD

[Uos74] Katsuji Uosaki. Some generalizations of dynamic stochastic approximation processes. Annals of Statistics, 2(5):1042–1048, September 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176342825.

Uthoff:1973:MPS

[Uth73] Vincent A. Uthoff. The most powerful scale and location invariant test of the normal versus the double exponential. Annals of Statistics, 1(1):170–174, January 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1193342395.

Vardeman:1978:ASF

[Var78] Stephen B. Vardeman. Admissible solutions of finite state sequence compound decision problems. Annals of Statistics, 6(3):673–679, May 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176344211.

Vardi:1979:AOS

[Var79a] Y. Vardi. Asymptotic optimal sequential estimation: The Poisson case. Annals of Statistics, 7(5):1040–1051, September 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176344787.

Vardi:1979:AOC

[Var79b] Y. Vardi. Asymptotic optimality of certain sequential estimators. Annals of Statistics, 7(5):1034–1039, September 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176344786.

vanHouwelingen:1976:MEB

[vH76] J. C. van Houwelingen. Monotone empirical Bayes tests for the continuous one-parameter exponential family. Annals of Statistics, 4(5):981–989, September 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176343596.

Vitale:1979:RGM

[Vit79] Richard A. Vitale. Regression with given marginals. Annals of Statistics, 7(3):653–658, May 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176344685.

VanRyzin:1977:EBA

[VS77] J. Van Ryzin and V. Susarla. On the empirical Bayes approach to multiple decision problems. Annals of Statistics, 5(1):172–181, January 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343750.

vanZuylen:1976:SPE

[vZ76] M. C. A. van Zuylen. Some properties of the empirical distribution function in the non-I.I.D. case. Annals of Statistics, 4(2):406–408, March 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343417.

Wagner:1973:DEB

[Wag73] T. J. Wagner. Deleted estimates of the Bayes risk. Annals of Statistics, 1(2):359–362, March 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176342374.

Wahba:1974:RDS

[Wah74] Grace Wahba. Regression design for some equivalence classes of kernels. Annals of Statistics, 2(5):925–934, September 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176342814.

Wahba:1975:ISM

[Wah75a] Grace Wahba. Interpolating spline methods for density estimation I. Equi-spaced knots. Annals of Statistics, 3(1):30–48, January 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176342998.

Wahba:1975:OCP

[Wah75b] Grace Wahba. Optimal convergence properties of variable knot, kernel, and orthogonal series methods for density estimation. Annals of Statistics, 3(1):15–29, January 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176342997.

Walter:1977:PHS

[Wal77] Gilbert G. Walter. Properties of Hermite series estimation of probability density. Annals of Statistics, 5(6):1258–1264, November 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176344013.

Worthley:1974:GAC

[WB74] Reginald Worthley and K. S. Banerjee. A general approach to confounding plans in mixed factorial experiments when the number of levels of a factor is any positive integer. Annals of Statistics, 2 (3):579–585, May 1974. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176342720.

Walter:1979:PDE

[WB79] G. Walter and J. Blum. Probability density estimation using delta sequences. Annals of Statistics, 7(2):328–340, March 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176344617.

Weed:1974:STS

[WBG74] Harrison D. Weed, Ralph A. Bradley, and Z. Govindarajulu. Stopping times of some one-sample sequential rank tests. Annals of Statistics, 2(6):1314–1322, November 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176342884.

Wegman:1979:RSR

[WD79] Edward J. Wegman and H. I. Davies. Remarks on some recursive estimators of a probability density. Annals of Statistics, 7(2):316– 327, March 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176344616.

Weiss:1974:ASR

[Wei74] Lionel Weiss. The asymptotic sufficiency of a relatively small number of order statistics in tests of fit. Annals of Statistics, 2(4): 795–802, July 1974. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176342766.

Wei:1978:ABC

[Wei78] L. J. Wei. The adaptive biased coin design for sequential experiments. Annals of Statistics, 6(1):92–100, January 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176344068.

Wei:1979:GPU

[Wei79] L. J. Wei. The generalized Pólya’s urn design for sequential medical trials. Annals of Statistics, 7(2):291–296, March 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176344614.

Wellner:1977:DRL

[Wel77a] Jon A. Wellner. Distributions related to linear bounds for the empirical distribution function. Annals of Statistics, 5(5):1003–1016, September 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343955.

Wellner:1977:GCT

[Wel77b] Jon A. Wellner. A Glivenko–Cantelli theorem and strong laws of large numbers for functions of order statistics. Annals of Statistics, 5 (3):473–480, May 1977. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.

aos/1176343844. See correction [Wel78a].

Wellner:1977:LIL

[Wel77c] Jon A. Wellner. A law of the iterated logarithm for functions of order statistics. Annals of Statistics, 5(3):481–494, May 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343845.

Wellner:1978:CGC

[Wel78a] Jon A. Wellner. Correction to: A Glivenko–Cantelli theorem and strong laws of large numbers for functions of order statistics. Annals of Statistics, 6(6):1394, November 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176344384. See [Wel77b].

Wells:1978:MUD

[Wel78b] Daniel R. Wells. A monotone unimodal distribution which is not central convex unimodal. Annals of Statistics, 6(4):926–931, July 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176344268.

Wellner:1979:PTD

[Wel79] Jon A. Wellner. Permutation tests for directional data. Annals of Statistics, 7(5):929–943, September 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176344779.

West:1979:ULP

[Wes79] Sandra A. West. Upper and lower probability inferences for the logistic function. Annals of Statistics, 7(2):400–413, March 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176344623.

Whitt:1976:BDG

[Whi76] Ward Whitt. Bivariate distributions with given marginals. Annals of Statistics, 4(6):1280–1289, November 1976. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176343660.

Wieand:1976:CUW

[Wie76] Harry S. Wieand. A condition under which the Pitman and Bahadur approaches to efficiency coincide. Annals of Statistics, 4(5):1003– 1011, September 1976. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176343600.

Wijsman:1960:MPS

[Wij60] Robert A. Wijsman. A monotonicity property of the sequential probability ratio test. Annals of Mathematical Statistics, 31(3):677– 684, September 1960. CODEN AASTAD. ISSN 0003-4851 (print),

2168-8990 (electronic). URL http://projecteuclid.org/euclid.

aoms/1177705794. See correction [Wij75b].

Wijsman:1973:ACR

[Wij73] R. A. Wijsman. On the attainment of the Cramér–Rao lower bound. Annals of Statistics, 1(3):538–542, May 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176342419.

Wijsman:1975:EBS

[Wij75a] R. A. Wijsman. Exponentially bounded stopping time of the sequential t-test. Annals of Statistics, 3(4):1006–1010, July 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343204.

Wijsman:1975:NCM

[Wij75b] Robert A. Wijsman. Notes: Correction to “A Monotonicity Property of the Sequential Probability Ratio Test”. Annals of Statistics, 3(3):796, May 1975. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176343148. See [Wij60].

Wijsman:1977:GTA

[Wij77] R. A. Wijsman. A general theorem with applications on exponentially bounded stopping time, without moment conditions. Annals of Statistics, 5(2):292–315, March 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176343796.

Wijsman:1979:CAS

[Wij79] Robert A. Wijsman. Constructing all smallest simultaneous confidence sets in a given class with applications to MANOVA. Annals of Statistics, 7(5):1003–1018, September 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176344784.

Wind:1973:EBA

[Win73] Serge L. Wind. An empirical Bayes approach to multiple linear regression. Annals of Statistics, 1(1):93–103, January 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1193342385.

Winter:1975:RSC

[Win75] B. B. Winter. Rate of strong consistency of two nonparametric density estimators. Annals of Statistics, 3(3):759–766, May 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343142.

Withers:1975:CEP

[Wit75] C. S. Withers. Convergence of empirical processes of mixing rv’s on [0, 1]. Annals of Statistics, 3(5):1101–1108, September 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343242.

Weinert:1974:SIR

[WK74] Howard L. Weinert and Thomas Kailath. Stochastic interpretations and recursive algorithms for spline functions. Annals of Statistics, 2 (4):787–794, July 1974. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176342765.

Woodroofe:1974:MLE

[Woo74] Michael Woodroofe. Maximum likelihood estimation of translation parameter of truncated distribution II. Annals of Statistics, 2(3): 474–488, May 1974. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176342708.

Woodroofe:1977:SOA

[Woo77] Michael Woodroofe. Second order approximations for sequential point and interval estimation. Annals of Statistics, 5(5):984–995, September 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176343953.

Woodroofe:1978:LDL

[Woo78] Michael Woodroofe. Large deviations of likelihood ratio statistics with applications to sequential testing. Annals of Statistics, 6(1): 72–84, January 1978. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176344066.

Wu:1978:SAA

[Wu78] Chien-Fu Wu. Some algorithmic aspects of the theory of optimal designs. Annals of Statistics, 6(6):1286–1301, November 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176344374.

Weiss:1973:MLE

[WW73] L. Weiss and J. Wolfowitz. Maximum likelihood estimation of a translation parameter of a truncated distribution. Annals of Statistics, 1(5):944–947, September 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176342515.

Wu:1978:CGS

[WW78] Chien-Fu Wu and Henry P. Wynn. The convergence of general step-length algorithms for regular optimum design criteria. Annals of Statistics, 6(6):1273–1285, November 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176344373.

Wynn:1977:CSF

[Wyn77] H. P. Wynn. Convex sets of finite population plans. Annals of Statistics, 5(2):414–418, March 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176343809.

Yahav:1979:SP

[Yah79] Joseph A. Yahav. On a screening problem. Annals of Statistics, 7 (5):1140–1143, September 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.org/

euclid.aos/1176344797.

Yakowitz:1974:MHT

[Yak74] S. Yakowitz. Multiple hypothesis testing by finite memory algorithms. Annals of Statistics, 2(2):323–336, March 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176342666.

Yakowitz:1979:NEM

[Yak79] Sidney Yakowitz. Nonparametric estimation of Markov transition functions. Annals of Statistics, 7(3):671–679, May 1979. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176344687.

Yang:1977:GDT

[Yan77] S. S. Yang. General distribution theory of the concomitants of order statistics. Annals of Statistics, 5(5):996–1002, September 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic).

URL http://projecteuclid.org/euclid.aos/1176343954.

Yang:1978:EBF

[Yan78] Grace L. Yang. Estimation of a biometric function. Annals of Statistics, 6(1):112–116, January 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176344070.

Yohai:1977:ABL

[YM77] Victor J. Yohai and Ricardo A. Maronna. Asymptotic behavior of least-squares estimates for autoregressive processes with infinite variances. Annals of Statistics, 5(3):554–560, May 1977. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343855.

Yohai:1979:ABE

[YM79] Victor J. Yohai and Ricardo A. Maronna. Asymptotic behavior of M-estimators for the linear model. Annals of Statistics, 7(2): 258–268, March 1979. CODEN ASTSC7. ISSN 0090-5364 (print),

2168-8966 (electronic). URL http://projecteuclid.org/euclid.
aos/1176344610.

Yohai:1973:AOB

[Yoh73] Victor J. Yohai. Asymptotically optimal Bayes sequential design of experiments for estimation. Annals of Statistics, 1(5):822–837, September 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176342504.

Yohai:1974:REL

[Yoh74] Victor J. Yohai. Robust estimation in the linear model. Annals of Statistics, 2(3):562–567, May 1974. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://projecteuclid.

org/euclid.aos/1176342717.

Zamir:1973:NVG

[Zam73] Shmuel Zamir. On the notion of value for games with infinitely many stages. Annals of Statistics, 1(4):791–796, July 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176342477.

Zehnwirth:1975:MIE

[Zeh75] Benjamin Zehnwirth. Minimax interval estimators of location parameters. Annals of Statistics, 3(2):451–456, March 1975. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL

http://projecteuclid.org/euclid.aos/1176343070.

Zidek:1973:ESP

[Zid73] James V. Zidek. Estimating the scale parameter of the exponential distribution with unknown location. Annals of Statistics, 1(2):264– 278, March 1973. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-

8966 (electronic). URL http://projecteuclid.org/euclid.aos/
1176342364.

Zidek:1978:DUR

[Zid78] Jim Zidek. Deriving unbiased risk estimators of multinormal mean and regression coefficient estimators using zonal polynomials. Annals of Statistics, 6(4):769–782, July 1978. CODEN ASTSC7. ISSN 0090-5364 (print), 2168-8966 (electronic). URL http://

projecteuclid.org/euclid.aos/1176344251.